### v1

# Learning in Stackelberg Mean Field Games: A Non-Asymptotic Analysis

Abstract
We study policy optimization in Stackelberg mean field games (MFGs), a hierarchical framework for modeling the strategic interaction between a single leader and an infinitely large population of homogeneous followers. The objective can be formulated as a structured bi-level optimization problem, in which the leader needs to learn a policy maximizing its reward, anticipating the response of the followers. Existing methods for solving these (and related) problems often rely on restrictive independence assumptions between the leader’s and followers’ objectives, use samples inefficiently due to nested-loop algorithm structure, and lack finite-time convergence guarantees. To address these limitations, we propose AC-SMFG, a single-loop actor-critic algorithm that operates on continuously generated Markovian samples. The algorithm alternates between (semi-)gradient updates for the leader, a representative follower, and the mean field, and is simple to implement in practice. We establish the finite-time and finite-sample convergence of the algorithm to a stationary point of the Stackelberg objective. To our knowledge, this is the first Stackelberg MFG algorithm with non-asymptotic convergence guarantees. Our key assumption is a “gradient alignment” condition, which requires that the full policy gradient of the leader can be approximated by a partial component of it, relaxing the existing leader-follower independence assumption. Simulation results in a range of well-established economics environments demonstrate that AC-SMFG outperforms existing multi-agent and MFG learning baselines in policy quality and convergence speed.
1 Introduction
Mean field games (MFGs) provide a framework for studying the strategic interaction among an infinite number of rational agents, with a wide range of applications in resource allocation (Li et al., 2020), telecommunication (Narasimha et al., 2019), and power system optimization (Alasseur et al., 2020). An important extension of MFGs is Stackelberg mean field games (SMFGs), incorporating a hierarchical structure where a single leader agent influences a population of follower agents and enjoys a first-mover advantage. A prominent example of SMFG is optimal liquidation (Almgren and Chriss, 1997), in which a large institutional investor (leader) strategically sells assets, e.g., central bank treasuries, in a market while accounting for the reactions of smaller traders (followers) who adjust their behavior, e.g., bond purchasing, in response (Chen et al., 2024). Other examples arise in public policy domains such as taxation, mortgage regulation, and epidemic control (Zheng et al., 2020; Mi et al., 2024; Aurell et al., 2022), where the leader, typically a government, designs interventions to promote social welfare while anticipating and influencing the collective response of the population.
Despite the increasing number of empirical studies applying the SMFG framework to real-world problems, there remains a lack of practical and theoretically grounded algorithms for learning in such hierarchical environments. The central challenge lies in the inherently coupled dynamics among the leader, a representative follower, and the mean field representing the collective behavior of the follower population. This interdependence gives rise to a complex bi-level structure: the leader’s objective depends on the equilibrium response of the follower population, which in turn is shaped by the leader’s policy. Existing approaches (Dayanıklı and Laurière, 2024; Cui et al., 2024) to SMFGs (and the closely related major-minor MFGs) need to make strong assumptions to decouple the leader and followers, lack non-asymptotic guarantees for finding a solution, and empirically exhibit slow convergence or tend towards sub-optimal limit points, possibly due to cyclic behaviors. In contrast, our work relaxes the assumptions and establishes a fast finite-sample rate of convergence for a simple actor-critic algorithm to the stationary point of the SMFG objective.
Figure 1: Theoretical convergence rates. The proposed algorithm is the first to have non-asymptotic convergence guarantees (black). Existing algorithms in similar settings are only known to converge asymptotically (red). Convergence rate of a single-loop algorithm for bi-level optimization (Hong et al., 2023) under a non-convex upper-level objective and lower-level strong convexity (blue).
1.1 Main Contributions
We introduce AC-SMFG, a single-loop actor-critic algorithm for SMFGs which operates under continuously generated Markovian samples. Our first key contribution is to characterize the finite-time and finite-sample complexity of AC-SMFG. We achieve this through a multi-time-scale analysis, where the updates of the leader’s policy, the follower’s policy, and the mean field are coordinated on distinct time scales via properly chosen step sizes. We show that AC-SMFG converges to a stationary point of the Stackelberg objective with rate $\widetilde{{\cal O}}(k^{-1/2})$, under a new assumption termed “gradient alignment". Intuitively, the assumption allows the leader to improve its objective by acting to the best response from the followers. As the algorithm draws exactly two samples in each iteration, this translates to a sample complexity of the same order. This is the first time that non-asymptotic convergence guarantees are established for any SMFG algorithm, and also the first time that a single-loop SMFG algorithm is analyzed.
Notably, our convergence rate surpasses that of a comparable single-loop algorithm for bi-level optimization under lower-level strong convexity (Hong et al., 2023) ($\widetilde{{\cal O}}(k^{-1/2})$ vs $\widetilde{{\cal O}}(k^{-2/5})$). The policy optimization problem in a SMFG can be viewed as bi-level optimization, where the upper-level objective is non-convex and the lower-level problem is to solve a standard MFG that does not exhibit any convexity. The reason that we achieve a better rate in a more challenging setting is two-fold. First, as a key technical innovation, we develop an analytical argument to handle bi-level optimization with lower-level Polyak-Lojasiewicz (PL) condition, which the MFG objective satisfies with respect to the follower’s softmax policy parameter under regularization (Mei et al., 2020). This technique allows us to obtain the same convergence rate in SMFGs as in bi-level optimization with lower-level strong convexity, and may be of independent interest and applicable to general bi-level optimization methods. Second, we make the rate improvement by incorporating the latest advances in multi-time-scale stochastic approximation (Han et al., 2024), leveraging the smoothness condition of the reward and transition.
We support the merit of proposed method through numerical simulations. We apply AC-SMFG to a range of environments inspired by real-world economic scenarios involving a prominent leader and a population of followers. Previous agent-based modeling (ABM) approaches to these problems simulate a large number of agents with complex interactions, leading to computational inefficiency and lack of theoretical guarantees (Evans et al., 2025). In contrast, we show that the SMFG formulation offers a more tractable alternative by summarizing the collective follower behavior through the mean field, thereby enabling faster computation of a (possibly local) optimum for the leader’s policy. Compared to the existing SMFG algorithms, AC-SMFG exhibits a significantly faster convergence rate.
1.2 Related Literature on Stackelberg Mean Field Games
(i) Continuous-time setting: (Djete, 2023) is among the first works to study SMFGs and to provide a characterization of the problem structure. Building on this, Dayanıklı and Laurière (2024) introduces a penalty-based approach that reformulates a SMFG as a single-level mean field control problem, offering a conceptual simplification. However, this approach does not come with a provably convergent learning algorithm, leaving open the question of how to efficiently and reliably solve SMFGs. (ii) Discrete-time setting: (Guo et al., 2022) consider a finite-horizon setup, and propose a minimax optimization framework (known model) for finding the policy of the leader considering a worst-case objective for the followers. Cui et al. (2024) consider finding an equilibrium in a (related, but distinct setup) of major-minor MFGs and proposes a nested-loop algorithm based on fictitious play, which approximately alternates best-response updates between the leader and the follower. Such a nested-loop structure often poses practical challenges, as it requires the inner loop (best response computation) to converge before each outer-loop update. This results in increased computational burden and reduced flexibility, particularly in large-scale or online settings where full convergence at each iteration may be infeasible or undesirable. Moreover, Cui et al. (2024) establishes only asymptotic convergence, offering limited insight into the algorithm’s performance in practice. Compared to the related literature discussed above, our paper is the first work to propose a single-loop algorithm for learning equilibria in SMFGs, supported by a non-asymptotic analysis.
2 Stackelberg Mean Field Game: Formulation
[…]
In this way, we can compactly represent a SMFG by the tuple $({\cal S},{\cal A},{\cal B},{\cal P},r_{f},r_{l},\gamma)$, where $\mathcal{P}:{\cal S}\times{\cal A}\times{\cal B}\times\Delta_{{\cal S}}\rightarrow\Delta_{{\cal S}}$ is the transition kernel and $\gamma\in(0,1)$ is the discount factor. The reward function of a representative follower is $r_{f}:\mathcal{S}\times{\cal A}\times{\cal B}\times\Delta_{\mathcal{S}}\rightarrow[0,1]$, with the understanding here that $\mathcal{S}$ is restricted to $\mathcal{S}_{f}$. Similarly, the reward of the leader $r_{l}:\mathcal{S}\times{\cal B}\times\Delta_{\mathcal{S}}\rightarrow[0,1]$ does not depend on the state of the representative follower $s_{f}$, but only on the mean field. The state and action spaces of all players is assumed to be finite to enable the analysis, but continuous space approximations are studied experimentally in Section˜5.
Followers’ Interaction.
The followers play a mean field game in response to a policy $\phi:{\cal S}\rightarrow\Delta_{{\cal B}}$ of the leader. As in a standard MFG, let the representative follower take actions according to a randomized policy $\pi:{\cal S}\rightarrow\Delta_{{\cal A}}$. Given a leader-follower policy pair $(\phi,\pi)$ and a mean field $\mu$, the sequentially generated states form a Markov chain, for which the transition matrix is $P^{\pi,\phi,\mu}\in\mathbb{R}^{|{\cal S}|\times|{\cal S}|}$. The matrix is entry-wise expressed as $P^{\pi,\phi,\mu}(s^{\prime}\mid s)=\sum_{a\in{\cal A},b\in{\cal B}}{\cal P}^{\mu}(s^{\prime}\mid s,a,b)\pi(a\mid s)\phi(b\mid s)$. We denote by $\nu^{\pi,\phi,\mu}$ the stationary distribution of the Markov chain, which is the singular vector of $P^{\pi,\phi,\mu}$ associated with the (only) largest singular value $1^{\prime\prime}$ (under an ergodicity assumption). Let the discounted visitation/occupancy measure $d_{\rho}^{\pi,\,\phi,\,\mu}\triangleq\mathbb{E}_{s_{0}\sim\rho}[d_{s}^{\pi,\,\phi,\,\mu}]$ under the initial state distribution $\rho$, where
$d_{s}^{\pi,\,\phi,\,\mu}\triangleq(1-\gamma)\mathbb{E}_{\pi,\phi,{\cal P}^{\mu}}[\textstyle\sum_{k=0}^{\infty}\gamma^{k}{\mathbf{1}}(s_{k}=s)\mid s_{0}=s],$
where the expectation is taken over
$a_{k}\sim\pi(\cdot\mid s_{k}),b_{k}\sim\phi(\cdot\mid s_{k}),s_{k+1}\sim{\cal P}^{\mu}(\cdot\mid s_{k},a_{k},b_{k}).$
Under $(\phi,\pi,\mu)$, the follower expects to collect the following (regularized) cumulative reward
$J_{f}(\pi,\phi,\mu)$ $\triangleq\textstyle\mathbb{E}_{\pi,\phi,{\cal P}^{\mu}}\left[\sum_{k=0}^{\infty}\gamma^{k}\big(r_{f}(s_{k},a_{k},b_{k},\mu)-\tau\log\pi(a_{k}\mid s_{k})\big)\mid s_{0}\sim\rho\right]$
$\textstyle=\frac{1}{1-\gamma}\mathbb{E}_{s\sim d_{\rho}^{\pi,\,\phi,\,\mu},\,a\sim\pi(\cdot\mid s),\,b\sim\phi(\cdot\mid s)}[r_{f}(s,a,b,\mu)+\tau E(\pi,s)],$ (1)
where the entropy function $E(\pi,s)\triangleq-\sum_{a}\pi(a\mid s)\log\pi(a\mid s)$, and regularization weight $\tau\geq 0$. Define the follower’s policy as $\pi^{\star}:\Delta_{\cal B}^{\cal S}\times\Delta_{\cal S}\rightarrow\Delta_{{\cal A}}^{{\cal S}}$ such that
$\pi^{\star}(\phi,\mu)$ $\triangleq\textstyle\mathop{\rm argmax}_{\pi\in\Delta_{\cal A}^{\cal S}}J_{f}(\pi,\phi,\mu),\quad\forall\mu.$ (2)
Let $\mu^{\star}:\Delta_{{\cal B}}^{{\cal S}}\rightarrow\Delta_{{\cal S}}$ denote the mapping from the leader’s policy to the mean field induced by the leader. It is known from the literature on standard MFGs that $\mu^{\star}(\phi)$ satisfies for all $\phi$
$\mu^{\star}(\phi)\triangleq\nu^{\pi^{\star}(\phi,\mu^{\star}(\phi)),\,\phi,\,\mu^{\star}(\phi)}.$ (3)
With some abuse of notation, we write
$\pi^{\star}(\phi)\triangleq\pi^{\star}(\phi,\mu^{\star}(\phi)),$ (4)
Fixing the leader’s policy to $\phi$ reduces the environment to a standard MFG among the followers. Using the terminology from the MFG literature, we refer to $(\pi^{\star}(\phi),\mu^{\star}(\phi))$ as the mean field equilibrium (MFE) for such MFG. We will later impose an assumption which guarantees that $\pi^{\star}(\phi)$ and $\mu^{\star}(\phi)$ are unique and well-defined.
Leader’s Interaction and Game Objective.
Given a leader’s policy $\phi$ and the follower’s mean field $\mu$, the leader’s cumulative reward is
$\textstyle J_{l}(\phi,\mu)\triangleq\mathbb{E}_{\pi,\phi,{\cal P}^{\mu}}\left[\sum_{k=0}^{\infty}\gamma^{k}r_{l}(s_{k},b_{k},\mu)\mid s_{0}\sim\rho\right]=\frac{1}{1-\gamma}\mathbb{E}_{s\sim d_{\rho}^{\pi,\,\phi,\,\mu},\,b\sim\phi(\cdot\mid s)}[r_{l}(s,b,\mu)].$ (5)
If the mean field $\mu$ were fixed, the aim of the leader would be to find a policy $\phi$ that maximizes $J_{l}(\phi,\mu)$. However, the stable mean field changes with $\phi$ as the followers try to best respond to the leader and each other. To solve a SMFG is to find an optimal policy for the leader, given that the followers best respond. We define $\Phi(\phi)=J_{l}(\phi,\mu^{\star}(\phi))$ for all $\phi$ and express the objective as follows
$\textstyle\phi^{\star}\triangleq\mathop{\rm argmax}_{\phi\in\Delta_{\cal B}^{\cal S}}\Phi(\phi)=\mathop{\rm argmax}_{\phi\in\Delta_{\cal B}^{\cal S}}J_{l}(\phi,\mu^{\star}(\phi)).$ (6)
3 Algorithm
The algorithm developed in this work is based on the principle of independent learning. To motivate our approach, consider a scenario where oracle knowledge is available on $\mu^{\star}(\phi)$ for any $\phi$. In this case, we can optimize $\phi$ iteratively with gradient descent – we maintain $\phi_{k}$ (where $k$ is the iteration index) and update $\phi_{k}$ in the direction of the gradient $\nabla\Phi(\phi_{k})$, which can be evaluated using $\mu^{\star}(\phi_{k})$.
In practice, oracle access to $\mu^{\star}(\phi_{k})$ is unavailable, and we must solve the lower-level MFG to approximate it. To this end, we introduce the iterates $\pi_{k}$ and $\hat{\mu}_{k}$ as estimates of $\pi^{\star}(\phi_{k})$ and $\mu^{\star}(\phi_{k})$, and refine them via (semi-)gradient descent. This naturally leads to a nested-loop structure, where the follower policy and mean field are updated to convergence in the inner loop to support the outer-loop gradient computation. Our approach eliminates this nested structure: we perform alternating updates of the leader, follower, and mean field in a single loop, using appropriately chosen step sizes to implicitly approximate a nested loop. The updates are formally stated in Algorithm 1.
We use a tabular softmax policy parameterization and maintain $\theta\in\mathbb{R}^{|{\cal S}|\times{\cal A}}$, $\omega\in\mathbb{R}^{|{\cal S}|\times{\cal B}}$ that encode the policies according to
$\phi_{\omega}(b\mid s)=\frac{\exp(\omega(s,b))}{\sum_{b^{\prime}\in{\cal B}}\exp(\omega(s,a^{\prime}))},\quad\pi_{\theta}(a\mid s)=\frac{\exp(\theta(s,a))}{\sum_{a^{\prime}\in{\cal A}}\exp(\theta(s,a^{\prime}))}.$
The tabular softmax parameterization is considered for the purpose of mathematical analysis – optimizing over the space of softmax parameters allows us to exploit the PL structure that each player’s objective observes with respect to its parameter (Mei et al., 2020). In practice, the proposed algorithm can be applied with function approximations such as neural networks, as discussed in detail in Remark 3. We can express the policy gradients of $\omega$ and $\theta$ in the closed form below (Sutton et al., 1999).
$\nabla_{\omega}J_{l}(\phi_{\omega},\mu)=\mathbb{E}_{\pi,\phi,{\cal P}^{\mu}}\left[\big(r_{l}(s,b,\mu)+\gamma V_{l}^{\phi_{\omega},\mu}(s^{\prime})-V_{l}^{\phi_{\omega},\mu}(s)\big)\nabla_{\omega}\log\phi_{\omega}(b\mid s)\right],$
$\nabla_{\theta}J_{f}(\pi_{\theta},\phi,\mu)\hskip-2.0pt=\hskip-2.0pt\mathbb{E}_{\pi,\phi,{\cal P}^{\mu}}\hskip-2.0pt\left[\big(r_{f}(s,a,b,\mu)\hskip-2.0pt-\hskip-2.0pt\tau\log\pi(a\mid s)\hskip-2.0pt+\hskip-2.0pt\gamma V_{f}^{\pi_{\theta},\phi,\mu}(s^{\prime})\hskip-2.0pt-\hskip-2.0ptV_{f}^{\pi_{\theta},\phi,\mu}(s)\big)\nabla_{\theta}\log\pi_{\theta}(a\hskip-2.0pt\mid\hskip-2.0pts)\right].$
The policy updates (7), (8) of Algorithm 2 are exactly based on the policy gradient expressions above, substituting in the latest leader’s policy parameter $\omega_{k}$, follower’s policy parameter $\theta_{k}$, and mean field iterate $\hat{\mu}_{k}$. Here the value functions are defined as
$\textstyle V_{l}^{\phi,\mu}\hskip-1.0pt(s)\hskip-2.0pt=\hskip-2.0pt\mathbb{E}[\sum_{k}^{\infty}\hskip-2.0pt\gamma^{k}r_{l}(s,b,\mu)\hskip-2.0pt\mid\hskip-2.0pts_{0}\hskip-2.0pt=\hskip-2.0pts],\;V_{f}^{\pi,\phi,\mu}(s)\hskip-2.0pt=\hskip-2.0pt\mathbb{E}[\sum_{k}^{\infty}\hskip-2.0pt\gamma^{k}(r_{f}(s,a,b,\mu)\hskip-2.0pt-\hskip-2.0pt\tau\log\pi(a\hskip-1.0pt\mid\hskip-1.0pts))\hskip-2.0pt\mid\hskip-2.0pts_{0}\hskip-2.0pt=\hskip-2.0pts].$
Algorithm 1 requires access to $V_{l}^{\phi_{\omega_{k}},\hat{\mu}_{k}}$ and $V_{f}^{\pi_{\theta_{k}},\phi_{\omega_{k}},\hat{\mu}_{k}}$ for policy gradient evaluation. Since these value functions are not directly available, we estimate them by $V_{l,k}$ and $V_{f,k}$, which are updated via temporal difference (TD) learning in (10). Note that to ensure stability we leverage projections in (9) and (10): $\Pi_{\Delta_{\cal S}}$ denotes the projection to the probability simplex over the state space, and $\Pi_{B_{V}}$ denotes the element-wise projection to the interval $0,B_{V}]$, where $B_{V}=1/(1-\gamma)+\tau\log|{\cal A}|$.
Our method exemplifies independent learning in the sense that $\omega_{k},\theta_{k},\hat{\mu}_{k}$ are each updated to optimize its own objective without any knowledge of the other two, though the samples used for the updates are from an environment determined by all variables. We note that the algorithm is single-loop and uses two trajectories of continuously generated samples [^1] The second trajectory of samples is used solely to estimate the mean field and may be eliminated in applications where there exists a generative model/oracle for the mean field., making it significantly more practical than the existing methods with complex nested loops (Dayanıklı and Laurière, 2024; Cui et al., 2024).
Algorithm 1 Single loop Actor-Critic Algorithm for Stackelberg Mean Field Games (AC-SMFG)
1:Initialize: leader’s policy parameter $\omega_{0}$, follower’s policy parameter $\theta_{0}$, value function estimates $\hat{V}_{l,0},\hat{V}_{f,0}$, mean field estimate $\hat{\mu}_{0}\in\Delta_{{\cal S}}$, initial state ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}s_{0}},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bar{s}_{0}}\sim\rho$, step sizes $\zeta_{k}\leq\xi_{k}\leq\alpha_{k}\leq\beta_{k}$.
2:for iteration $k=0,1,2,$... do
3:  Sample path 1 for tracking the discounted occupancy measure:Follower and leader take actions ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}a_{k}}\sim\pi_{\theta_{k}}(\cdot\mid{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}s_{k}})$, ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}b_{k}}\sim\phi_{\omega_{k}}(\cdot\mid{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}s_{k}})$, receive rewards $r_{f}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}s_{k}},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}a_{k}},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}b_{k}},\hat{\mu}_{k})$, and observe the next state according to the transition probability
$s_{k+1}\sim\begin{cases}\rho,&\text{with probability }1-\gamma\\ {\cal P}^{\hat{\mu}_{k}}(\cdot\mid{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}s_{k}},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}a_{k}},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}b_{k}}),&\text{with probability }\gamma\end{cases}$
4:  Sample path 2 for tracking the stationary distribution (mean field):Follower and leader take actions ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bar{a}_{k}}\sim\pi_{\theta_{k}}(\cdot\mid{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bar{s}_{k}})$, ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bar{b}_{k}}\sim\phi_{\omega_{k}}(\cdot\mid{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bar{s}_{k}})$, and observe the next state ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bar{s}_{k+1}}\sim{\cal P}^{\hat{\mu}_{k}}(\cdot\mid{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bar{s}_{k}},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bar{a}_{k}},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bar{b}_{k}})$
5:  Leader’s policy update:
$\omega_{k+1}=\omega_{k}+\zeta_{k}\nabla_{\omega}\log\phi_{\omega_{k}}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}b_{k}}\mid{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}s_{k}})\Big(r_{l}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}s_{k}},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}b_{k}},\hat{\mu}_{k})+\gamma\hat{V}_{l,k}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}s_{k+1}})-\hat{V}_{l,k}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}s_{k}})\Big)$ (7)
6:  Follower’s policy (actor) update:
$\theta_{k+1}\hskip-2.0pt=\hskip-2.0pt\theta_{k}\hskip-2.0pt+\hskip-2.0pt\alpha_{k}\nabla_{\theta}\hskip-2.0pt\log\pi_{\theta_{k}}\hskip-1.0pt(a_{k}\hskip-2.0pt\mid\hskip-2.0pt{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}s_{k}})\big(r_{f}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}s_{k}},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}a_{k}},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}b_{k}},\hat{\mu}_{k})\hskip-2.0pt+\hskip-2.0pt\tau E(\pi_{\theta_{k}},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}s_{k}})\hskip-2.0pt+\hskip-2.0pt\gamma\hat{V}_{f,k}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}s_{k+1}})\hskip-2.0pt-\hskip-2.0pt\hat{V}_{f,k}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}s_{k}})\big)$ (8)
7:  Mean field update:
$\hat{\mu}_{k+1}=\Pi_{\Delta_{{\cal S}}}\big(\hat{\mu}_{k}+\xi_{k}(e_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bar{s}_{k}}}-\hat{\mu}_{k})\big)$ (9)
8:  Value function (critic) update:
$\hat{V}_{l,k+1}$ $=\Pi_{B_{V}}\Big(\hat{V}_{l,k}+\beta_{k}e_{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}s_{k}}}\big(r_{l}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}s_{k}},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}b_{k}},\hat{\mu}_{k})+\gamma\hat{V}_{l,k}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}s_{k+1}})-\hat{V}_{l,k}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}s_{k}})\big)\Big)$ (10)
$\hat{V}_{f,k+1}\hskip-2.0pt$ $=\hskip-2.0pt\Pi_{B_{V}}\Big(\hat{V}_{f,k}\hskip-2.0pt+\hskip-2.0pt\beta_{k}e_{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}s_{k}}}\hskip-2.0pt\big(r_{f}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}s_{k}},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}a_{k}},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}b_{k}},\hat{\mu}_{k})+\tau E(\pi_{\theta_{k}},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}s_{k}})+\gamma\hat{V}_{f,k}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}s_{k+1}})-\hat{V}_{f,k}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}s_{k}})\big)\Big)$
9:end for
4 Convergence Analysis
In this section, we establish the convergence of AC-SMFG. We first introduce the main technical assumptions, which can be segregated into the standard ones (Assumptions 15) required for finite-sample analysis of standard MFGs, and the additional regularity assumption, Assumption 6 (termed gradient alignment), required to deal with the more challenging setting of SMFGs. To illustrate that the gradient alignment assumption is weaker than those in the existing literature, we provide a non-trivial example that satisfies our assumption but not the stronger ones in prior works.
Assumption 1 (Lipschitz and Smooth Transition and Reward)
There exist bounded constants $L_{P},L_{r}$ such that for all $\pi,\phi,\mu_{1},\mu_{2},s,a,b$
$\|P^{\pi,\phi,\mu_{1}}-P^{\pi,\phi,\mu_{2}}\|\leq L_{P}\|\mu_{1}-\mu_{2}\|,\;\|\nabla_{\mu}P^{\pi,\phi,\mu_{1}}-\nabla_{\mu}P^{\pi,\phi,\mu_{2}}\|\leq L_{P}\|\mu_{1}-\mu_{2}\|,$
$|r_{l}(s,b,\mu_{1})-r_{l}(s,b,\mu_{2})|\leq L_{r}\|\mu_{1}-\mu_{2}\|,\;|r_{f}(s,a,b,\mu_{1})-r_{f}(s,a,b,\mu_{2})|\leq L_{r}\|\mu_{1}-\mu_{2}\|,$
$\|\nabla_{\mu}r_{l}(s,b,\mu_{1})-\nabla_{\mu}r_{l}(s,b,\mu_{2})\|\leq L_{r}\|\mu_{1}-\mu_{2}\|,$
$\|\nabla_{\mu}r_{f}(s,a,b,\mu_{1})-\nabla_{\mu}r_{f}(s,a,b,\mu_{2})\|\leq L_{r}\|\mu_{1}-\mu_{2}\|.$
The Lipschitz continuity is a common assumption made in the literature on MFGs (Cui and Koeppl, 2021b; Zeng et al., 2025; Cui et al., 2024). We additionally assume that they are smooth.
Assumption 2 (Lipschitz Best Response)
The best response operators $\mu^{\star}$ and $\pi^{\star}$ are Lipschitz and $\mu^{\star}$ has Lipschitz gradients, i.e. there exists a constant $L\in[1,\infty)$ such that for all $\phi,\phi^{\prime},\mu,\mu^{\prime}$
$\begin{gathered}\|\mu^{\star}(\phi)-\mu^{\star}(\phi^{\prime})\|\leq L\|\phi-\phi^{\prime}\|,\quad\|\nabla_{\phi}\mu^{\star}(\phi)-\nabla_{\phi}\mu^{\star}(\phi^{\prime})\|\leq L\|\phi-\phi^{\prime}\|,\\ \|\pi^{\star}(\phi,\mu)-\pi^{\star}(\phi^{\prime},\mu^{\prime})\|\leq L(\|\phi-\phi^{\prime}\|+\|\mu-\mu^{\prime}\|).\end{gathered}$ (13)
The condition (13) in fact can be shown to follow from the Lipschitz continuity and smoothness of the transition and reward function in Assumption 1. We impose Assumption 2 directly for simplicity.
Assumption 3 (Exploration)
There exists a constant $\rho_{\min}>0$ such that the initial state distribution satisfies $\rho(s)\geq\rho_{\min},\forall s\in{\cal S}$. In addition, the follower’s policy iterates are uniformly bounded away from zero, i.e. there exists a constant $p_{\min}$ such that $\pi_{\theta_{k}}(a\mid s)\geq p_{\min}$ for all $k\geq 0,s\in{\cal S},a\in{\cal A}$.
The first part of Assumption 3 implies that the discounted occupancy measure $d_{\rho}^{\pi,\phi,\mu}$ is bounded away from zero for any $\pi,\phi,\mu$ and is a standard assumption in RL (Agarwal et al., 2021; Mei et al., 2020), while the second part can be readily satisfied by capping the norm of $\theta_{k}$ away from infinity. The assumption guarantee sufficient exploration of every state and action pair.
Assumption 4 (Contraction)
There exists a constant $\delta\in(0,1)$ such that for any $\phi,\mu$
$\|\nu^{\pi^{\star}(\phi,\mu_{1}),\phi,\mu_{1}}-\nu^{\pi^{*}(\phi,\mu_{2}),\phi,\mu_{2}}\|\leq\delta\|\mu_{1}-\mu_{2}\|.$ (14)
Assumption 4 states that in a standard MFG determined by any fixed $\phi$, an “optimality-consistency operator contraction condition” holds, which implies the uniqueness of the MFE. This is a key assumption made in the existing literature on MFGs (Xie et al., 2021; Zaman et al., 2023; Yardim et al., 2023) and can be guaranteed when the regularization weight $\tau$ is large enough.
Assumption 5 (Uniform Geometric Ergodicity)
For any $\pi,\phi,\mu$, the Markov chain $\{s_{t}\}$ generated by $P^{\pi,\,\mu}$ according to $s_{t+1}\sim P^{\pi,\,\phi,\,\mu}(\cdot\mid s_{t})$ is irreducible and aperiodic. In addition, we have
$\sup_{s}d_{\text{TV}}\big(\mathbb{P}(s_{t}=\cdot\mid s_{0}=s),\nu^{\pi,\,\phi,\,\mu}\big)\leq C_{0}C_{1}^{t},\,\forall t\geq 0,$
for some constants $C_{0}\geq 1$ and $C_{1}\in(0,1)$, where $d_{TV}$ denotes the total variation distance.
Assumption 5 is again standard in analyzing stochastic approximation and RL algorithms under Markovian samples (Zou et al., 2019; Wu et al., 2020; Wang et al., 2024).
Assumption 6 (Gradient Alignment)
There exist positive bounded constants $\eta_{1},\eta_{2}$ such that $\forall\omega$
$\eta_{1}\langle\nabla_{\omega}\Phi(\phi_{\omega}),\nabla_{\omega}J_{l}(\phi_{\omega},\mu)\mid_{\mu=\mu^{\star}(\phi_{\omega})}\rangle\geq\|\nabla_{\omega}\Phi(\phi_{\omega})\|^{2},$ (15)
$\eta_{2}\langle\nabla_{\omega}\Phi(\phi_{\omega}),\nabla_{\omega}J_{l}(\phi_{\omega},\mu)\mid_{\mu=\mu^{\star}(\phi_{\omega})}\rangle\geq\|\nabla_{\omega}J_{l}(\phi_{\omega},\mu)\mid_{\mu=\mu^{\star}(\phi_{\omega})}\|^{2}.$ (16)
Note that $\nabla_{\omega}J_{l}(\phi_{\omega},\mu)\mid_{\mu=\mu^{\star}(\phi_{\omega})}$ is a partial component of the full gradient $\nabla_{\omega}\Phi(\phi_{\omega})$. The condition states that the two gradients always make an acute angle. Without loss of generality, we assume $\eta_{1},\eta_{2}\geq 1$, as if the conditions hold with $\eta_{1},\eta_{2}<1$ they also hold with $\eta_{1},\eta_{2}=1$. Conceptually, the condition allows the leader to improve its objective by acting to the best response from the followers. Prior works on SMFGs and the closely related major-minor MFGs need to make assumptions of a similar nature, but more restrictive. For example, Cui et al. (2024) by their Assumptions 4.c) and 4.d) assumes that the leader’s reward is independent of the mean field and that the follower’s reward and transition are independent of the leader’s action. Under such conditions, the leader and follower are effectively decoupled, reducing the hierarchical structure of the problem to a near-independent setting. In contrast, with the gradient alignment assumption we permit our leader to depend on $\mu$, and follower rewards and transitions to depend on $b$. In Appendix E, we present a non-trivial MFG which satisfies the gradient alignment assumption but not the leader-follower independence condition in Cui et al. (2024).
4.1 Algorithm Complexity
Each variable in Algorithm 1 converges in the sense that an associated residual (measure of sub-optimality gap) decays to zero. The residual of the leader is the squared policy gradient norm. The mean field convergence is measured by its deviation from the (unique) equilibrium of a standard MFG induced by leader’s latest policy iterate. The convergence of the follower is assessed by the objective function gap relative to the best response against $\pi_{k},\hat{\mu}_{k}$. Finally, the value function estimates are evaluated by their $\ell_{2}$ distance to the true value functions under the latest policy and mean field iterates. The proof of Theorem 1 relies on analyzing the convergence of all variables jointly, through a coupled Lyapunov function.
$\begin{gathered}\varepsilon_{k}^{\phi}=\|\nabla_{\omega}\Phi(\phi_{\omega_{k}})\|^{2},\quad\varepsilon_{k}^{\pi}=J_{f}(\pi^{\star}(\phi_{\omega_{k}},\hat{\mu}_{k}),\phi_{\omega_{k}},\hat{\mu}_{k})-J_{f}(\pi_{\theta_{k}},\phi_{\omega_{k}},\hat{\mu}_{k}),\\ \varepsilon_{k}^{\mu}=\|\hat{\mu}_{k}-\mu^{\star}(\phi_{\omega_{k}})\|^{2},\quad\varepsilon_{l,k}^{V}=\|\hat{V}_{l,k}-V_{l}^{\phi_{\omega_{k}},\,\hat{\mu}_{k}}\|^{2},\quad\varepsilon_{f,k}^{V}=\|\hat{V}_{f,k}-V_{f}^{\pi_{\theta_{k}},\phi_{\omega_{k}},\,\hat{\mu}_{k}}\|^{2}.\end{gathered}$ (19)
Theorem 1
Consider the iterates of Algorithm 1 under the step sizes
$\zeta_{k}=\frac{c_{\zeta}}{(k+1)^{1/2}},\quad\alpha_{k}=\frac{c_{\alpha}}{(k+1)^{1/2}},\quad\xi_{k}=\frac{c_{\xi}}{(k+1)^{1/2}},\quad\beta_{k}=\frac{c_{\beta}}{(k+1)^{1/2}},$
with the properly selected constants $c_{\zeta},c_{\alpha},c_{\xi},c_{\beta}$. Under Assumptions 1-5, we have for all $k>0$,
$\min_{t<k}\mathbb{E}[\varepsilon_{t}^{\phi}]\leq\widetilde{{\cal O}}\left(\frac{1}{(k+1)^{\frac{1}{2}}}\right).$
As the residuals are all non-negative, Theorem 1 implies that the best iterate of Algorithm 1 converges to a stationary point of the leader’s objective with rate $\widetilde{{\cal O}}(k^{-1/2})$, where $\widetilde{{\cal O}}$ hides constants and logarithmic factors of $k$. As exactly two samples are drawn in each iteration, this translates to a sample complexity of the same order. To our knowledge, this is the first result providing a finite-time and sample complexity for learning in Stackelberg/major-minor mean field games. In contrast to the most relevant prior work (Cui et al., 2024), which establishes the asymptotic convergence of a nested-loop fictitious play method, our paper analyzes a single-loop algorithm and provides non-asymptotic guarantees. This rate improves upon that of a two-time-scale algorithm for bi-level optimization with a non-convex upper-level objective and lower-level strong convexity (Hong et al., 2023), which achieves $\widetilde{{\cal O}}(1/k^{2/5})$. Our improvement is obtained in a more challenging setting – the lower-level problem is not a strongly convex function but a MFG involving two coupled variables: the follower’s policy and the mean field. The MFG does not exhibit any convex structure, and only a non-uniform PL condition holds with respect to the follower’s policy under regularization. A key technical innovation made in our work is to handle such non-convexity, which we sketch in Remark 2, and which we believe is of independent interest and applicable to the analysis of general bi-level optimization methods with lower-level PL condition.
Note that our analysis relies on a sufficiently large regularization weight $\tau$ to ensure that Assumption 4 is satisfied. As a result, our method does not find the equilibrium of the original (unregularized) problem. In fact, solving unregularized MFGs – even in the standard, non-hierarchical setting – remains an open problem. Our work inherits this fundamental limitation.
Remark 1
It is known in the standard RL literature that the cumulative reward observes a “gradient domination” condition with respect to the policy parameter. Using the notation of our paper, it means that any stationary point (where $\nabla_{\omega}J_{l}(\phi_{\omega},\mu)=0$) is globally optimal under a fixed $\mu$. We are not able to extend the argument here and have to settle for convergence to a stationary point, because $\Phi$ is a composite function involving $\mu^{\star}$ and hence does not observe the gradient domination condition.
Remark 2
Two techniques allow us to achieve the convergence rate of $\widetilde{{\cal O}}(k^{-1/2})$, surpassing that in Hong et al. (2023) under lower-level strong convexity.
To analyze a bi-level optimization algorithm, we need to bound the residual of the follower’s policy and show its iteration-wise reduction. If the lower-level objective $J_{f}$ were strongly convex, we would select the residual to be $\tilde{\epsilon}_{k}^{\pi}=\|\pi_{\theta_{k}}-\pi^{\star}(\phi_{\omega_{k}},\hat{\mu}_{k})\|^{2}$ and show that $\tilde{\epsilon}_{k+1}^{\pi}-\tilde{\epsilon}_{k}^{\pi}$ is approximately negative. As the learning target shifts from iteration $k$ to $k+1$, bounding $\tilde{\epsilon}_{k+1}^{\pi}-\tilde{\epsilon}_{k}^{\pi}$ requires controlling the drift $\|\pi^{\star}(\phi_{\omega_{k+1}},\hat{\mu}_{k+1})-\pi^{\star}(\phi_{\omega_{k}},\hat{\mu}_{k})\|^{2}$, which can be easily bounded by
${\cal O}(\|\phi_{\omega_{k+1}}-\phi_{\omega_{k}}\|^{2}+\|\hat{\mu}_{k+1}-\hat{\mu}_{k}\|^{2})$ (20)
under the Lipschitz continuity of $\pi^{\star}$. However, without strong convexity, we cannot consider the residual $\tilde{\epsilon}_{k}^{\pi}$ (as $\pi^{\star}$ may not be unique in general) and instead employ $\epsilon_{k}^{\pi}$ defined in (19), a residual in the value function space. We would like to bound the learning target shift $J_{f}(\pi^{\star}(\phi_{\omega_{k+1}},\hat{\mu}_{k+1}),\phi_{\omega_{k+1}},\hat{\mu}_{k+1})-J_{f}(\pi^{\star}(\phi_{\omega_{k}},\hat{\mu}_{k}),\phi_{\omega_{k}},\hat{\mu}_{k})$, preferably still by (20) to avoid rate deterioration. However, the Lipschitz continuity of $J_{f}$ only yields a loose bound on the order of ${\cal O}(\|\phi_{\omega_{k+1}}-\phi_{\omega_{k}}\|+\|\hat{\mu}_{k+1}-\hat{\mu}_{k}\|)$, where the norms are not squared. Detailed in Proposition 3 (proof in Section C.3), we develop a careful error decomposition scheme to break down the residual difference $\epsilon_{k+1}^{\pi}-\epsilon_{k}^{\pi}$, and tightly bound the decomposed terms based on the PL condition, thereby avoiding the suboptimal bound implied by naive application of the Lipschitz property of $J_{f}$.
The technique highlighted above allows us to achieve the same convergence rate in SMFGs as in bi-level optimization with lower-level strong convexity. We further improve upon the rate in Hong et al. (2023), leveraging a recent advance in the multi-time-scale stochastic approximation literature. Specifically, Han et al. (2024) shows that a faster convergence rate can be achieved under lower-level strong convexity, if the lower-level learning target (equivalent to operators $\mu^{\star},\pi^{\star}$ in our context) has Lipschitz gradients. Our work adapts and extends the argument to the case where the lower-level objective is nonconvex.
5 Experiments
(a) Market Entrance (sell skew)
(b) Shop Positioning ($c=0.05$)
(c) Equilibrium Pricing ($i=5$)
Figure 2: Convergence across environments, demonstrating bootstrapped mean and 5% confidence interval across 30 runs for the leader reward (left) and follower rewards (right).
We conduct a comprehensive evaluation of the proposed methodology across a diverse set of canonical MFGs. Specifically, we extend three environments from MFGLib (Guo et al., 2023a), each exhibiting varying degrees of complexity. These environments, ordered by increasing difficulty, include Left-Right, Beach Bar, and Equilibrium pricing (Guo et al., 2023b), with descriptions in Appendix F. We also include in the appendix additional plots visualizing the evolution of the mean field as learning proceeds. All source code is available in the supplementary material.
Comparisons. The proposed approach is compared to the existing state-of-the-art methods for the discrete time setting: 1) a nested-loop MFG algorithm which alternates between training the leader and representative follower to convergence (OneByOne); 2) a weighted OneByOne update as presented in Algorithm 1 of (Cui et al., 2024) (CuiKoeppl); 3) a MARL implementation, where the leader and followers are updated simultaneously with PPO, as presented in (Evans et al., 2025) (ADAGE). In the MARL setting, there are two policies, one for the leader, and a shared policy (Vadori et al., 2020) for the followers, to keep the representation as fair as possible when compared to mean fields, and to prevent the need to learn $N$ independent follower policies.
(a) Market entrance
(b) Shop positioning
(c) Equilibrium Price
Figure 3: Equilibria Analysis
Market Entrance.
The first environment is a market entrance scenario inspired by the Left-Right game (Cui and Koeppl, 2021a), akin to minority games in physics and El Farol problems in complex systems, as detailed in Appendix F.1. Figure 3(a) illustrates the impact of various leader configurations, achieved by altering the leader’s objective, demonstrating how the leader learns to execute optimal actions while followers adapt effectively, steering the mean field towards the target state (see also Figure 6). The top row of Figure 3(a) highlights the leader’s influence on the resulting mean fields, while the bottom row depicts the market skew. The leader shifts the market dynamics to align with the desired behavior. In each scenario, we observe convergence towards a suitable MFE, with the MF adjusting appropriately to the task at hand. These findings underscore how follower behavior successfully adapts in response to the leader’s actions, consistently achieving appropriate equilibria.
Shop Positioning.
The shop positioning environment (Appendix F.2) is modeled after the beach bar MFG, where followers strive to position themselves near a desirable location $\star$ while avoiding crowded areas, and a leader is tasked with establishing a new location to compete with $\star$. We investigate how varying the crowd aversion parameter $c$ among followers influences both the resulting mean field and the leader’s chosen location. When crowd aversion is absent ($c=0$), the leader strategically places the new location at the same spot as $\star$, capitalizing on the existing customer base, aligning with known models of spatial competition where competitors cluster nearby (Figure 3(b) left). As followers’ focus shifts entirely to crowd aversion ($c\to\infty$, Figure 3(b) right), the leader places the shop uniformly across the state space, preventing any single state from becoming overly crowded. For intermediate crowd aversions ($0<c<\infty$), we observe a spectrum of behaviors, balancing between the two limiting equilibria. In each scenario, the MFE is appropriately adjusted. These findings demonstrate how the leader’s actions adapt automatically in response to the followers’ behavior.
Equilibrium Price.
Finally, we examine an equilibrium pricing environment (Appendix F.3), as outlined in (Guo et al., 2023a, b; Cousin et al., 2011). In this environment, followers are homogeneous firms producing an identical product, with prices determined by the endogenous supply-demand equilibrium. These firms must effectively manage their inventory, production, and replenishment processes, and the leaders role is to incentivize firms to maintain a specific target inventory level. We explore various target scenarios, ranging from promoting lean operations with minimal inventories (suitable for predictable periods with stable demand) to encouraging robust firms with substantial inventory reserves (to withstand crises or demand surges). As illustrated in Figure 3(c), the leader can effectively shift the mean field of the followers to achieve the desired inventory levels. This demonstrates the leader’s ability to implement appropriate penalties or bonuses to guide followers towards either building a buffer (Figure 3(c) right) or maintaining leaner operations (Figure 3(c) left).
(a) Leader rewards
(b) Follower rewards
(c) Inventory costs $c_{2}$
(d) Mean Inventory
Figure 4: Function approximation with continuous states/actions in Equilibrium Pricing (with $i=0$)
Key simulation findings. The convergence of the leader and follower rewards is compared in Figure 2. Overall, the proposed approach affords significantly faster leader convergence, and smoother convergence trajectories, characterized by reduced variability across runs and diminished intra-run volatility. Notably, the more the followers’ behavior is influenced by the leader’s actions, the greater the enhancement observed with the proposed approach. Breaking down the results per environment, we see significantly faster convergence for the proposed approach in the market entrance game than all the comparisons (Figure 2(a)), while also resulting in an improved final equilibria (here, governed by the leaders reward which attempts to shape the equilibrium). In the shop positioning game (Figure 2(b)), we see significantly faster convergence for the leader, albeit with slightly slower follower convergence, yet still showing substantial improvements over the MARL scenario for both leaders and followers. The slower follower convergence here is because the leader has a smaller impact on the followers behaviour, due to their being the fixed desirable location, so only a small part of the environment is changing (the additional location). In the equilibrium pricing game (Figure 2(c)), the proposed approach yields significantly better leader and follower rewards than existing SMFG algorithms. This environment is particularly challenging and exceeds the representational capacity of purely tabular methods due to the continuous state and action spaces. ADAGE leverages function approximation, parameterizing the leader’s and followers’ policies by a neural neural work. We discuss in Remark 3 how function approximation integrates with the proposed SMFG approach – we can parameterize the policies of the leader and representative follower by a neural network in a standard way, whereas for the mean field we make a distributional assumption and learn the parameters of the distribution. Specifically in the equilibrium pricing experiment, we assume that the mean field follows a Gaussian distribution and updates the mean and standard deviation of the mean field in AC-SMFG. The assumption may not be exactly accurate, which is a possible cause of the gap in leader’s rewards between ADAGE and AC-SMFG. Nevertheless, the approximation suffices to yield competitive performance, and the proposed algorithm closely tracks the performance of ADAGE and outperforms other SMFG baselines.
Remark 3 (Continuous State and Action Spaces & Function approximation)
In Algorithm 1, we see that there are three key features that can be substituted with function approximation techniques rather than the exact update schemes presented: 1) the updates of the policies, 2) the updates of the value functions, 3) the updates of the mean field. Here, we show how each of these can be achieved. For the leader and follower, we can convert the tabular policy and value function into continuous representations with neural network function approximations (with an actor network and critic network) – this is relatively straightforward and similar to how continuous state and action spaces are handled in standard RL. What poses more challenge is the mean field update: in the tabular case we track the mean field by a probability vector of dimension $|{\cal S}|$, but this does not scale to the large/continuous state space setting. For simplicity, in the experiments of this paper, we consider a Gaussian mean field $\mathcal{N}(\nu,\sigma)$ and update the distribution parameters towards the sampled state $s$ with
$\nu_{k+1}=\nu_{k}+\xi_{k}(s-\nu_{k}),\quad\sigma_{k+1}=\sigma_{k}+\xi_{k}((s-\nu_{k+1})^{2}-\sigma_{k}).$
However, more complex schemes could be utilized. For demonstration, we take the Equilibrium Price environment and assume that the leader takes a continuous action $c_{2}\in[-1,1]$ to set the cost term, and the follower’s inventory (and thus the state space) is positive, ordinal, and unbounded. An overview of the resulting learning dynamics is displayed in Figure 4, which shows rapid learning and matches or even surpasses the results shown in Figure 8(a) of the appendix. This highlights the compatibility and effectiveness of our approach when integrated with function approximation in complex environments, underscoring the practical versatility and applicability of our method across diverse settings.
6 Conclusion & Discussions
We proposed a single-loop actor-critic algorithm for learning equilibria in SMFGs, and provided the first known non-asymptotic analysis. Numerical simulations show the algorithm’s superior performance in various realistic problems, highlighting its potential for practical applications in hierarchical decision-making environments. Our findings contribute to the growing body of research on SMFGs, offering a scalable and theoretically grounded solution for learning equilibria in complex SMFGs.
Finally, we acknowledge certain limitations of the SMFG framework in modeling practical problems, particularly when compared to agent-based modeling (ABM) approaches that simulate a finite number of followers directly. One limitation stems from the assumption of homogeneity among followers, which is embedded in the mean field formulation. Another key consideration is that MFGs serve as asymptotic approximations of $N$-agent Markov games, with the approximation error diminishing as $N$ increases. In scenarios involving a small number of follower agents, the approximation error may become non-negligible, and alternative modeling paradigms such as ABM may be more appropriate.
Disclaimer
This paper was prepared for informational purposes [“in part” if the work is collaborative with external partners] by the Artificial Intelligence Research group of JPMorgan Chase & Co. and its affiliates (“JP Morgan”) and is not a product of the Research Department of JP Morgan. JP Morgan makes no representation and warranty whatsoever and disclaims all liability, for the completeness, accuracy or reliability of the information contained herein. This document is not intended as investment research or investment advice, or a recommendation, offer or solicitation for the purchase or sale of any security, financial instrument, financial product or service, or to be used in any way for evaluating the merits of participating in any transaction, and shall not constitute a solicitation under any jurisdiction or to any person, if such solicitation under such jurisdiction or to such person would be unlawful.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: We clearly state the main contributions in the abstract and introduction, which is to provide the first algorithm for solving Stackelberg mean field games with non-asymptotic convergence analysis, and to make the analysis under a relaxed assumption on the interaction between the leader and followers.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: The main limitation lies in the gradient alignment assumption, and we clearly discuss how it is restrictive but still relaxes the prior ones, which are even stronger.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: We clearly state all assumptions in Section 4. We provide the complete proofs of all theoretical results in the appendix.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We provide all source code for reproducing the experimental results.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: We plan to make the code publicly available. In the review phase the code is submitted as a part of the supplementary material.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: We provide the source code and the most important experiment details in the appendix.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: The figures show the boostrapped mean result over 30 repetitions, with a 5% confidence interval displayed as the shaded region.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [N/A]
  39. Justification: Numerical simulations conducted are small-scale and require minimal computational resources.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: The authors confirm that the research conforms with the Code of Ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [N/A]
  49. Justification: The authors do not anticipate any social impact of the work at the moment.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: No such threat is present to our knowledge.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: To implement the baseline algorithms, we used code from published works such as Evans et al. [2025]. These works are properly credited in our paper.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: We have properly documented and added comments for the source code.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: No crowdsourcing and/or research with human subjects involved.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: No such risk is present.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: We did not use LLM to generate any significant content of the paper.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.
Appendix
Appendix A Notations and Frequently Used Inequalities
We introduce some shorthand notations frequently used in the rest of the paper. The operators below abstract the (semi-)gradient of the leader, follower, mean field, and value functions.
$\begin{gathered}D(\omega,\mu,V_{l},s,b,s^{\prime})=\nabla_{\omega}\log\phi_{\omega}(b\mid s)\Big(r_{l}(s,b,\mu)+\gamma V_{l}(s^{\prime})-V_{l}(s)\Big),\\ F(\theta,\mu,V_{f},s,a,b,s^{\prime})=\nabla_{\theta}\log\pi_{\theta}(a\mid s)\Big(r_{f}(s,a,b,\mu)+\tau E(\pi_{\theta},s)+\gamma V_{f}(s^{\prime})-V_{f}(s)\Big),\\ H(\mu,s)=e_{s}-\mu,\\ G_{l}(\omega,\mu,V_{l},s,a,b,s^{\prime})=e_{s}\Big(r_{l}(s,b,\mu)+\gamma V_{l}(s^{\prime})-V_{l}(s)\Big),\\ G_{f}(\theta,\mu,V_{f},s,a,b,s^{\prime})=e_{s}\Big(r_{f}(s,a,b,\mu)+\tau E(\pi_{\theta},s)+\gamma V_{f}(s^{\prime})-V_{f}(s)\Big).\end{gathered}$ (26)
Given the notations in (26), we can re-write the variable updates in Algorithm 1.
$\omega_{k+1}=\omega_{k}+\zeta_{k}D(\omega_{k},\hat{\mu}_{k},\hat{V}_{l,k},s_{k},b_{k},s_{k}^{\prime}),$ (27)
$\theta_{k+1}=\theta_{k}+\alpha_{k}F(\theta_{k},\hat{\mu}_{k},\hat{V}_{f,k},s_{k},a_{k},b_{k},s_{k}^{\prime}),$ (28)
$\hat{\mu}_{k+1}=\Pi_{\Delta_{{\cal S}}}\Big(\hat{\mu}_{k}+\xi_{k}H(\hat{\mu}_{k},\bar{s}_{k})\Big),$ (29)
$\hat{V}_{l,k+1}=\Pi_{B_{V}}\Big(\hat{V}_{l,k}+\beta_{k}G_{l}(\omega_{k},\hat{\mu}_{k},\hat{V}_{l,k},s_{k},a_{k},b_{k},s_{k}^{\prime})\Big),$ (30)
$\hat{V}_{f,k+1}=\Pi_{B_{V}}\Big(\hat{V}_{f,k}+\beta_{k}G_{f}(\theta_{k},\hat{\mu}_{k},\hat{V}_{f,k},s_{k},a_{k},b_{k},s_{k}^{\prime})\Big).$ (31)
We also define the expected versions of the (semi-)gradient operators, where the expectation is taken over the stochastic samples from the stationary distribution.
$\bar{D}(\omega,\mu,V_{l})\triangleq\mathbb{E}_{s\sim d_{\rho}^{\pi,\phi_{\omega},\mu},a\sim\pi(\cdot\mid s),b\sim\phi_{\omega}(\cdot\mid s),s^{\prime}\sim{\cal P}^{\mu}(\cdot\mid s,a,b)}[D(\omega,\mu,V_{l},s,b,s^{\prime})],$ (32)
$\bar{F}(\theta,\omega,\mu,V_{f})\triangleq\mathbb{E}_{s\sim d_{\rho}^{\pi_{\theta},\phi_{\omega},\mu},a\sim\pi_{\theta}(\cdot\mid s),b\sim\phi_{\omega}(\cdot\mid s),s^{\prime}\sim{\cal P}^{\mu}(\cdot\mid s,a,b)}[F(\theta,\mu,V_{f},s,a,b,s^{\prime})],$ (33)
$\bar{H}(\pi,\omega,\mu)\triangleq\mathbb{E}_{\bar{s}\sim\nu^{\pi,\phi_{\omega},\mu}}[H(\mu,\bar{s})],$ (34)
$\bar{G}_{l}(\omega,\mu,V_{l})\triangleq\mathbb{E}_{s\sim d_{\rho}^{\pi,\phi_{\omega},\mu},a\sim\pi(\cdot\mid s),b\sim\phi_{\omega}(\cdot\mid s),s^{\prime}\sim{\cal P}^{\mu}(\cdot\mid s,a,b)}[G_{l}(\omega,\mu,V_{l},s,a,b,s^{\prime})],$ (35)
$\bar{G}_{f}(\theta,\omega,\mu,V_{f})\triangleq\mathbb{E}_{s\sim d_{\rho}^{\pi_{\theta},\phi_{\omega},\mu},a\sim\pi_{\theta}(\cdot\mid s),b\sim\phi_{\omega}(\cdot\mid s),s^{\prime}\sim{\cal P}^{\mu}(\cdot\mid s,a,b)}[G_{f}(\theta,\mu,V_{f},s,a,b,s^{\prime})].$ (36)
We re-iterate that due to the state separation discussed in Section 2, $\bar{D}$ and $\bar{G}_{l}$ are independent of the single follower’s policy $\pi$, though $\pi$ appears on the right hand side of (32) and (35).
We define the filtration ${\cal F}_{k}=\{s_{0},a_{0},b_{0},s_{0}^{\prime}\cdots,s_{k},a_{k},b_{k},s_{k}^{\prime},\bar{s}_{0},\cdots,\bar{s}_{k}\}$.
Unless otherwise noted, we use $\|\cdot\|$ to denote the $\ell_{2}$ norm of a vector and the operator norm of a matrix.
We may use $\nabla_{\omega}\Phi(\phi_{\omega})$ and $\nabla_{\omega}J_{l}(\phi_{\omega},\mu^{\star}(\phi_{\omega}))$ interchangeably in the rest of the paper, which is the true gradient of the leader. The partial gradient is denoted by $\nabla_{\omega}J_{l}(\phi_{\omega},\mu)\mid_{\mu=\mu^{\star}(\phi_{\omega})}$.
We also introduce a few frequently used lemmas. The first is on the Lipschitz continuity and smoothness of value functions.
Lemma 1
We have for any $\theta,\theta^{\prime},\omega,\omega^{\prime},\mu,\mu^{\prime}$
$|J_{f}(\pi_{\theta},\phi_{\omega},\mu)-J_{f}(\pi_{\theta^{\prime}},\phi_{\omega^{\prime}},\mu^{\prime})|\leq L_{V}(\|\theta-\theta^{\prime}\|+\|\phi_{\omega}-\phi_{\omega^{\prime}}\|+\|\mu-\mu^{\prime}\|),$ (37)
$\|V_{f}^{\pi_{\theta},\phi_{\omega},\mu}-V_{f}^{\pi_{\theta^{\prime}},\phi_{\omega^{\prime}},\mu^{\prime}}\|\leq L_{V}(\|\theta-\theta^{\prime}\|+\|\phi_{\omega}-\phi_{\omega^{\prime}}\|+\|\mu-\mu^{\prime}\|),$ (38)
$\|\nabla_{\theta}V_{f}^{\pi_{\theta},\phi_{\omega},\mu}-\nabla_{\theta}V_{f}^{\pi_{\theta^{\prime}},\phi_{\omega},\mu}\|\leq L_{VV}\|\theta-\theta^{\prime}\|,\;\|\nabla_{\omega}V_{f}^{\pi_{\theta},\phi_{\omega},\mu}-\nabla_{\omega}V_{f}^{\pi_{\theta},\phi_{\omega^{\prime}},\mu}\|\leq L_{VV}\|\omega-\omega^{\prime}\|,$ (39)
$\|\nabla_{\mu}V_{f}^{\pi_{\theta},\phi_{\omega},\mu}-\nabla_{\mu}V_{f}^{\pi_{\theta},\phi_{\omega},\mu^{\prime}}\|\leq L_{VV}\|\mu-\mu^{\prime}\|,$ (40)
$|J_{l}(\phi_{\omega},\mu)-J_{l}(\phi_{\omega^{\prime}},\mu^{\prime})|\leq L_{V}(\|\omega-\omega^{\prime}\|+\|\mu-\mu^{\prime}\|),$ (41)
$\|V_{l}^{\phi_{\omega},\mu}-V_{l}^{\phi_{\omega^{\prime}},\mu^{\prime}}\|\leq L_{V}(\|\omega-\omega^{\prime}\|+\|\mu-\mu^{\prime}\|),$ (42)
$\|\nabla_{\omega}V_{l}^{\phi_{\omega},\mu}-\nabla_{\omega}V_{l}^{\phi_{\omega^{\prime}},\mu}\|\leq L_{VV}\|\omega-\omega^{\prime}\|,\;\|\nabla_{\mu}V_{l}^{\phi_{\omega},\mu}-\nabla_{\mu}V_{l}^{\phi_{\omega},\mu^{\prime}}\|\leq L_{VV}\|\mu-\mu^{\prime}\|,$ (43)
where $L_{V}=\max\Big\{\frac{2}{(1-\gamma)^{2}},\frac{L_{P}\gamma\sqrt{|{\cal S}|}}{(1-\gamma)^{2}}+\frac{L_{r}}{1-\gamma}\Big\}$ and $L_{VV}=\max\Big\{\frac{8}{(1-\gamma)^{3}},\frac{2\gamma^{2}\sqrt{|{\cal S}|}L_{P}^{2}}{(1-\gamma)^{3}}+\frac{2\gamma\sqrt{|{\cal S}|}L_{P}}{(1-\gamma)^{2}}+2\gamma L_{P}L_{r}\Big\}$.
The second lemma guarantees the Lipschitz continuity of the stationary distribution (of the current and next states jointly) with respect to the policies and mean field.
Lemma 2
For any $\theta_{1},\theta_{2},\omega_{1},\omega_{2},\mu_{1},\mu_{2}$, we have
$\Big\|\sum_{s}\big(d_{\rho}^{\pi_{\theta_{1}},\,\phi_{\omega_{1}}\,\mu_{1}}(s)P^{\pi_{\theta_{1}},\,\phi_{\omega_{1}}\,\mu_{1}}(\cdot\mid s)-d_{\rho}^{\pi_{\theta_{2}},\,\phi_{\omega_{2}}\,\mu_{2}}(s)P^{\pi_{\theta_{2}},\,\phi_{\omega_{2}}\,\mu_{2}}(\cdot\mid s)\big)\Big\|$
$\leq\frac{1}{1-\gamma}\Big(L_{P}\|\mu_{1}-\mu_{2}\|+\|\theta_{1}-\theta_{2}\|+\|\omega_{1}-\omega_{2}\|\Big).$
The next lemma establishes the boundedness of the (semi-)gradient operators introduced in (26).
Lemma 3
Recall that $B_{V}$ is the entry-wise upper bound on the magnitude of the value function. We define the constants
$\begin{gathered}B_{D}=2(1+\gamma)B_{V}+2,\quad B_{F}=2(1+\gamma)B_{V}+2\tau\log|{\cal A}|+2,\\ B_{H}=2,\quad B_{G}=(1+\gamma)B_{V}+\tau\log|{\cal A}|+1.\end{gathered}$ (46)
We have for all $\theta,\omega,\mu,V_{l},V_{f},s,a,b,s^{\prime}$
$\|D(\omega,\mu,V_{l},s,b,s^{\prime})\|\leq B_{D},\quad\|F(\theta,\mu,V_{f},s,a,b,s^{\prime})\|\leq B_{F},$
$\|H(\mu,s)\|\leq B_{H},\quad\|G_{l}(\omega,\mu,V_{l},s,b,s^{\prime})\|\leq B_{G},\quad\|G_{f}(\theta,\mu,V_{f},s,a,b,s^{\prime})\|\leq B_{G}.$
Lemma 4 further shows that the (semi-)gradient operators introduced in (26) are Lipschitz.
Lemma 4
We define the constants
$\begin{gathered}L_{D}=L_{F}=\max\left\{1+\tau\log|{\cal A}|+(1+\gamma)B_{V}+\frac{(4+8\log|{\cal A}|)\tau}{(1-\gamma)^{3}}+\frac{B_{F}}{1-\gamma},2L_{r}+\frac{B_{F}L_{P}}{1-\gamma},\frac{B_{F}}{1-\gamma},4\right\},\\ L_{H}=\max\left\{\frac{L_{P}}{1-\gamma},1\right\},\quad L_{G}=\max\left\{\frac{(4+8\log|{\cal A}|)\tau}{(1-\gamma)^{3}}+\frac{B_{G}}{1-\gamma},L_{r}+\frac{B_{G}L_{P}}{1-\gamma},\frac{B_{G}}{1-\gamma},2\right\}.\end{gathered}$ (49)
We have for all $\theta_{1},\theta_{2},\pi_{1},\pi_{2},\omega_{1},\omega_{2},\mu_{1},\mu_{2},V_{1},V_{2}$
$\|\bar{D}(\omega_{1},\mu_{1},V_{1})-\bar{D}(\omega_{2},\mu_{2},V_{2})\|\leq L_{D}\left(\|\omega_{1}-\omega_{2}\|+\|\mu_{1}-\mu_{2}\|+\|V_{1}-V_{2}\|\right),$
$\|\bar{F}(\theta_{1},\omega_{1},\mu_{1},V_{1})-\bar{F}(\theta_{2},\omega_{2},\mu_{2},V_{2})\|\leq L_{F}\left(\|\theta_{1}-\theta_{2}\|+\|\omega_{1}-\omega_{2}\|+\|\mu_{1}-\mu_{2}\|+\|V_{1}-V_{2}\|\right),$
$\|\bar{H}(\pi_{1},\omega_{1},\mu_{1})-\bar{H}(\pi_{2},\omega_{2},\mu_{2})\|\leq L_{H}\left(\|\pi_{1}-\pi_{2}\|+\|\omega_{1}-\omega_{2}\|+\|\mu_{1}-\mu_{2}\|\right),$
$\|\bar{G}_{f}(\theta_{1},\omega_{1},\mu_{1},V_{1})-\bar{G}_{f}(\theta_{2},\omega_{2},\mu_{2},V_{2})\|\leq L_{G}\left(\|\theta_{1}-\theta_{2}\|+\|\omega_{1}-\omega_{2}\|+\|\mu_{1}-\mu_{2}\|+\|V_{1}-V_{2}\|\right),$
$\|\bar{G}_{l}(\omega_{1},\mu_{1},V_{1})-\bar{G}_{l}(\omega_{2},\mu_{2},V_{2})\|\leq L_{G}\left(\|\omega_{1}-\omega_{2}\|+\|\mu_{1}-\mu_{2}\|+\|V_{1}-V_{2}\|\right).$
Appendix B Proof of Main Results
Algorithm 2 Actor-Critic Algorithm for Hierarchical Mean Field Games (Simplified for Analysis)
1:Initialize: leader’s policy parameter $\omega_{0}$, follower’s policy parameter $\theta_{0}$, value function estimate $\hat{V}_{l,0},\hat{J}_{l,0},\hat{V}_{f,0},\hat{J}_{f,0}$, mean field estimate $\hat{\mu}_{0}\in\Delta_{{\cal S}}$, initial state $s_{0},\bar{s}_{0}\sim\rho$
2:for iteration $k=0,1,2,$... do
3:  Sample 1 for tracking the discounted occupancy measure:Get samples $s_{k}\sim d_{\rho}^{\pi_{\theta_{k}},\phi_{\omega_{k}},\hat{\mu}_{k}}$, $a_{k}\sim\phi_{\omega_{k}}(\cdot\mid s_{k})$, $b_{k}\sim\phi_{\omega_{k}}(\cdot\mid s_{k})$, $s_{k}^{\prime}\sim{\cal P}^{\hat{\mu}_{k}}(\cdot\mid s_{k},a_{k},b_{k})$ and receive reward $r_{f}(s_{k},a_{k},b_{k},\hat{\mu}_{k})$, $r_{l}(s_{k},b_{k},\hat{\mu}_{k})$.
4:  Sample 2 for tracking the stationary distribution (mean field):Get sample $\bar{s}_{k}\sim\nu^{\pi_{\theta_{k}},\phi_{\omega_{k}},\hat{\mu}_{k}}$.
5:  Leader’s policy update:
$\omega_{k+1}=\omega_{k}+\zeta_{k}\nabla_{\omega}\log\phi_{\omega_{k}}(b_{k}\mid s_{k})\Big(r_{l}(s_{k},b_{k},\hat{\mu}_{k})+\gamma\hat{V}_{l,k}(s_{k}^{\prime})-\hat{V}_{l,k}(s_{k})\Big)$ (50)
6:  Follower’s policy (actor) update:
$\theta_{k+1}\hskip-2.0pt=\hskip-2.0pt\theta_{k}\hskip-2.0pt+\hskip-2.0pt\alpha_{k}\nabla_{\theta}\hskip-2.0pt\log\pi_{\theta_{k}}\hskip-1.0pt(a_{k}\hskip-2.0pt\mid\hskip-2.0pts_{k})\big(r_{f}(s_{k},a_{k},b_{k},\hat{\mu}_{k})\hskip-2.0pt+\hskip-2.0pt\tau E(\pi_{\theta_{k}},s_{k})\hskip-2.0pt+\hskip-2.0pt\gamma\hat{V}_{f,k}(s_{k}^{\prime})\hskip-2.0pt-\hskip-2.0pt\hat{V}_{f,k}(s_{k})\big)$ (51)
7:  Mean field update:
$\hat{\mu}_{k+1}=\Pi_{\Delta_{{\cal S}}}\big(\hat{\mu}_{k}+\xi_{k}(e_{\bar{s}_{k}}-\hat{\mu}_{k})\big)$ (52)
8:  Value function (critic) update:
$\hat{V}_{l,k+1}$ $=\Pi_{B_{V}}\Big(\hat{V}_{l,k}+\beta_{k}e_{s_{k}}\big(r_{l}(s_{k},b_{k},\hat{\mu}_{k})+\gamma\hat{V}_{l,k}(s_{k}^{\prime})-\hat{V}_{l,k}(s_{k})\big)\Big)$ (53)
$\hat{V}_{f,k+1}\hskip-2.0pt$ $=\hskip-2.0pt\Pi_{B_{V}}\Big(\hat{V}_{f,k}\hskip-2.0pt+\hskip-2.0pt\beta_{k}e_{s_{k}}\hskip-2.0pt\big(r_{f}(s_{k},a_{k},b_{k},\hat{\mu}_{k})+\tau E(\pi_{\theta_{k}},s_{k})+\gamma\hat{V}_{f,k}(s_{k}^{\prime})-\hat{V}_{f,k}(s_{k})\big)\Big)$
9:end for
We analyze a slightly simplified version of Algorithm 1, which we present in Algorithm 2. The only difference between them is that Algorithm 2 replaces continuously generated Markovian samples with i.i.d. samples from the stationary distribution. Note that stochastic approximation and RL algorithms have been extensively studied under Markovian samples [Zou et al., 2019, Wu et al., 2020], and it is well-known that Markovian sampling only incurs an additional logarithm factor in the convergence rates. This simplification allows us to focus the presentation on the innovations we develop for MFGs, without diverting attention to well-understood technical details on Markovian samples.
We state the theorem on the convergence of Algorithm 2 below, with the exact conditions that the step sizes need to satisfy.
Theorem 2
Consider the iterates of Algorithm 2 under the step sizes
$\zeta_{k}=\frac{c_{\zeta}}{(k+1)^{1/2}},\quad\alpha_{k}=\frac{c_{\alpha}}{(k+1)^{1/2}},\quad\xi_{k}=\frac{c_{\xi}}{(k+1)^{1/2}},\quad\beta_{k}=\frac{c_{\beta}}{(k+1)^{1/2}},$
with the constants $c_{\zeta},c_{\alpha},c_{\xi},c_{\beta}$ selected such that
$\begin{gathered}\zeta_{l}\leq\xi_{k}\leq\alpha_{k}\leq\beta_{k}\leq 1,\quad\xi_{k}\leq\min\{\frac{1-\delta}{8L^{2}L_{H}^{2}},\frac{1}{1-\delta}\},\\ \alpha_{k}\leq\min\{\frac{1-\gamma}{4L_{VV}(B_{F}+B_{D}+B_{H})^{2}},\frac{2}{B_{D}^{2}+B_{H}^{2}}\},\quad\beta_{k}\leq\frac{1-\gamma}{L_{G}^{2}},\\ \frac{\zeta_{k}}{\xi_{k}}\leq\min\Big\{\frac{1-\delta}{64L^{2}\eta_{1}\eta_{2}^{2}},\frac{1-\delta}{24\eta_{1}L_{D}^{2}L_{V}^{2}},\frac{1-\delta}{8\sqrt{6}LL_{V}L_{D}},\frac{\eta_{1}(1-\delta)}{16L^{2}},\frac{1-\delta}{16B_{D}^{2}L},\frac{B_{H}}{B_{D}L},\frac{LL_{H}}{3L_{V}L_{D}}\Big\},\\ \frac{\zeta_{k}}{\alpha_{k}}\leq\min\Big\{\frac{(1-\gamma)\tau^{2}\rho_{\min}^{3}p_{\min}^{2}}{16|{\cal S}|L_{V}^{2}\eta_{1}\eta_{2}^{2}},\frac{(1-\delta)(1-\gamma)\tau^{2}\rho_{\min}^{3}p_{\min}^{2}}{192|{\cal S}|L^{2}L_{V}^{4}L_{D}^{2}L_{H}^{2}},\frac{(1-\gamma)\eta_{1}\tau^{2}\rho_{\min}^{3}p_{\min}^{2}L_{D}^{2}}{4|{\cal S}|L_{V}^{2}}\Big\},\\ \frac{\zeta_{k}}{\beta_{k}}\leq\min\Big\{\frac{1-\gamma}{96L_{V}^{2}\eta_{1}\eta_{2}^{2}},\frac{(1-\delta)(1-\gamma)}{384L^{2}L_{V}^{2}L_{H}^{2}},\frac{1-\gamma}{32\eta_{1}L_{D}^{2}},\frac{1-\gamma}{2\sqrt{6}L_{V}^{2}}\Big\},\\ \frac{\xi_{k}}{\alpha_{k}}\leq\min\Big\{\frac{(1-\gamma)(1-\delta)\tau\rho_{\min}^{3}p_{\min}^{2}}{352|{\cal S}|L_{H}^{2}},\frac{(1-\gamma)^{2}\tau\rho_{\min}^{2}p_{\min}^{2}}{384|{\cal S}|L_{V}^{2}L_{H}^{2}},\frac{(1-\gamma)\tau^{2}\rho_{\min}^{3}p_{\min}^{2}}{4|{\cal S}|L_{V}L_{D}L_{H}},\frac{B_{F}(B_{D}+B_{H})}{L_{V}(2B_{H}^{2}+3L_{V})}\Big\},\\ \frac{\alpha_{k}}{\beta_{k}}\leq\min\Big\{\frac{1-\gamma}{384L_{V}^{2}},\frac{1-\gamma}{8\sqrt{6}L_{V}L_{D}},\frac{2B_{G}}{\sqrt{6L_{V}^{2}+L_{VV}}(B_{F}+B_{D}+B_{H})},\frac{1-\gamma}{4\sqrt{6}L_{V}L_{D}}\Big\}.\end{gathered}$ (61)
Then, under Assumptions 1-4 and 6, we have for all $k>0$,
$\min_{t<k}\mathbb{E}[\varepsilon_{t}^{\phi}]\leq\frac{16}{c_{\zeta}(k+1)^{1/2}}\Big(J_{l}(\phi^{\star},\mu^{\star}(\phi^{\star}))-J(\phi_{\omega_{0}},\mu^{\star}(\phi_{\omega_{0}}))+\varepsilon_{0}^{\mu}+\varepsilon_{0}^{\pi}+\varepsilon_{l,0}^{V}+\varepsilon_{f,0}^{V}\Big)$
$+\frac{128c_{\alpha}\log(k+1)}{c_{\zeta}(k+1)^{1/2}}.\hskip-20.0pt$
We note that the step sizes satisfying (61) always exist, and the conditions are met by making the step sizes sufficiently small. To find the appropriate step sizes, we first make $c_{\beta}$ small enough that $\beta_{k}$ satisfies all of its upper bounds, then $c_{\xi}$ small enough with respect to $\beta_{k}$, followed by $c_{\alpha}$ with respect to $\xi_{k}$ and $\beta_{k}$, and eventually $c_{\zeta}$. Note that $c_{\zeta},c_{\alpha},c_{\xi},c_{\beta}$ are all chosen as polynomial functions of the structural parameters, and do not depend on $k$.
The proof of Theorem 1 relies on intermediate convergence bound on the iteration-wise convergence of the leader’s policy parameter, mean field, follower’s policy parameter, and value function estimates. We state the intermediate results in the propositions below and defer their proofs to Section C.
Proposition 1 (Leader Convergence)
Under Assumptions 1-4 and 6 and step sizes satisfying (61), we have for all $k\geq 0$,
$\mathbb{E}[J_{l}(\phi_{\omega_{k}},\mu^{\star}(\phi_{\omega_{k}}))-J_{l}(\phi_{\omega_{k+1}},\mu^{\star}(\phi_{\omega_{k+1}}))]$
$\hskip 50.0pt\leq-\frac{\zeta_{k}}{2\eta_{1}}\mathbb{E}[\varepsilon_{k}^{\phi}]+2\eta_{1}L_{D}^{2}L_{V}^{2}\zeta_{k}\mathbb{E}[\varepsilon_{k}^{\mu}]+2\eta_{1}L_{D}^{2}\zeta_{k}\mathbb{E}[\varepsilon_{l,k}^{V}]+\frac{L_{V}B_{D}^{2}\zeta_{k}^{2}}{2}.$
Proposition 2 (Mean Field Convergence)
Under Assumptions 1-4 and 6 and step sizes satisfying (61), we have for all $k\geq 0$,
$\mathbb{E}[\varepsilon_{k+1}^{\mu}]$ $\leq(1-\frac{1-\delta}{4}\xi_{k})\mathbb{E}[\varepsilon_{k}^{\mu}]+\frac{11L_{H}^{2}\xi_{k}}{(1-\delta)\tau\rho_{\min}}\mathbb{E}[\varepsilon_{k}^{\pi}]+\frac{8L^{2}\eta_{2}^{2}\zeta_{k}^{2}}{(1-\delta)\xi_{k}}\mathbb{E}[\varepsilon_{k}^{\phi}]$
$\hskip 50.0pt+\frac{32L^{2}L_{D}^{2}\zeta_{k}^{2}}{(1-\delta)\xi_{k}}\mathbb{E}[\varepsilon_{l,k}^{V}]+16B_{H}^{2}\xi_{k}^{2}.$
Proposition 3 (Follower Convergence)
Under Assumptions 1-4 and 6 and step sizes satisfying (61), we have for all $k\geq 0$,
$\mathbb{E}[\varepsilon_{k+1}^{\pi}]$ $\leq\mathbb{E}[\varepsilon_{k}^{\pi}]-\frac{\alpha_{k}}{16}\mathbb{E}[\|\nabla_{\theta}J_{f}(\pi_{\theta_{k}},\phi_{\omega_{k}},\hat{\mu}_{k})\|^{2}]+\frac{2|{\cal S}|L_{V}^{2}\eta_{2}^{2}\zeta_{k}^{2}}{(1-\gamma)\tau^{2}\rho_{\min}^{3}p_{\min}^{2}\alpha_{k}}\mathbb{E}[\varepsilon_{k}^{\phi}]$
$\hskip 20.0pt+\frac{16|{\cal S}|L^{2}L_{V}^{4}L_{D}^{2}L_{H}^{2}\xi_{k}^{2}}{(1-\gamma)\tau^{2}\rho_{\min}^{3}p_{\min}^{2}\alpha_{k}}\mathbb{E}[\varepsilon_{k}^{\mu}]+\frac{8|{\cal S}|L_{V}^{2}\zeta_{k}^{2}}{(1-\gamma)\tau^{2}\rho_{\min}^{3}p_{\min}^{2}\alpha_{k}}\mathbb{E}[\varepsilon_{l,k}^{V}]$
$\hskip 20.0pt+2L_{V}^{2}\alpha_{k}\mathbb{E}[\varepsilon_{f,k}^{V}]+3B_{F}(B_{D}+B_{H})\alpha_{k}\xi_{k}.$
Proposition 4 (Value Function Convergence)
Under Assumptions 1-4 and 6 and step sizes satisfying (61), we have for all $k\geq 0$,
$\mathbb{E}[\varepsilon_{f,k+1}^{V}]$ $\leq(1-\frac{(1-\gamma)\beta_{k}}{4})\mathbb{E}[\varepsilon_{f,k}^{V}]+\frac{6L_{V}^{2}\alpha_{k}^{2}}{(1-\gamma)\beta_{k}}\mathbb{E}[\|\nabla_{\theta}J_{f}(\pi_{\theta_{k}},\phi_{\omega_{k}},\hat{\mu}_{k})\|^{2}]+\frac{6L_{V}^{2}L_{H}^{2}\xi_{k}^{2}}{(1-\gamma)\beta_{k}}\mathbb{E}[\varepsilon_{k}^{\pi}]$
$\hskip 20.0pt+\frac{6L_{V}^{2}\eta_{2}^{2}\zeta_{k}^{2}}{(1-\gamma)\beta_{k}}\mathbb{E}[\varepsilon_{k}^{\phi}]+\frac{16L^{2}L_{V}^{2}L_{H}^{2}\xi_{k}^{2}}{(1-\gamma)\beta_{k}}\mathbb{E}[\varepsilon_{k}^{\mu}]+\frac{24L_{V}^{2}L_{D}^{2}\alpha_{k}^{2}}{(1-\gamma)\beta_{k}}\mathbb{E}[\varepsilon_{l,k}^{V}]+12B_{G}^{2}\beta_{k}^{2},$
$\mathbb{E}[\varepsilon_{l,k+1}^{V}]$ $\leq(1-\frac{(1-\gamma)\beta_{k}}{4})\mathbb{E}[\varepsilon_{l,k}^{V}]+\frac{6L_{V}^{2}\alpha_{k}^{2}}{(1-\gamma)\beta_{k}}\mathbb{E}[\|\nabla_{\theta}J_{f}(\pi_{\theta_{k}},\phi_{\omega_{k}},\hat{\mu}_{k})\|^{2}]+\frac{6L_{V}^{2}L_{H}^{2}\xi_{k}^{2}}{(1-\gamma)\beta_{k}}\mathbb{E}[\varepsilon_{k}^{\pi}]$
$\hskip 20.0pt+\frac{6L_{V}^{2}\eta_{2}^{2}\zeta_{k}^{2}}{(1-\gamma)\beta_{k}}\mathbb{E}[\varepsilon_{k}^{\phi}]+\frac{16L^{2}L_{V}^{2}L_{H}^{2}\xi_{k}^{2}}{(1-\gamma)\beta_{k}}\mathbb{E}[\varepsilon_{k}^{\mu}]+\frac{6L_{V}^{4}\zeta_{k}^{2}}{(1-\gamma)\beta_{k}}\mathbb{E}[\varepsilon_{f,k}^{V}]+12B_{G}^{2}\beta_{k}^{2}.$
B.1 Proof of Theorem 2
Collecting the results from Propositions 1-4, we have
$\mathbb{E}\left[\Big(J_{l}(\phi_{\omega_{k}},\mu^{\star}(\phi_{\omega_{k}}))-J_{l}(\phi_{\omega_{k+1}},\mu^{\star}(\phi_{\omega_{k+1}}))\Big)+\varepsilon_{k+1}^{\mu}+\varepsilon_{k+1}^{\pi}+\varepsilon_{f,k+1}^{V}+\varepsilon_{l,k+1}^{V}\right]$
$\leq-\frac{\zeta_{k}}{2\eta_{1}}\mathbb{E}[\varepsilon_{k}^{\phi}]+2\eta_{1}L_{D}^{2}L_{V}^{2}\zeta_{k}\mathbb{E}[\varepsilon_{k}^{\mu}]+2\eta_{1}L_{D}^{2}\zeta_{k}\mathbb{E}[\varepsilon_{l,k}^{V}]+\frac{L_{V}B_{D}^{2}\zeta_{k}^{2}}{2}$
$\hskip 20.0pt+(1-\frac{1-\delta}{4}\xi_{k})\mathbb{E}[\varepsilon_{k}^{\mu}]+\frac{11L_{H}^{2}\xi_{k}}{(1-\delta)\tau\rho_{\min}}\mathbb{E}[\varepsilon_{k}^{\pi}]+\frac{8L^{2}\eta_{2}^{2}\zeta_{k}^{2}}{(1-\delta)\xi_{k}}\mathbb{E}[\varepsilon_{k}^{\phi}]+\frac{32L^{2}L_{D}^{2}\zeta_{k}^{2}}{(1-\delta)\xi_{k}}\mathbb{E}[\varepsilon_{l,k}^{V}]+16B_{H}^{2}\xi_{k}^{2}$
$\hskip 20.0pt+\mathbb{E}[\varepsilon_{k}^{\pi}]-\frac{\alpha_{k}}{16}\mathbb{E}[\|\nabla_{\theta}J_{f}(\pi_{\theta_{k}},\phi_{\omega_{k}},\hat{\mu}_{k})\|^{2}]+\frac{2|{\cal S}|L_{V}^{2}\eta_{2}^{2}\zeta_{k}^{2}}{(1-\gamma)\tau^{2}\rho_{\min}^{3}p_{\min}^{2}\alpha_{k}}\mathbb{E}[\varepsilon_{k}^{\phi}]+\frac{16|{\cal S}|L^{2}L_{V}^{4}L_{D}^{2}L_{H}^{2}\xi_{k}^{2}}{(1-\gamma)\tau^{2}\rho_{\min}^{3}p_{\min}^{2}\alpha_{k}}\mathbb{E}[\varepsilon_{k}^{\mu}]$
$\hskip 20.0pt+\frac{8|{\cal S}|L_{V}^{2}\zeta_{k}^{2}}{(1-\gamma)\tau^{2}\rho_{\min}^{3}p_{\min}^{2}\alpha_{k}}\mathbb{E}[\varepsilon_{l,k}^{V}]+2L_{V}^{2}\alpha_{k}\mathbb{E}[\varepsilon_{f,k}^{V}]+3B_{F}(B_{D}+B_{H})\alpha_{k}\xi_{k}$
$\hskip 20.0pt+(1-\frac{(1-\gamma)\beta_{k}}{4})\mathbb{E}[\varepsilon_{f,k}^{V}]+\frac{6L_{V}^{2}\alpha_{k}^{2}}{(1-\gamma)\beta_{k}}\mathbb{E}[\|\nabla_{\theta}J_{f}(\pi_{\theta_{k}},\phi_{\omega_{k}},\hat{\mu}_{k})\|^{2}]+\frac{6L_{V}^{2}L_{H}^{2}\xi_{k}^{2}}{(1-\gamma)\beta_{k}}\mathbb{E}[\varepsilon_{k}^{\pi}]$
$\hskip 20.0pt+\frac{6L_{V}^{2}\eta_{2}^{2}\zeta_{k}^{2}}{(1-\gamma)\beta_{k}}\mathbb{E}[\varepsilon_{k}^{\phi}]+\frac{16L^{2}L_{V}^{2}L_{H}^{2}\xi_{k}^{2}}{(1-\gamma)\beta_{k}}\mathbb{E}[\varepsilon_{k}^{\mu}]+\frac{24L_{V}^{2}L_{D}^{2}\alpha_{k}^{2}}{(1-\gamma)\beta_{k}}\mathbb{E}[\varepsilon_{l,k}^{V}]+12B_{G}^{2}\beta_{k}^{2}$
$\hskip 20.0pt+(1-\frac{(1-\gamma)\beta_{k}}{4})\mathbb{E}[\varepsilon_{l,k}^{V}]+\frac{6L_{V}^{2}\alpha_{k}^{2}}{(1-\gamma)\beta_{k}}\mathbb{E}[\|\nabla_{\theta}J_{f}(\pi_{\theta_{k}},\phi_{\omega_{k}},\hat{\mu}_{k})\|^{2}]+\frac{6L_{V}^{2}L_{H}^{2}\xi_{k}^{2}}{(1-\gamma)\beta_{k}}\mathbb{E}[\varepsilon_{k}^{\pi}]$
$\hskip 20.0pt+\frac{6L_{V}^{2}\eta_{2}^{2}\zeta_{k}^{2}}{(1-\gamma)\beta_{k}}\mathbb{E}[\varepsilon_{k}^{\phi}]+\frac{16L^{2}L_{V}^{2}L_{H}^{2}\xi_{k}^{2}}{(1-\gamma)\beta_{k}}\mathbb{E}[\varepsilon_{k}^{\mu}]+\frac{6L_{V}^{4}\zeta_{k}^{2}}{(1-\gamma)\beta_{k}}\mathbb{E}[\varepsilon_{f,k}^{V}]+12B_{G}^{2}\beta_{k}^{2}$
$\leq\underbrace{\Big(-\frac{\zeta_{k}}{2\eta_{1}}+\frac{8L^{2}\eta_{2}^{2}\zeta_{k}^{2}}{(1-\delta)\xi_{k}}+\frac{2|{\cal S}|L_{V}^{2}\eta_{2}^{2}\zeta_{k}^{2}}{(1-\gamma)\tau^{2}\rho_{\min}^{3}p_{\min}^{2}\alpha_{k}}+\frac{12L_{V}^{2}\eta_{2}^{2}\zeta_{k}^{2}}{(1-\gamma)\beta_{k}}\Big)}_{T_{1}}\mathbb{E}[\varepsilon_{k}^{\phi}]$
$\hskip 20.0pt+\mathbb{E}[\varepsilon_{k}^{\mu}]+\underbrace{\Big(-\frac{1-\delta}{4}\xi_{k}+2\eta_{1}L_{D}^{2}L_{V}^{2}\zeta_{k}+\frac{16|{\cal S}|L^{2}L_{V}^{4}L_{D}^{2}L_{H}^{2}\xi_{k}^{2}}{(1-\gamma)\tau^{2}\rho_{\min}^{3}p_{\min}^{2}\alpha_{k}}+\frac{32L^{2}L_{V}^{2}L_{H}^{2}\xi_{k}^{2}}{(1-\gamma)\beta_{k}}\Big)}_{T_{2}}\mathbb{E}[\varepsilon_{k}^{\mu}]$
$\hskip 20.0pt+\mathbb{E}[\varepsilon_{k}^{\pi}]+\underbrace{\Big(-\frac{\alpha_{k}}{16}+\frac{12L_{V}^{2}\alpha_{k}^{2}}{(1-\gamma)\beta_{k}}\Big)\mathbb{E}[\|\nabla_{\theta}J_{f}(\pi_{\theta_{k}},\phi_{\omega_{k}},\hat{\mu}_{k})\|^{2}]+\Big(\frac{11L_{H}^{2}\xi_{k}}{(1-\delta)\tau\rho_{\min}}+\frac{12L_{V}^{2}L_{H}^{2}\xi_{k}^{2}}{(1-\gamma)\beta_{k}}\Big)\mathbb{E}[\varepsilon_{k}^{\pi}]}_{T_{3}}$
$\hskip 20.0pt+\mathbb{E}[\varepsilon_{l,k}^{V}]+\underbrace{\Big(-\frac{1-\gamma}{4}\beta_{k}+2\eta_{1}L_{D}^{2}\zeta_{k}+\frac{32L^{2}L_{D}^{2}\zeta_{k}^{2}}{(1-\delta)\xi_{k}}+\frac{8|{\cal S}|L_{V}^{2}\zeta_{k}^{2}}{(1-\gamma)\tau^{2}\rho_{\min}^{3}p_{\min}^{2}\alpha_{k}}+\frac{24L_{V}^{2}L_{D}^{2}\alpha_{k}^{2}}{(1-\gamma)\beta_{k}}\Big)}_{T_{4}}\mathbb{E}[\varepsilon_{l,k}^{V}]$
$\hskip 20.0pt+\mathbb{E}[\varepsilon_{f,k}^{V}]+\underbrace{\Big(-\frac{1-\gamma}{4}\beta_{k}+2L_{V}^{2}\alpha_{k}+\frac{6L_{V}^{4}\zeta_{k}^{2}}{(1-\gamma)\beta_{k}}\Big)}_{T_{5}}\mathbb{E}[\varepsilon_{f,k}^{V}]$
$\hskip 20.0pt+\underbrace{\frac{B_{D}^{2}L_{V}\zeta_{k}^{2}}{2}+16B_{H}^{2}\xi_{k}^{2}+3B_{F}(B_{D}+B_{H})\alpha_{k}\xi_{k}+24B_{G}^{2}\beta_{k}^{2}}_{T_{6}}.$ (62)
We bound each term of (62) individually. First, we can ensure $T_{1}\leq-\frac{\zeta_{k}}{8\eta_{1}}$ by choosing the step sizes such that every positive term is no larger than $\frac{\zeta_{k}}{8\eta_{1}}$. The required step size conditions are
$\frac{\zeta_{k}}{\xi_{k}}\leq\frac{1-\delta}{64L^{2}\eta_{1}\eta_{2}^{2}},\quad\frac{\zeta_{k}}{\alpha_{k}}\leq\frac{(1-\gamma)\tau^{2}\rho_{\min}^{3}p_{\min}^{2}}{16|{\cal S}|L_{V}^{2}\eta_{1}\eta_{2}^{2}},\quad\frac{\zeta_{k}}{\beta_{k}}\leq\frac{1-\gamma}{96L_{V}^{2}\eta_{1}\eta_{2}^{2}}.$
For $T_{2}$, we choose the step sizes such that each positive term is no larger than $\frac{1-\delta}{12}$, making $T_{2}$ non-positive. The step size conditions are
$\frac{\zeta_{k}}{\xi_{k}}\leq\frac{1-\delta}{24\eta_{1}L_{D}^{2}L_{V}^{2}},\quad\frac{\zeta_{k}}{\alpha_{k}}\leq\frac{(1-\delta)(1-\gamma)\tau^{2}\rho_{\min}^{3}p_{\min}^{2}}{192|{\cal S}|L^{2}L_{V}^{4}L_{D}^{2}L_{H}^{2}},\quad\frac{\zeta_{k}}{\beta_{k}}\leq\frac{(1-\delta)(1-\gamma)}{384L^{2}L_{V}^{2}L_{H}^{2}}.$
To bound $T_{3}$, note that under entropy regularization, $J_{f}$ satisfies a non-uniform PL condition with respect to the follower’s policy [Mei et al., 2020]. Specifically, for any $\theta,\phi,\mu$,
$\begin{gathered}\|\nabla_{\theta}J_{f}(\pi_{\theta},\phi,\mu)\|^{2}\geq\frac{2(1-\gamma)\tau\rho_{\min}^{2}p_{\min}^{2}}{|{\cal S}|}\Big(J_{f}(\pi^{\star}(\phi,\mu),\phi,\mu)-J_{f}(\pi_{\theta},\phi,\mu)\Big).\end{gathered}$ (64)
The inequality in this specific form is adapted from Zeng et al. [2022][Lemma 4].
Under the step size conditions $\frac{\alpha_{k}}{\beta_{k}}\leq\frac{1-\gamma}{384L_{V}^{2}}$, $\xi_{k}\leq\beta_{k}$, and $\frac{\xi_{k}}{\alpha_{k}}\leq\min\{\frac{(1-\gamma)(1-\delta)\tau\rho_{\min}^{3}p_{\min}^{2}}{352|{\cal S}|L_{H}^{2}},\frac{(1-\gamma)^{2}\tau\rho_{\min}^{2}p_{\min}^{2}}{384|{\cal S}|L_{V}^{2}L_{H}^{2}}\}$, (64) implies
$T_{3}$ $\leq-\frac{\alpha_{k}}{32}\mathbb{E}[\|\nabla_{\theta}J_{f}(\pi_{\theta_{k}},\phi_{\omega_{k}},\hat{\mu}_{k})\|^{2}]+\Big(\frac{11L_{H}^{2}\xi_{k}}{(1-\delta)\tau\rho_{\min}}+\frac{12L_{V}^{2}L_{H}^{2}\xi_{k}^{2}}{(1-\gamma)\beta_{k}}\Big)\mathbb{E}[\varepsilon_{k}^{\pi}]$
$\leq\Big(-\frac{(1-\gamma)\tau\rho_{\min}^{2}p_{\min}^{2}}{16|{\cal S}|}\alpha_{k}+\frac{11L_{H}^{2}\xi_{k}}{(1-\delta)\tau\rho_{\min}}+\frac{12L_{V}^{2}L_{H}^{2}\xi_{k}^{2}}{(1-\gamma)\beta_{k}}\Big)\mathbb{E}[\varepsilon_{k}^{\pi}]$
$\leq 0.$
For $T_{4}$, each positive term is no larger than $\frac{1-\gamma}{16}$ under the step size conditions
$\frac{\zeta_{k}}{\beta_{k}}\leq\frac{1-\gamma}{32\eta_{1}L_{D}^{2}},\quad\frac{\zeta_{k}}{\xi_{k}}\leq\frac{\eta_{1}(1-\delta)}{16L^{2}},\quad\frac{\zeta_{k}}{\alpha_{k}}\leq\frac{(1-\gamma)\eta_{1}\tau^{2}\rho_{\min}^{3}p_{\min}^{2}L_{D}^{2}}{4|{\cal S}|L_{V}^{2}},\quad\frac{\alpha_{k}}{\beta_{k}}\leq\frac{1-\gamma}{8\sqrt{6}L_{V}L_{D}}.$
This ensures $T_{4}\leq 0$.
Similarly, we have $T_{5}\leq 0$ under the conditions
$\frac{\alpha_{k}}{\beta_{k}}\leq\frac{1-\gamma}{16L_{V}^{2}},\quad\frac{\zeta_{k}}{\beta_{k}}\leq\frac{1-\gamma}{2\sqrt{6}L_{V}^{2}}.$
Finally, we choose the step sizes such that each term in $T_{6}$ is no larger than $\alpha_{k}^{2}$
$\frac{\zeta_{k}}{\alpha_{k}}\leq B_{D}\sqrt{\frac{L_{V}}{2}},\quad\frac{\xi_{k}}{\alpha_{k}}\leq\min\{\frac{1}{4B_{H}},\frac{1}{B_{F}(B_{D}+B_{H})}\},\quad\frac{\beta_{k}}{\alpha_{k}}\leq\frac{1}{2\sqrt{6}B_{G}}.$
Plugging the bounds on $T_{1}$-$T_{6}$ into (62), we have
$\mathbb{E}\left[\Big(J_{l}(\phi_{\omega_{k}},\mu^{\star}(\phi_{\omega_{k}}))-J_{l}(\phi_{\omega_{k+1}},\mu^{\star}(\phi_{\omega_{k+1}}))\Big)+\varepsilon_{k+1}^{\mu}+\varepsilon_{k+1}^{\pi}+\varepsilon_{f,k+1}^{V}+\varepsilon_{l,k+1}^{V}\right]$
$\leq-\frac{\zeta_{k}}{8\eta_{1}}\mathbb{E}[\varepsilon_{k}^{\phi}]+\mathbb{E}[\varepsilon_{k}^{\mu}+\varepsilon_{k}^{\pi}+\varepsilon_{l,k}^{V}+\varepsilon_{f,k}^{V}]+4\alpha_{k}^{2}.$ (65)
Re-arranging the terms and plugging in the step size, we have
$\frac{c_{\zeta}}{8(k+1)^{1/2}}\mathbb{E}[\varepsilon_{k}^{\phi}]$ $\leq\mathbb{E}[-J_{l}(\phi_{\omega_{k}},\mu^{\star}(\phi_{\omega_{k}}))+\varepsilon_{k}^{\mu}+\varepsilon_{k}^{\pi}+\varepsilon_{l,k}^{V}+\varepsilon_{f,k}^{V}]$
$\hskip 20.0pt-\mathbb{E}[-J_{l}(\phi_{\omega_{k+1}},\mu^{\star}(\phi_{\omega_{k+1}}))+\varepsilon_{k+1}^{\mu}+\varepsilon_{k+1}^{\pi}+\varepsilon_{f,k+1}^{V}+\varepsilon_{l,k+1}^{V}]+\frac{4c_{\alpha}}{(k+1)}.$
Recall that $\phi^{\star}$ is the optimal solution to the Stackelberg objective in (6). The inequality above implies
$\sum_{t=0}^{k-1}\frac{c_{\zeta}}{8(t+1)^{1/2}}\mathbb{E}[\varepsilon_{t}^{\phi}]$
$\leq\sum_{t=0}^{k-1}\frac{c_{\zeta}}{8(t+1)^{1/2}}\mathbb{E}[\varepsilon_{t}^{\phi}]+\mathbb{E}[\varepsilon_{k+1}^{\mu}+\varepsilon_{k+1}^{\pi}+\varepsilon_{f,k+1}^{V}+\varepsilon_{l,k+1}^{V}]$
$\leq J_{l}(\phi^{\star},\mu^{\star}(\phi^{\star}))-J(\phi_{\omega_{0}},\mu^{\star}(\phi_{\omega_{0}}))+\varepsilon_{0}^{\mu}+\varepsilon_{0}^{\pi}+\varepsilon_{l,0}^{V}+\varepsilon_{f,0}^{V}+\sum_{t=0}^{k-1}\frac{4c_{\alpha}}{(t+1)}.$ (66)
The following relationships are standard results: for any $k\geq 0$, we have
$\sum_{t=0}^{k-1}\frac{1}{(t+1)^{1/2}}\geq\frac{(k+1)^{1/2}}{2},\quad\sum_{t=0}^{k-1}\frac{1}{t+1}\leq 2\log(k+1).$ (67)
Combining (66) and (67),
$\min_{t<k}\mathbb{E}[\varepsilon_{t}^{\phi}]$ $\leq\frac{\sum_{t=0}^{k-1}\frac{c_{\zeta}}{8(t+1)^{1/2}}\mathbb{E}[\varepsilon_{t}^{\phi}]}{\sum_{t^{\prime}=0}^{k-1}\frac{c_{\zeta}}{8(t^{\prime}+1)^{1/2}}}$
$\leq\frac{16}{c_{\zeta}(k+1)^{1/2}}\Big(J_{l}(\phi^{\star},\mu^{\star}(\phi^{\star}))-J(\phi_{\omega_{0}},\mu^{\star}(\phi_{\omega_{0}}))+\varepsilon_{0}^{\mu}+\varepsilon_{0}^{\pi}+\varepsilon_{l,0}^{V}+\varepsilon_{f,0}^{V}\Big)$
$\hskip 250.0pt+\frac{128c_{\alpha}\log(k+1)}{c_{\zeta}(k+1)^{1/2}}.$
$\blacksquare$
Appendix C Proof of Propositions
C.1 Proof of Proposition 1
By the $L_{V}$-smoothness of the function $J_{l}$
$J_{l}(\phi_{\omega_{k}},\mu^{\star}(\phi_{\omega_{k}}))-J_{l}(\phi_{\omega_{k+1}},\mu^{\star}(\phi_{\omega_{k+1}}))$
$\leq-\langle\nabla_{\omega}J_{l}(\phi_{\omega_{k}},\mu^{\star}(\phi_{\omega_{k}}),\omega_{k+1}-\omega_{k}\rangle+\frac{L_{V}}{2}\|\omega_{k+1}-\omega_{k}\|^{2}$
$=-\zeta_{k}\langle\nabla_{\omega}J_{l}(\phi_{\omega_{k}},\mu^{\star}(\phi_{\omega_{k}}),D(\omega_{k},\hat{\mu}_{k},\hat{V}_{l,k},s_{k},b_{k},s_{k}^{\prime})\rangle+\frac{L_{V}\zeta_{k}^{2}}{2}\|D(\omega_{k},\hat{\mu}_{k},\hat{V}_{l,k},s_{k},b_{k},s_{k}^{\prime})\|^{2}$
$=-\zeta_{k}\langle\nabla_{\omega}J_{l}(\phi_{\omega_{k}},\mu^{\star}(\phi_{\omega_{k}}),D(\omega_{k},\hat{\mu}_{k},\hat{V}_{l,k},s_{k},b_{k},s_{k}^{\prime})-\bar{D}(\omega_{k},\hat{\mu}_{k},\hat{V}_{l,k})\rangle$
$\hskip 20.0pt-\zeta_{k}\langle\nabla_{\omega}J_{l}(\phi_{\omega_{k}},\mu^{\star}(\phi_{\omega_{k}}),\bar{D}(\omega_{k},\hat{\mu}_{k},\hat{V}_{l,k})\rangle+\frac{L_{V}\zeta_{k}^{2}}{2}\|D(\omega_{k},\hat{\mu}_{k},\hat{V}_{l,k},s_{k},b_{k},s_{k}^{\prime})\|^{2}.$ (68)
Taking the expectation and plugging in the bound on $D$ from Lemma 3, we have
$\mathbb{E}[J_{l}(\phi_{\omega_{k}},\mu^{\star}(\phi_{\omega_{k}}))-J_{l}(\phi_{\omega_{k+1}},\mu^{\star}(\phi_{\omega_{k+1}}))]$
$\leq-\zeta_{k}\mathbb{E}[\langle\nabla_{\omega}J_{l}(\phi_{\omega_{k}},\mu^{\star}(\phi_{\omega_{k}}),\bar{D}(\omega_{k},\hat{\mu}_{k},\hat{V}_{l,k})\rangle]+\frac{L_{V}B_{D}^{2}\zeta_{k}^{2}}{2}$
$\leq-\frac{\zeta_{k}}{\eta_{1}}\mathbb{E}[\|\nabla_{\omega}J_{l}(\phi_{\omega_{k}},\mu^{\star}(\phi_{\omega_{k}})\|^{2}]+\frac{L_{V}B_{D}^{2}\zeta_{k}^{2}}{2}$
$\hskip 20.0pt+\zeta_{k}\mathbb{E}[\langle\nabla_{\omega}J_{l}(\phi_{\omega_{k}},\mu^{\star}(\phi_{\omega_{k}}),\bar{D}\left(\omega_{k},\mu^{\star}(\phi_{\omega_{k}}),V_{l}^{\phi_{\omega_{k}},\,\mu^{\star}(\phi_{\omega_{k}})}\right)-\bar{D}(\omega_{k},\hat{\mu}_{k},\hat{V}_{l,k})\rangle],$ (69)
where the first inequality follows from
$\mathbb{E}[\langle\nabla_{\omega}J_{l}(\phi_{\omega_{k}},\mu^{\star}(\phi_{\omega_{k}}),D(\omega_{k},\hat{\mu}_{k},\hat{V}_{l,k},s_{k},b_{k},s_{k}^{\prime})-\bar{D}(\omega_{k},\hat{\mu}_{k},\hat{V}_{l,k})\rangle]$
$=\mathbb{E}[\langle\nabla_{\omega}J_{l}(\phi_{\omega_{k}},\mu^{\star}(\phi_{\omega_{k}}),\mathbb{E}[D(\omega_{k},\hat{\mu}_{k},\hat{V}_{l,k},s_{k},b_{k},s_{k}^{\prime})-\bar{D}(\omega_{k},\hat{\mu}_{k},\hat{V}_{l,k})]\mid{\cal F}_{k-1}\rangle]=0.$
and the second inequality follows from Assumption  6 and the relationship
$\nabla_{\omega}J(\phi_{\omega},\mu)\mid_{\mu=\mu^{\star}(\phi_{\omega})}\,=\bar{D}(\omega,\mu^{\star}(\phi_{\omega}),V_{l}^{\phi_{\omega},\,\mu^{\star}(\phi_{\omega})}).$
To bound the last term on the right hand side of (69), we apply Young’s inequality
$\zeta_{k}\langle\nabla_{\omega}J_{l}(\phi_{\omega_{k}},\mu^{\star}(\phi_{\omega_{k}}),\bar{D}\left(\omega_{k},\mu^{\star}(\phi_{\omega_{k}}),V_{l}^{\phi_{\omega_{k}},\,\mu^{\star}(\phi_{\omega_{k}})}\right)-\bar{D}(\omega_{k},\hat{\mu}_{k},\hat{V}_{l,k})\rangle$
$\leq\frac{\zeta_{k}}{2\eta_{1}}\|\nabla_{\omega}J(\phi_{\omega_{k}},\mu^{\star}(\phi_{\omega_{k}}))\|^{2}+\frac{\eta_{1}\zeta_{k}}{2}\left\|\bar{D}\left(\omega_{k},\mu^{\star}(\phi_{\omega_{k}}),V_{l}^{\phi_{\omega_{k}},\,\mu^{\star}(\phi_{\omega_{k}})}\right)-\bar{D}(\omega_{k},\hat{\mu}_{k},\hat{V}_{l,k})\right\|^{2}$
$\leq\frac{\zeta_{k}}{2\eta_{1}}\|\nabla_{\omega}J(\phi_{\omega_{k}},\mu^{\star}(\phi_{\omega_{k}}))\|^{2}+\eta_{1}L_{D}^{2}\zeta_{k}\|\hat{V}_{l,k}-V_{l}^{\phi_{\omega_{k}},\,\mu^{\star}(\phi_{\omega_{k}})}\|^{2}+\eta_{1}L_{D}^{2}\zeta_{k}\|\hat{\mu}_{k}-\mu^{\star}(\phi_{\omega_{k}})\|^{2}$
$\leq\frac{\zeta_{k}}{2\eta_{1}}\varepsilon_{k}^{\phi}+2\eta_{1}L_{D}^{2}\zeta_{k}\|\hat{V}_{l,k}-V_{l}^{\phi_{\omega_{k}},\,\hat{\mu}_{k}}\|^{2}+\eta_{1}L_{D}^{2}(L_{V}^{2}+1)\zeta_{k}\|\hat{\mu}_{k}-\mu^{\star}(\phi_{\omega_{k}})\|^{2}$
$\leq\frac{\zeta_{k}}{2\eta_{1}}\varepsilon_{k}^{\phi}+2\eta_{1}L_{D}^{2}L_{V}^{2}\zeta_{k}\varepsilon_{k}^{\mu}+2\eta_{1}L_{D}^{2}\zeta_{k}\varepsilon_{l,k}^{V},$ (70)
where the second and third inequalities follow from the Lipschitz continuity established in Lemma 4.
Substituting (70) into (69),
$\mathbb{E}[J_{l}(\phi_{\omega_{k}},\mu^{\star}(\phi_{\omega_{k}}))-J_{l}(\phi_{\omega_{k+1}},\mu^{\star}(\phi_{\omega_{k+1}}))]$
$\leq-\frac{\zeta_{k}}{\eta_{1}}\mathbb{E}[\varepsilon_{k}^{\phi}]+\frac{L_{V}B_{D}^{2}\zeta_{k}^{2}}{2}+\zeta_{k}\mathbb{E}[\langle\nabla_{\omega}J_{l}(\phi_{\omega_{k}},\mu^{\star}(\phi_{\omega_{k}}),\bar{D}\left(\omega_{k},\mu^{\star}(\phi_{\omega_{k}}),V_{l}^{\phi_{\omega_{k}},\,\mu^{\star}(\phi_{\omega_{k}})}\right)-\bar{D}(\omega_{k},\hat{\mu}_{k},\hat{V}_{l,k})\rangle]$
$\leq-\frac{\zeta_{k}}{2\eta_{1}}\mathbb{E}[\varepsilon_{k}^{\phi}]+2\eta_{1}L_{D}^{2}L_{V}^{2}\zeta_{k}\mathbb{E}[\varepsilon_{k}^{\mu}]+2\eta_{1}L_{D}^{2}\zeta_{k}\mathbb{E}[\varepsilon_{l,k}^{V}]+\frac{L_{V}B_{D}^{2}\zeta_{k}^{2}}{2}.$
$\blacksquare$
C.2 Proof of Proposition 2
The analysis employs the following lemma, which bounds an intermediate cross term. We provide the proof of the lemma in Section D.5.
Lemma 5
Under the assumptions and step sizes of Proposition 2, we have for all $k\geq 0$,
$\mathbb{E}[\langle\hat{\mu}_{k}-\mu^{\star}(\phi_{\omega_{k}})+\xi_{k}\bar{H}(\pi^{\star}(\phi_{\omega_{k}},\hat{\mu}_{k}),\omega_{k},\hat{\mu}_{k}),\mu^{\star}(\phi_{\omega_{k+1}})-\mu^{\star}(\phi_{\omega_{k}})\rangle]$
$\leq\frac{(1-\delta)\xi_{k}}{8}\mathbb{E}[\|\hat{\mu}_{k}-\mu^{\star}(\phi_{\omega_{k}})+\xi_{k}\bar{H}(\pi^{\star}(\phi_{\omega_{k}}\|^{2}]+\frac{48L^{2}L_{V}^{2}L_{D}^{2}\zeta_{k}^{2}}{(1-\delta)\xi_{k}}\mathbb{E}[\varepsilon_{k}^{\mu}]$
$\hskip 20.0pt+\frac{8L^{2}\eta_{2}^{2}\zeta_{k}^{2}}{(1-\delta)\xi_{k}}\mathbb{E}[\varepsilon_{k}^{\phi}]+\frac{32L^{2}L_{D}^{2}\zeta_{k}^{2}}{(1-\delta)\xi_{k}}\mathbb{E}[\varepsilon_{l,k}^{V}]+B_{D}^{2}L\zeta_{k}^{2}.$
We define the shorthand notation $\Delta h_{k}\triangleq H(\hat{\mu}_{k},\bar{s}_{k})-\bar{H}(\pi_{\theta_{k}},\omega_{k},\hat{\mu}_{k})$. Note that $\mathbb{E}[\Delta h_{k}]=0$ and $\|\Delta h_{k}\|^{2}\leq 4B_{H}^{2}$.
By the definition of $\varepsilon_{k}^{\mu}$,
$\varepsilon_{k+1}^{\mu}$ $=\|\hat{\mu}_{k+1}-\mu^{\star}(\phi_{\omega_{k+1}})\|^{2}$
$=\|\Pi_{\Delta_{{\cal S}}}(\hat{\mu}_{k}+\xi_{k}H(\hat{\mu}_{k},\bar{s}_{k}))-\mu^{\star}(\phi_{\omega_{k+1}})\|^{2}$
$\leq\|\hat{\mu}_{k}+\xi_{k}H(\hat{\mu}_{k},\bar{s}_{k})-\mu^{\star}(\phi_{\omega_{k+1}})\|^{2}$
$=\Big\|\hat{\mu}_{k}-\mu^{\star}(\phi_{\omega_{k}})+\xi_{k}\Delta h_{k}+\xi_{k}\bar{H}(\pi^{\star}(\phi_{\omega_{k}},\hat{\mu}_{k}),\omega_{k},\hat{\mu}_{k})$
$\hskip 20.0pt+\xi_{k}\left(\bar{H}(\pi_{\theta_{k}},\omega_{k},\hat{\mu}_{k})-\bar{H}(\pi^{\star}(\phi_{\omega_{k}},\hat{\mu}_{k}),\omega_{k},\hat{\mu}_{k})\right)-\left(\mu^{\star}(\phi_{\omega_{k+1}})-\mu^{\star}(\phi_{\omega_{k}})\right)\Big\|^{2}$
$\leq\|\hat{\mu}_{k}-\mu^{\star}(\phi_{\omega_{k}})+\xi_{k}\bar{H}(\pi^{\star}(\phi_{\omega_{k}},\hat{\mu}_{k}),\omega_{k},\hat{\mu}_{k})\|^{2}+3\xi_{k}^{2}\|\Delta h_{k}\|^{2}+3\|\mu^{\star}(\phi_{\omega_{k+1}})-\mu^{\star}(\phi_{\omega_{k}})\|^{2}$
$\hskip 20.0pt+3\xi_{k}^{2}\|\bar{H}(\pi_{\theta_{k}},\omega_{k},\hat{\mu}_{k})-\bar{H}(\pi^{\star}(\phi_{\omega_{k}},\hat{\mu}_{k}),\omega_{k},\hat{\mu}_{k})\|^{2}$
$\hskip 20.0pt+2\xi_{k}\langle\hat{\mu}_{k}-\mu^{\star}(\phi_{\omega_{k}})+\xi_{k}\bar{H}(\pi^{\star}(\phi_{\omega_{k}},\hat{\mu}_{k}),\omega_{k},\hat{\mu}_{k}),\Delta h_{k}\rangle$
$\hskip 20.0pt+2\langle\hat{\mu}_{k}-\mu^{\star}(\phi_{\omega_{k}})+\xi_{k}\bar{H}(\pi^{\star}(\phi_{\omega_{k}},\hat{\mu}_{k}),\omega_{k},\hat{\mu}_{k}),\mu^{\star}(\phi_{\omega_{k+1}})-\mu^{\star}(\phi_{\omega_{k}})\rangle$
$\hskip 20.0pt+2\xi_{k}\langle\hat{\mu}_{k}-\mu^{\star}(\phi_{\omega_{k}})+\xi_{k}\bar{H}(\pi^{\star}(\phi_{\omega_{k}},\hat{\mu}_{k}),\omega_{k},\hat{\mu}_{k}),\bar{H}(\pi^{\star}(\phi_{\omega_{k}},\hat{\mu}_{k}),\omega_{k},\hat{\mu}_{k})-\bar{H}(\pi_{\theta_{k}},\omega_{k},\hat{\mu}_{k})\rangle,$ (71)
where the first inequality is due to the fact that projection to a convex set is a non-expansive operator.
As $\bar{H}(\pi^{\star}(\phi),\phi,\mu^{\star}(\phi))=0$ for any $\phi$, we have for the first term of (71),
$\|\hat{\mu}_{k}-\mu^{\star}(\phi_{\omega_{k}})+\xi_{k}\bar{H}(\pi^{\star}(\phi_{\omega_{k}},\hat{\mu}_{k}),\omega_{k},\hat{\mu}_{k})\|^{2}$
$=\|\hat{\mu}_{k}-\mu^{\star}(\phi_{\omega_{k}})+\xi_{k}\bar{H}(\pi^{\star}(\phi_{\omega_{k}},\hat{\mu}_{k}),\omega_{k},\hat{\mu}_{k})-\xi_{k}\bar{H}(\pi^{\star}(\phi_{\omega_{k}}),\omega_{k},\mu^{\star}(\phi_{\omega_{k}}))\|^{2}$
$=\|\hat{\mu}_{k}-\mu^{\star}(\phi_{\omega_{k}})\|^{2}+\xi_{k}^{2}\|\bar{H}(\pi^{\star}(\phi_{\omega_{k}},\hat{\mu}_{k}),\omega_{k},\hat{\mu}_{k})-\bar{H}(\pi^{\star}(\phi_{\omega_{k}}),\omega_{k},\mu^{\star}(\phi_{\omega_{k}}))\|^{2}$
$\hskip 20.0pt+\xi_{k}\langle\hat{\mu}_{k}-\mu^{\star}(\phi_{\omega_{k}}),\nu^{\pi^{\star}(\phi_{\omega_{k}},\hat{\mu}_{k}),\omega_{k},\hat{\mu}_{k}}-\hat{\mu}_{k}-\nu^{\pi^{\star}(\phi_{\omega_{k}}),\omega_{k},\mu^{\star}(\phi_{\omega_{k}})}+\mu^{\star}(\phi_{\omega_{k}})\rangle$
$\leq(1-\xi_{k})\|\hat{\mu}_{k}-\mu^{\star}(\phi_{\omega_{k}})\|^{2}+2L_{H}^{2}\xi_{k}^{2}\|\hat{\mu}_{k}-\mu^{\star}(\phi_{\omega_{k}})\|^{2}+2L_{H}^{2}\xi_{k}^{2}\|\pi^{\star}(\phi_{\omega_{k}},\hat{\mu}_{k})-\pi^{\star}(\phi_{\omega_{k}})\|^{2}$
$\hskip 20.0pt+\xi_{k}\|\hat{\mu}_{k}-\mu^{\star}(\phi_{\omega_{k}})\|\|\nu^{\pi^{\star}(\phi_{\omega_{k}},\hat{\mu}_{k}),\omega_{k},\hat{\mu}_{k}}-\nu^{\pi^{\star}(\phi_{\omega_{k}}),\omega_{k},\mu^{\star}(\phi_{\omega_{k}})}\|$
$\leq\Big(1-(1-\delta)\xi_{k}+4L^{2}L_{H}^{2}\xi_{k}^{2}\Big)\|\hat{\mu}_{k}-\mu^{\star}(\phi_{\omega_{k}})\|^{2}$
$\leq(1-\frac{1-\delta}{2}\xi_{k})\varepsilon_{k}^{\mu},$ (72)
where the second equation plugs in the definition of $\bar{H}$ in (34), and the first inequality is due to the Lipschitz continuity of $\bar{H}$, the second inequality follows from Assumption 4 and the Lipschitz continuity of $\pi^{\star}$ from Assumption 2, and the final inequality is due to the step size condition $\xi_{k}\leq\frac{1-\delta}{8L^{2}L_{H}^{2}}$.
For the fourth term of (71),
$2\xi_{k}\mathbb{E}[\langle\hat{\mu}_{k}-\mu^{\star}(\phi_{\omega_{k}})+\xi_{k}\bar{H}(\pi^{\star}(\phi_{\omega_{k}},\hat{\mu}_{k}),\omega_{k},\hat{\mu}_{k}),\Delta h_{k}\rangle]$
$\leq 2\xi_{k}\mathbb{E}[\langle\hat{\mu}_{k}-\mu^{\star}(\phi_{\omega_{k}})+\xi_{k}\bar{H}(\pi^{\star}(\phi_{\omega_{k}},\hat{\mu}_{k}),\omega_{k},\hat{\mu}_{k}),\mathbb{E}[\Delta h_{k}\mid{\cal F}_{k-1}]\rangle]$
$=0.$ (73)
For the fifth term of (71), we have from Lemma 5
$2\langle\hat{\mu}_{k}-\mu^{\star}(\phi_{\omega_{k}})+\xi_{k}\bar{H}(\pi^{\star}(\phi_{\omega_{k}},\hat{\mu}_{k}),\omega_{k},\hat{\mu}_{k}),\mu^{\star}(\phi_{\omega_{k+1}})-\mu^{\star}(\phi_{\omega_{k}})\rangle$
$\leq\frac{(1-\delta)\xi_{k}}{8}\mathbb{E}[\|\hat{\mu}_{k}-\mu^{\star}(\phi_{\omega_{k}})+\xi_{k}\bar{H}(\pi^{\star}(\phi_{\omega_{k}},\hat{\mu}_{k}),\omega_{k},\hat{\mu}_{k})\|^{2}]+\frac{48L^{2}L_{V}^{2}L_{D}^{2}\zeta_{k}^{2}}{(1-\delta)\xi_{k}}\mathbb{E}[\varepsilon_{k}^{\mu}]$
$\hskip 20.0pt+\frac{8L^{2}\eta_{2}^{2}\zeta_{k}^{2}}{(1-\delta)\xi_{k}}\mathbb{E}[\varepsilon_{k}^{\phi}]+\frac{32L^{2}L_{D}^{2}\zeta_{k}^{2}}{(1-\delta)\xi_{k}}\mathbb{E}[\varepsilon_{l,k}^{V}]+B_{D}^{2}L\zeta_{k}^{2}$
$\leq\frac{1-\delta}{8}\xi_{k}\mathbb{E}[\varepsilon_{k}^{\mu}]+\frac{48L^{2}L_{V}^{2}L_{D}^{2}\zeta_{k}^{2}}{(1-\delta)\xi_{k}}\mathbb{E}[\varepsilon_{k}^{\mu}]+\frac{8L^{2}\eta_{2}^{2}\zeta_{k}^{2}}{(1-\delta)\xi_{k}}\mathbb{E}[\varepsilon_{k}^{\phi}]+\frac{32L^{2}L_{D}^{2}\zeta_{k}^{2}}{(1-\delta)\xi_{k}}\mathbb{E}[\varepsilon_{l,k}^{V}]+B_{D}^{2}L\zeta_{k}^{2}$
$\leq\frac{1-\delta}{4}\xi_{k}\mathbb{E}[\varepsilon_{k}^{\mu}]+\frac{8L^{2}\eta_{2}^{2}\zeta_{k}^{2}}{(1-\delta)\xi_{k}}\mathbb{E}[\varepsilon_{k}^{\phi}]+\frac{32L^{2}L_{D}^{2}\zeta_{k}^{2}}{(1-\delta)\xi_{k}}\mathbb{E}[\varepsilon_{l,k}^{V}]+B_{D}^{2}L\zeta_{k}^{2},$ (74)
where the second inequality plugs in (72), and the last inequality is due to the step size condition $\frac{\zeta_{k}}{\xi_{k}}\leq\frac{1-\delta}{8\sqrt{6}LL_{V}L_{D}}$.
Similarly, for the final term of (71),
$2\xi_{k}\langle\hat{\mu}_{k}-\mu^{\star}(\phi_{\omega_{k}})+\xi_{k}\bar{H}(\pi^{\star}(\phi_{\omega_{k}},\hat{\mu}_{k}),\omega_{k},\hat{\mu}_{k}),\bar{H}(\pi^{\star}(\phi_{\omega_{k}},\hat{\mu}_{k}),\omega_{k},\hat{\mu}_{k})-\bar{H}(\pi_{\theta_{k}},\omega_{k},\hat{\mu}_{k})\rangle$
$\leq\frac{1-\delta}{8}\xi_{k}\|\hat{\mu}_{k}-\mu^{\star}(\phi_{\omega_{k}})+\xi_{k}\bar{H}(\pi^{\star}(\phi_{\omega_{k}},\hat{\mu}_{k}),\omega_{k},\hat{\mu}_{k})\|^{2}$
$\hskip 20.0pt+\frac{8}{1-\delta}\xi_{k}\|\bar{H}(\pi^{\star}(\phi_{\omega_{k}},\hat{\mu}_{k}),\omega_{k},\hat{\mu}_{k})-\bar{H}(\pi_{\theta_{k}},\omega_{k},\hat{\mu}_{k})\|^{2}$
$\leq\frac{1-\delta}{8}\xi_{k}\varepsilon_{k}^{\mu}+\frac{8}{1-\delta}\xi_{k}\|\bar{H}(\pi^{\star}(\phi_{\omega_{k}},\hat{\mu}_{k}),\omega_{k},\hat{\mu}_{k})-\bar{H}(\pi_{\theta_{k}},\omega_{k},\hat{\mu}_{k})\|^{2}.$ (75)
Taking expectation and substituting (72)-(75) into (71),
$\mathbb{E}[\varepsilon_{k+1}^{\mu}]$ $\leq(1-\frac{1-\delta}{2}\xi_{k})\mathbb{E}[\varepsilon_{k}^{\mu}]+3\xi_{k}^{2}\cdot 4B_{H}^{2}+3\mathbb{E}[\|\mu^{\star}(\phi_{\omega_{k+1}})-\mu^{\star}(\phi_{\omega_{k}})\|^{2}]$
$\hskip 20.0pt+3\xi_{k}^{2}\mathbb{E}[\|\bar{H}(\pi_{\theta_{k}},\omega_{k},\hat{\mu}_{k})-\bar{H}(\pi^{\star}(\phi_{\omega_{k}},\hat{\mu}_{k}),\omega_{k},\hat{\mu}_{k})\|^{2}]$
$\hskip 20.0pt+\frac{1-\delta}{4}\xi_{k}\mathbb{E}[\varepsilon_{k}^{\mu}]+\frac{8L^{2}\eta_{2}^{2}\zeta_{k}^{2}}{(1-\delta)\xi_{k}}\mathbb{E}[\varepsilon_{k}^{\phi}]+\frac{32L^{2}L_{D}^{2}\zeta_{k}^{2}}{(1-\delta)\xi_{k}}\mathbb{E}[\varepsilon_{l,k}^{V}]+B_{D}^{2}L\zeta_{k}^{2}$
$\hskip 20.0pt+\frac{1-\delta}{8}\xi_{k}\mathbb{E}[\varepsilon_{k}^{\mu}]+\frac{8}{1-\delta}\xi_{k}\mathbb{E}[\|\bar{H}(\pi^{\star}(\phi_{\omega_{k}},\hat{\mu}_{k}),\omega_{k},\hat{\mu}_{k})-\bar{H}(\pi_{\theta_{k}},\omega_{k},\hat{\mu}_{k})\|^{2}]$
$\leq(1-\frac{1-\delta}{8}\xi_{k})\mathbb{E}[\varepsilon_{k}^{\mu}]+12B_{H}^{2}\xi_{k}^{2}+3L^{2}\mathbb{E}[\|\omega_{k+1}-\omega_{k}\|^{2}]$
$\hskip 20.0pt+\frac{8L^{2}\eta_{2}^{2}\zeta_{k}^{2}}{(1-\delta)\xi_{k}}\mathbb{E}[\varepsilon_{k}^{\phi}]+\frac{32L^{2}L_{D}^{2}\zeta_{k}^{2}}{(1-\delta)\xi_{k}}\mathbb{E}[\varepsilon_{l,k}^{V}]+B_{D}^{2}L\zeta_{k}^{2}$
$\hskip 20.0pt+\frac{11\xi_{k}}{1-\delta}\mathbb{E}[\|\bar{H}(\pi_{\theta_{k}},\omega_{k},\hat{\mu}_{k})-\bar{H}(\pi^{\star}(\phi_{\omega_{k}},\hat{\mu}_{k}),\omega_{k},\hat{\mu}_{k})\|^{2}]$
$\leq(1-\frac{1-\delta}{8}\xi_{k})\mathbb{E}[\varepsilon_{k}^{\mu}]+\frac{11L_{H}^{2}\xi_{k}}{1-\delta}\mathbb{E}[\|\pi_{\theta_{k}}-\pi^{\star}(\phi_{\omega_{k}},\hat{\mu}_{k})\|^{2}]+12B_{H}^{2}\xi_{k}^{2}+3B_{D}^{2}L^{2}\zeta_{k}^{2}$
$\hskip 20.0pt+\frac{8L^{2}\eta_{2}^{2}\zeta_{k}^{2}}{(1-\delta)\xi_{k}}\mathbb{E}[\varepsilon_{k}^{\phi}]+\frac{32L^{2}L_{D}^{2}\zeta_{k}^{2}}{(1-\delta)\xi_{k}}\mathbb{E}[\varepsilon_{l,k}^{V}]+B_{D}^{2}L\zeta_{k}^{2}$
$\leq(1-\frac{1-\delta}{8}\xi_{k})\mathbb{E}[\varepsilon_{k}^{\mu}]+\frac{11L_{H}^{2}\xi_{k}}{1-\delta}\mathbb{E}[\|\pi_{\theta_{k}}-\pi^{\star}(\phi_{\omega_{k}},\hat{\mu}_{k})\|^{2}]$
$\hskip 20.0pt+\frac{8L^{2}\eta_{2}^{2}\zeta_{k}^{2}}{(1-\delta)\xi_{k}}\mathbb{E}[\varepsilon_{k}^{\phi}]+\frac{32L^{2}L_{D}^{2}\zeta_{k}^{2}}{(1-\delta)\xi_{k}}\mathbb{E}[\varepsilon_{l,k}^{V}]+16B_{H}^{2}\xi_{k}^{2},$ (76)
where the second inequality follows from the step size condition $\xi_{k}\leq\frac{1}{1-\delta}$, and the third inequality is a result of the Lipschitz continuity of $\mu^{\star}$ and $\bar{H}$, and the last inequality applies the step size condition $\frac{\zeta_{k}}{\xi_{k}}\leq\frac{B_{H}}{B_{D}L}$.
Note that the following inequality holds for any $\pi,\phi,\mu$ as a result of the entropy regularization (adapted from Lemma 1 of Zeng et al. [2022])
$\|\pi-\pi^{\star}(\phi,\mu)\|^{2}\leq\frac{1}{\tau\rho_{\min}}\Big(J_{f}(\pi^{\star}(\phi,\mu),\phi,\mu)-J_{f}(\pi,\phi,\mu)\Big).$ (77)
Combining (76) and (77),
$\mathbb{E}[\varepsilon_{k+1}^{\mu}]$ $\leq(1-\frac{1-\delta}{8}\xi_{k})\mathbb{E}[\varepsilon_{k}^{\mu}]+\frac{11L_{H}^{2}\xi_{k}}{(1-\delta)\tau\rho_{\min}}\mathbb{E}\Big[J_{f}(\pi^{\star}(\phi_{\omega_{k}},\hat{\mu}_{k}),\phi_{\omega_{k}},\hat{\mu}_{k})-J_{f}(\pi_{\theta_{k}},\phi_{\omega_{k}},\hat{\mu}_{k})\Big]$
$\hskip 20.0pt+\frac{8L^{2}\eta_{2}^{2}\zeta_{k}^{2}}{(1-\delta)\xi_{k}}\mathbb{E}[\varepsilon_{k}^{\phi}]+\frac{32L^{2}L_{D}^{2}\zeta_{k}^{2}}{(1-\delta)\xi_{k}}\mathbb{E}[\varepsilon_{l,k}^{V}]+16B_{H}^{2}\xi_{k}^{2}$
$=(1-\frac{1-\delta}{4}\xi_{k})\mathbb{E}[\varepsilon_{k}^{\mu}]+\frac{11L_{H}^{2}\xi_{k}}{(1-\delta)\tau\rho_{\min}}\mathbb{E}[\varepsilon_{k}^{\pi}]+\frac{8L^{2}\eta_{2}^{2}\zeta_{k}^{2}}{(1-\delta)\xi_{k}}\mathbb{E}[\varepsilon_{k}^{\phi}]+\frac{32L^{2}L_{D}^{2}\zeta_{k}^{2}}{(1-\delta)\xi_{k}}\mathbb{E}[\varepsilon_{l,k}^{V}]+16B_{H}^{2}\xi_{k}^{2}.$
$\blacksquare$
C.3 Proof of Proposition 3
The proof of the proposition relies on the intermediate result below which bounds an important cross term.
Lemma 6
Under the assumptions and step sizes of Proposition 3, we have for all $k\geq 0$
$-\mathbb{E}[\langle\left[\begin{array}[]{c}\nabla_{\omega}J_{f}(\pi_{\theta_{k}},\phi_{\omega_{k}},\hat{\mu}_{k})-\nabla_{\omega}J_{f}(\pi,\phi_{\omega_{k}},\hat{\mu}_{k})\mid_{\pi=\pi^{\star}(\phi_{\omega_{k}},\hat{\mu}_{k})}\\ \nabla_{\mu}J_{f}(\pi_{\theta_{k}},\phi_{\omega_{k}},\hat{\mu}_{k})-\nabla_{\mu}J_{f}(\pi,\phi_{\omega_{k}},\hat{\mu}_{k})\mid_{\pi=\pi^{\star}(\phi_{\omega_{k}},\hat{\mu}_{k})}\end{array}\right],\left[\begin{array}[]{c}\omega_{k+1}-\omega_{k}\\ \hat{\mu}_{k+1}-\hat{\mu}_{k}\end{array}\right]\rangle]$
$\leq\frac{(1-\gamma)\tau^{2}\rho_{\min}^{3}p_{\min}^{2}\alpha_{k}}{4|{\cal S}|}\mathbb{E}[\|\pi^{\star}(\phi_{\omega_{k}},\hat{\mu}_{k})-\pi_{\theta_{k}}\|^{2}]+\frac{2|{\cal S}|L_{V}^{2}L_{D}^{2}L_{H}^{2}\xi_{k}^{2}}{(1-\gamma)\tau^{2}\rho_{\min}^{3}p_{\min}^{2}\alpha_{k}}\mathbb{E}[\|\pi^{\star}(\phi_{\omega_{k},\hat{\mu}_{k}})-\pi_{\theta_{k}}\|^{2}]$
$\hskip 20.0pt+\frac{2|{\cal S}|L_{V}^{2}\eta_{2}^{2}\zeta_{k}^{2}}{(1-\gamma)\tau^{2}\rho_{\min}^{3}p_{\min}^{2}\alpha_{k}}\mathbb{E}[\varepsilon_{k}^{\phi}]+\frac{16|{\cal S}|L^{2}L_{V}^{4}L_{D}^{2}L_{H}^{2}\xi_{k}^{2}}{(1-\gamma)\tau^{2}\rho_{\min}^{3}p_{\min}^{2}\alpha_{k}}\mathbb{E}[\varepsilon_{k}^{\mu}]+\frac{8|{\cal S}|L_{V}^{2}\zeta_{k}^{2}}{(1-\gamma)\tau^{2}\rho_{\min}^{3}p_{\min}^{2}\alpha_{k}}\mathbb{E}[\varepsilon_{l,k}^{V}].$
We provide the proof of Lemma 6 in Section D.6.
As $J_{f}$ has $L_{V}$-Lipschitz gradients,
$J_{f}(\pi_{\theta_{k}},\phi_{\omega_{k+1}},\hat{\mu}_{k+1})-J_{f}(\pi_{\theta_{k+1}},\phi_{\omega_{k+1}},\hat{\mu}_{k+1})$
$\leq-\langle\nabla_{\theta}J_{f}(\pi_{\theta_{k}},\phi_{\omega_{k+1}},\hat{\mu}_{k+1}),\theta_{k+1}-\theta_{k}\rangle+\frac{L_{V}}{2}\|\theta_{k+1}-\theta_{k}\|^{2}$
$=-\alpha_{k}\langle\nabla_{\theta}J_{f}(\pi_{\theta_{k}},\phi_{\omega_{k+1}},\hat{\mu}_{k+1}),F(\theta_{k},\omega_{k},\hat{\mu}_{k},\hat{V}_{f,k},s_{k},a_{k},b_{k},s_{k}^{\prime})\rangle+\frac{L_{V}\alpha_{k}^{2}}{2}\|F(\theta_{k},\omega_{k},\hat{\mu}_{k},\hat{V}_{f,k},s_{k},a_{k},b_{k},s_{k}^{\prime})\|^{2}$
$=-\alpha_{k}\langle\nabla_{\theta}J_{f}(\pi_{\theta_{k}},\phi_{\omega_{k+1}},\hat{\mu}_{k+1}),F(\theta_{k},\omega_{k},\hat{\mu}_{k},\hat{V}_{f,k},s_{k},a_{k},b_{k},s_{k}^{\prime})-\bar{F}(\theta_{k},\omega_{k},\hat{\mu}_{k},\hat{V}_{f,k})\rangle$
$\hskip 20.0pt-\alpha_{k}\langle\nabla_{\theta}J_{f}(\pi_{\theta_{k}},\phi_{\omega_{k+1}},\hat{\mu}_{k+1}),\bar{F}(\theta_{k},\omega_{k},\hat{\mu}_{k},\hat{V}_{f,k})-\bar{F}(\theta_{k},\omega_{k},\hat{\mu}_{k},V_{f}^{\pi_{\theta_{k}},\phi_{\omega_{k}},\hat{\mu}_{k}})\rangle$
$\hskip 20.0pt-\alpha_{k}\langle\nabla_{\theta}J_{f}(\pi_{\theta_{k}},\phi_{\omega_{k+1}},\hat{\mu}_{k+1}),\nabla_{\theta}J_{f}(\pi_{\theta_{k}},\phi_{\omega_{k}},\hat{\mu}_{k})\rangle+\frac{L_{V}\alpha_{k}^{2}}{2}\|F(\theta_{k},\omega_{k},\hat{\mu}_{k},\hat{V}_{f,k},s_{k},a_{k},b_{k},s_{k}^{\prime})\|^{2},$ (78)
where the final equation follows from $\nabla_{\theta}J_{f}(\pi_{\theta},\phi_{\omega},\mu)=\bar{F}(\theta,\omega,\mu,V_{f}^{\pi_{\theta},\phi_{\omega},\mu})$ for all $\theta,\omega,\mu$.
For the first term of (78),
$-\alpha_{k}\mathbb{E}[\langle\nabla_{\theta}J_{f}(\pi_{\theta_{k}},\phi_{\omega_{k+1}},\hat{\mu}_{k+1}),F(\theta_{k},\omega_{k},\hat{\mu}_{k},\hat{V}_{f,k},s_{k},a_{k},b_{k},s_{k}^{\prime})-\bar{F}(\theta_{k},\omega_{k},\hat{\mu}_{k},\hat{V}_{f,k})\rangle]$
$=-\alpha_{k}\mathbb{E}[\langle\nabla_{\theta}J_{f}(\pi_{\theta_{k}},\phi_{\omega_{k}},\hat{\mu}_{k}),\mathbb{E}[F(\theta_{k},\omega_{k},\hat{\mu}_{k},\hat{V}_{f,k},s_{k},a_{k},b_{k},s_{k}^{\prime})-\bar{F}(\theta_{k},\omega_{k},\hat{\mu}_{k},\hat{V}_{f,k})\mid{\cal F}_{k-1}]\rangle]$
$\hskip 20.0pt+\alpha_{k}\mathbb{E}[\langle\nabla_{\theta}J_{f}(\pi_{\theta_{k}},\phi_{\omega_{k}},\hat{\mu}_{k})-\nabla_{\theta}J_{f}(\pi_{\theta_{k}},\phi_{\omega_{k+1}},\hat{\mu}_{k+1}),F(\theta_{k},\omega_{k},\hat{\mu}_{k},\hat{V}_{f,k},s_{k},a_{k},b_{k},s_{k}^{\prime})-\bar{F}(\theta_{k},\omega_{k},\hat{\mu}_{k},\hat{V}_{f,k})\rangle]$
$=\alpha_{k}\mathbb{E}[\langle\nabla_{\theta}J_{f}(\pi_{\theta_{k}},\phi_{\omega_{k}},\hat{\mu}_{k})-\nabla_{\theta}J_{f}(\pi_{\theta_{k}},\phi_{\omega_{k+1}},\hat{\mu}_{k+1}),F(\theta_{k},\omega_{k},\hat{\mu}_{k},\hat{V}_{f,k},s_{k},a_{k},b_{k},s_{k}^{\prime})-\bar{F}(\theta_{k},\omega_{k},\hat{\mu}_{k},\hat{V}_{f,k})\rangle]$
$\leq 2B_{F}\alpha_{k}\cdot L_{V}\mathbb{E}[\|\phi_{\omega_{k+1}}-\phi_{\omega_{k}}\|+\|\hat{\mu}_{k+1}-\hat{\mu}_{k}\|]$
$\leq 2B_{F}L_{V}\alpha_{k}(B_{D}\zeta_{k}+B_{H}\xi_{k})$
$\leq 2B_{F}(B_{D}+B_{H})\alpha_{k}\xi_{k},$ (79)
where the last inequality follows from $\zeta_{k}\leq\xi_{k}$.
For the second term of (78),
$-\alpha_{k}\langle\nabla_{\theta}J_{f}(\pi_{\theta_{k}},\phi_{\omega_{k+1}},\hat{\mu}_{k+1}),\bar{F}(\theta_{k},\omega_{k},\hat{\mu}_{k},\hat{V}_{f,k})-\bar{F}(\theta_{k},\omega_{k},\hat{\mu}_{k},V_{f}^{\pi_{\theta_{k}},\phi_{\omega_{k}},\hat{\mu}_{k}})\rangle$
$\leq\frac{\alpha_{k}}{8}\|\nabla_{\theta}J_{f}(\pi_{\theta_{k}},\phi_{\omega_{k+1}},\hat{\mu}_{k+1})\|^{2}+2\alpha_{k}\|\bar{F}(\theta_{k},\omega_{k},\hat{\mu}_{k},\hat{V}_{f,k})-\bar{F}(\theta_{k},\omega_{k},\hat{\mu}_{k},V_{f}^{\pi_{\theta_{k}},\phi_{\omega_{k}},\hat{\mu}_{k}})\|^{2}$
$\leq\frac{\alpha_{k}}{4}\|\nabla_{\theta}J_{f}(\pi_{\theta_{k}},\phi_{\omega_{k}},\hat{\mu}_{k})\|^{2}+\frac{\alpha_{k}}{4}\|\nabla_{\theta}J_{f}(\pi_{\theta_{k}},\phi_{\omega_{k+1}},\hat{\mu}_{k+1})-\nabla_{\theta}J_{f}(\pi_{\theta_{k}},\phi_{\omega_{k}},\hat{\mu}_{k})\|^{2}+2L_{V}^{2}\alpha_{k}\varepsilon_{f,k}^{V}$
$\leq\frac{\alpha_{k}}{4}\|\nabla_{\theta}J_{f}(\pi_{\theta_{k}},\phi_{\omega_{k}},\hat{\mu}_{k})\|^{2}+\frac{L_{V}^{2}\alpha_{k}}{2}\Big(\|\phi_{\omega_{k+1}}-\phi_{\omega_{k}}\|^{2}+\|\hat{\mu}_{k+1}-\hat{\mu}_{k}\|^{2}\Big)+2L_{V}^{2}\alpha_{k}\varepsilon_{f,k}^{V}$
$\leq\frac{\alpha_{k}}{4}\|\nabla_{\theta}J_{f}(\pi_{\theta_{k}},\phi_{\omega_{k}},\hat{\mu}_{k})\|^{2}+\frac{L_{V}^{2}\alpha_{k}}{2}(B_{D}^{2}\zeta_{k}^{2}+B_{H}^{2}\xi_{k}^{2})+2L_{V}^{2}\alpha_{k}\varepsilon_{f,k}^{V}$
$\leq\frac{\alpha_{k}}{4}\|\nabla_{\theta}J_{f}(\pi_{\theta_{k}},\phi_{\omega_{k}},\hat{\mu}_{k})\|^{2}+2L_{V}^{2}\alpha_{k}\varepsilon_{f,k}^{V}+L_{V}^{2}\xi_{k}^{2},$ (80)
where the third inequality is a result of the Lipschitz continuity of the value function, and the last inequality follows from the step size condition $\alpha_{k}\leq\frac{2}{B_{D}^{2}+B_{H}^{2}}$ and $\zeta_{k}\leq\xi_{k}$.
For the third term of (78),
$-\alpha_{k}\langle\nabla_{\theta}J_{f}(\pi_{\theta_{k}},\phi_{\omega_{k+1}},\hat{\mu}_{k+1}),\nabla_{\theta}J_{f}(\pi_{\theta_{k}},\phi_{\omega_{k}},\hat{\mu}_{k})\rangle$
$=-\alpha_{k}\|\nabla_{\theta}J_{f}(\pi_{\theta_{k}},\phi_{\omega_{k}},\hat{\mu}_{k})\|^{2}+\alpha_{k}\langle\nabla_{\theta}J_{f}(\pi_{\theta_{k}},\phi_{\omega_{k}},\hat{\mu}_{k})-\nabla_{\theta}J_{f}(\pi_{\theta_{k}},\phi_{\omega_{k+1}},\hat{\mu}_{k+1}),\nabla_{\theta}J_{f}(\pi_{\theta_{k}},\phi_{\omega_{k}},\hat{\mu}_{k})\rangle$
$\leq-\frac{\alpha_{k}}{2}\|\nabla_{\theta}J_{f}(\pi_{\theta_{k}},\phi_{\omega_{k}},\hat{\mu}_{k})\|^{2}+\frac{\alpha_{k}}{2}\|\nabla_{\theta}J_{f}(\pi_{\theta_{k}},\phi_{\omega_{k}},\hat{\mu}_{k})-\nabla_{\theta}J_{f}(\pi_{\theta_{k}},\phi_{\omega_{k+1}},\hat{\mu}_{k+1})\|^{2}$
$\leq-\frac{\alpha_{k}}{2}\|\nabla_{\theta}J_{f}(\pi_{\theta_{k}},\phi_{\omega_{k}},\hat{\mu}_{k})\|^{2}+L_{V}^{2}\alpha_{k}\Big(\|\phi_{\omega_{k+1}}-\phi_{\omega_{k}}\|^{2}+\|\hat{\mu}_{k+1}-\hat{\mu}_{k}\|^{2}\Big)$
$\leq-\frac{\alpha_{k}}{2}\|\nabla_{\theta}J_{f}(\pi_{\theta_{k}},\phi_{\omega_{k}},\hat{\mu}_{k})\|^{2}+L_{V}^{2}\alpha_{k}(B_{D}^{2}\zeta_{k}^{2}+B_{H}^{2}\xi_{k}^{2})$
$\leq-\frac{\alpha_{k}}{2}\|\nabla_{\theta}J_{f}(\pi_{\theta_{k}},\phi_{\omega_{k}},\hat{\mu}_{k})\|^{2}+2L_{V}^{2}\xi_{k}^{2},$ (81)
where the last inequality again uses the step size condition $\alpha_{k}\leq\frac{2}{B_{D}^{2}+B_{H}^{2}}$ and $\zeta_{k}\leq\xi_{k}$.
Substituting (79)-(81) into (78) and taking the expectation,
$\mathbb{E}[J_{f}(\pi_{\theta_{k}},\phi_{\omega_{k+1}},\hat{\mu}_{k+1})-J_{f}(\pi_{\theta_{k+1}},\phi_{\omega_{k+1}},\hat{\mu}_{k+1})]$
$\leq 2B_{F}(B_{D}+B_{H})\alpha_{k}\xi_{k}+\frac{\alpha_{k}}{4}\mathbb{E}[\|\nabla_{\theta}J_{f}(\pi_{\theta_{k}},\phi_{\omega_{k}},\hat{\mu}_{k})\|^{2}]+2L_{V}^{2}\alpha_{k}\mathbb{E}[\varepsilon_{f,k}^{V}]+L_{V}^{2}\xi_{k}^{2}$
$\hskip 20.0pt-\frac{\alpha_{k}}{2}\mathbb{E}[\|\nabla_{\theta}J_{f}(\pi_{\theta_{k}},\phi_{\omega_{k}},\hat{\mu}_{k})\|^{2}]+2L_{V}^{2}\xi_{k}^{2}$
$\leq-\frac{\alpha_{k}}{4}\mathbb{E}[\|\nabla_{\theta}J_{f}(\pi_{\theta_{k}},\phi_{\omega_{k}},\hat{\mu}_{k})\|^{2}]+2L_{V}^{2}\alpha_{k}\mathbb{E}[\varepsilon_{f,k}^{V}]+3L_{V}^{2}\xi_{k}^{2}+2B_{F}(B_{D}+B_{H})\alpha_{k}\xi_{k}.$ (82)
Under entropy regularization, $J_{f}$ satisfies a non-uniform PL condition with respect to the follower’s policy [Mei et al., 2020]. Specifically, for any $\theta,\phi,\mu$,
$\|\nabla_{\theta}J_{f}(\pi_{\theta},\phi,\mu)\|^{2}\geq\frac{2(1-\gamma)\tau\rho_{\min}^{2}p_{\min}^{2}}{|{\cal S}|}\Big(J_{f}(\pi^{\star}(\phi,\mu),\phi,\mu)-J_{f}(\pi_{\theta},\phi,\mu)\Big).$ (83)
The inequality in this specific form is adapted from Zeng et al. [2022][Lemma 4].
Combining (82) with (83) leads to
$\mathbb{E}[J_{f}(\pi_{\theta_{k}},\phi_{\omega_{k+1}},\hat{\mu}_{k+1})-J_{f}(\pi_{\theta_{k+1}},\phi_{\omega_{k+1}},\hat{\mu}_{k+1})]$
$\leq-\frac{\alpha_{k}}{16}\mathbb{E}[\|\nabla_{\theta}J_{f}(\pi_{\theta_{k}},\phi_{\omega_{k}},\hat{\mu}_{k})\|^{2}]+2L_{V}^{2}\alpha_{k}\mathbb{E}[\varepsilon_{f,k}^{V}]+3L_{V}^{2}\xi_{k}^{2}+2B_{F}(B_{D}+B_{H})\alpha_{k}\xi_{k}$
$\hskip 20.0pt-\frac{3(1-\gamma)\tau\rho_{\min}^{2}p_{\min}^{2}\alpha_{k}}{8|{\cal S}|}\Big(J_{f}(\pi^{\star}(\phi_{\omega_{k}},\hat{\mu}_{k}),\phi_{\omega_{k}},\hat{\mu}_{k})-J_{f}(\pi_{\theta_{k}},\phi_{\omega_{k}},\hat{\mu}_{k})\Big)$
$\leq-\frac{\alpha_{k}}{16}\mathbb{E}[\|\nabla_{\theta}J_{f}(\pi_{\theta_{k}},\phi_{\omega_{k}},\hat{\mu}_{k})\|^{2}]+2L_{V}^{2}\alpha_{k}\mathbb{E}[\varepsilon_{f,k}^{V}]+3L_{V}^{2}\xi_{k}^{2}+2B_{F}(B_{D}+B_{H})\alpha_{k}\xi_{k}$
$\hskip 20.0pt-\frac{3(1-\gamma)\tau^{2}\rho_{\min}^{3}p_{\min}^{2}\alpha_{k}}{8|{\cal S}|}\|\pi^{\star}(\phi_{\omega_{k}},\hat{\mu}_{k})-\pi_{\theta_{k}}\|^{2}.$ (84)
where the last inequality follows from (77).
Within this subsection, we denote $x_{k}=(\phi_{\omega_{k}},\hat{\mu}_{k})$. Due to the $L_{V}$-smoothness of the function $J_{f}$, we have
$J_{f}(\pi_{\theta_{k}},\phi_{\omega_{k}},\hat{\mu}_{k})-J_{f}(\pi_{\theta_{k}},\phi_{\omega_{k+1}},\hat{\mu}_{k+1})$
$\leq-\langle\nabla_{x}J_{f}(\pi_{\theta_{k}},\phi_{\omega_{k}},\hat{\mu}_{k}),x_{k+1}-x_{k}\rangle+\frac{L_{V}}{2}\|x_{k+1}-x_{k}\|^{2}$
$=-\langle\nabla_{x}J_{f}(\pi_{\theta_{k}},\phi_{\omega_{k}},\hat{\mu}_{k})-\nabla_{x}J_{f}(\pi,\phi_{\omega_{k}},\hat{\mu}_{k})\mid_{\pi=\pi^{\star}(\phi_{\omega_{k}},\hat{\mu}_{k})},x_{k+1}-x_{k}\rangle+\frac{L_{V}}{2}\|x_{k+1}-x_{k}\|^{2}$
$\hskip 20.0pt-\langle\nabla_{x}J_{f}(\pi,\phi_{\omega_{k}},\hat{\mu}_{k})\mid_{\pi=\pi^{\star}(\phi_{\omega_{k}},\hat{\mu}_{k})},x_{k+1}-x_{k}\rangle$
$=-\langle\nabla_{x}J_{f}(\pi_{\theta_{k}},\phi_{\omega_{k}},\hat{\mu}_{k})-\nabla_{x}J_{f}(\pi,\phi_{\omega_{k}},\hat{\mu}_{k})\mid_{\pi=\pi^{\star}(\phi_{\omega_{k}},\hat{\mu}_{k})},x_{k+1}-x_{k}\rangle+\frac{L_{V}}{2}\|x_{k+1}-x_{k}\|^{2}$
$\hskip 20.0pt-\langle\nabla_{x}J_{f}(\pi^{\star}(\phi_{\omega_{k}},\hat{\mu}_{k}),\phi_{\omega_{k}},\hat{\mu}_{k}),x_{k+1}-x_{k}\rangle$
$\leq-\langle\nabla_{x}J_{f}(\pi_{\theta_{k}},\phi_{\omega_{k}},\hat{\mu}_{k})-\nabla_{x}J_{f}(\pi,\phi_{\omega_{k}},\hat{\mu}_{k})\mid_{\pi=\pi^{\star}(\phi_{\omega_{k}},\hat{\mu}_{k})},x_{k+1}-x_{k}\rangle+\frac{L_{V}}{2}\|x_{k+1}-x_{k}\|^{2}$
$\hskip 20.0pt+\Big(J_{f}(\pi^{\star}(\phi_{\omega_{k}},\hat{\mu}_{k}),\phi_{\omega_{k}},\hat{\mu}_{k})-J_{f}(\pi^{\star}(\phi_{\omega_{k+1}},\hat{\mu}_{k+1}),\phi_{\omega_{k+1}},\hat{\mu}_{k+1})\Big)+\frac{L_{V}}{2}\|x_{k+1}-x_{k}\|^{2},$
where the second inequality follows from the fact that for an $L$-smooth function $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$ we have $f(x)-f(y)\geq-\langle\nabla f(x),y-x\rangle-\frac{L}{2}\|y-x\|^{2}$. The second equations above uses the relationship $\nabla_{x}J_{f}(\pi^{\star}(\phi_{\omega_{k}},\hat{\mu}_{k}),\phi_{\omega_{k}},\hat{\mu}_{k})=\nabla_{x}J_{f}(\pi,\phi_{\omega_{k}},\hat{\mu}_{k})\mid_{\pi=\pi^{\star}(\phi_{\omega_{k}},\hat{\mu}_{k})}$, which holds as the partial gradient with respect to $\pi$ is zero at the optimizer (under entropy regularization $\tau>0$, the policy $\pi^{\star}(\phi,\mu)$ is in the interior of the probability simplex for any $\phi,\mu$), i.e.
$\frac{\partial J_{f}(\pi^{\star}(\phi_{\omega_{k}},\hat{\mu}_{k}),\phi_{\omega_{k}},\hat{\mu}_{k})}{\partial\pi^{\star}(\phi_{\omega_{k}},\hat{\mu}_{k})}=0.$
Taking the expectation and plugging in the result from Lemma 6,
$\mathbb{E}[J_{f}(\pi_{\theta_{k}},\phi_{\omega_{k}},\hat{\mu}_{k})-J_{f}(\pi_{\theta_{k}},\phi_{\omega_{k+1}},\hat{\mu}_{k+1})]$
$\leq\frac{(1-\gamma)\tau^{2}\rho_{\min}^{3}p_{\min}^{2}\alpha_{k}}{4|{\cal S}|}\mathbb{E}[\|\pi^{\star}(\phi_{\omega_{k}},\hat{\mu}_{k})-\pi_{\theta_{k}}\|^{2}]+\frac{2|{\cal S}|L_{V}^{2}L_{D}^{2}L_{H}^{2}\xi_{k}^{2}}{(1-\gamma)\tau^{2}\rho_{\min}^{3}p_{\min}^{2}\alpha_{k}}\mathbb{E}[\|\pi^{\star}(\phi_{\omega_{k},\hat{\mu}_{k}})-\pi_{\theta_{k}}\|^{2}]$
$\hskip 20.0pt+\frac{2|{\cal S}|L_{V}^{2}\eta_{2}^{2}\zeta_{k}^{2}}{(1-\gamma)\tau^{2}\rho_{\min}^{3}p_{\min}^{2}\alpha_{k}}\mathbb{E}[\varepsilon_{k}^{\phi}]+\frac{16|{\cal S}|L^{2}L_{V}^{4}L_{D}^{2}L_{H}^{2}\xi_{k}^{2}}{(1-\gamma)\tau^{2}\rho_{\min}^{3}p_{\min}^{2}\alpha_{k}}\mathbb{E}[\varepsilon_{k}^{\mu}]+\frac{8|{\cal S}|L_{V}^{2}\zeta_{k}^{2}}{(1-\gamma)\tau^{2}\rho_{\min}^{3}p_{\min}^{2}\alpha_{k}}\mathbb{E}[\varepsilon_{l,k}^{V}]$
$\hskip 20.0pt+\Big(J_{f}(\pi^{\star}(\phi_{\omega_{k}},\hat{\mu}_{k}),\phi_{\omega_{k}},\hat{\mu}_{k})-J_{f}(\pi^{\star}(\phi_{\omega_{k+1}},\hat{\mu}_{k+1}),\phi_{\omega_{k+1}},\hat{\mu}_{k+1})\Big)+L_{V}\mathbb{E}[\|x_{k+1}-x_{k}\|^{2}].$ (85)
Now, we combine (84) and (85). Recall the definition of $\varepsilon_{k}^{\pi}$ in (19).
$\mathbb{E}[\varepsilon_{k+1}^{\pi}-\varepsilon_{k}^{\pi}]$
$=\mathbb{E}[J_{f}(\pi_{\theta_{k}},\phi_{\omega_{k}},\hat{\mu}_{k})-J_{f}(\pi_{\theta_{k}},\phi_{\omega_{k+1}},\hat{\mu}_{k+1})]$
$\hskip 20.0pt+\mathbb{E}[J_{f}(\pi_{\theta_{k}},\phi_{\omega_{k+1}},\hat{\mu}_{k+1})-J_{f}(\pi_{\theta_{k+1}},\phi_{\omega_{k+1}},\hat{\mu}_{k+1})]$
$\hskip 20.0pt+\mathbb{E}[J_{f}(\pi^{\star}(\phi_{\omega_{k+1}},\hat{\mu}_{k+1}),\phi_{\omega_{k+1}},\hat{\mu}_{k+1})-J_{f}(\pi^{\star}(\phi_{\omega_{k}},\hat{\mu}_{k}),\phi_{\omega_{k}},\hat{\mu}_{k})]$
$\leq\frac{(1-\gamma)\tau^{2}\rho_{\min}^{3}p_{\min}^{2}\alpha_{k}}{4|{\cal S}|}\mathbb{E}[\|\pi^{\star}(\phi_{\omega_{k}},\hat{\mu}_{k})-\pi_{\theta_{k}}\|^{2}]+\frac{2|{\cal S}|L_{V}^{2}L_{D}^{2}L_{H}^{2}\xi_{k}^{2}}{(1-\gamma)\tau^{2}\rho_{\min}^{3}p_{\min}^{2}\alpha_{k}}\mathbb{E}[\|\pi^{\star}(\phi_{\omega_{k},\hat{\mu}_{k}})-\pi_{\theta_{k}}\|^{2}]$
$\hskip 20.0pt+\frac{2|{\cal S}|L_{V}^{2}\eta_{2}^{2}\zeta_{k}^{2}}{(1-\gamma)\tau^{2}\rho_{\min}^{3}p_{\min}^{2}\alpha_{k}}\mathbb{E}[\varepsilon_{k}^{\phi}]+\frac{16|{\cal S}|L^{2}L_{V}^{4}L_{D}^{2}L_{H}^{2}\xi_{k}^{2}}{(1-\gamma)\tau^{2}\rho_{\min}^{3}p_{\min}^{2}\alpha_{k}}\mathbb{E}[\varepsilon_{k}^{\mu}]+\frac{8|{\cal S}|L_{V}^{2}\zeta_{k}^{2}}{(1-\gamma)\tau^{2}\rho_{\min}^{3}p_{\min}^{2}\alpha_{k}}\mathbb{E}[\varepsilon_{l,k}^{V}]$
$\hskip 20.0pt+L_{V}(\|\omega_{k+1}-\omega_{k}\|^{2}+\|\hat{\mu}_{k+1}-\hat{\mu}_{k}\|^{2})$
$\hskip 20.0pt-\frac{\alpha_{k}}{16}\mathbb{E}[\|\nabla_{\theta}J_{f}(\pi_{\theta_{k}},\phi_{\omega_{k}},\hat{\mu}_{k})\|^{2}]+2L_{V}^{2}\alpha_{k}\mathbb{E}[\varepsilon_{f,k}^{V}]+3L_{V}^{2}\xi_{k}^{2}+2B_{F}(B_{D}+B_{H})\alpha_{k}\xi_{k}$
$\hskip 20.0pt-\frac{3(1-\gamma)\tau^{2}\rho_{\min}^{3}p_{\min}^{2}\alpha_{k}}{8|{\cal S}|}\|\pi^{\star}(\phi_{\omega_{k}},\hat{\mu}_{k})-\pi_{\theta_{k}}\|^{2}$
$\leq-\frac{\alpha_{k}}{16}\mathbb{E}[\|\nabla_{\theta}J_{f}(\pi_{\theta_{k}},\phi_{\omega_{k}},\hat{\mu}_{k})\|^{2}]+2L_{V}^{2}\alpha_{k}\mathbb{E}[\varepsilon_{f,k}^{V}]$
$\hskip 20.0pt+\frac{2|{\cal S}|L_{V}^{2}\eta_{2}^{2}\zeta_{k}^{2}}{(1-\gamma)\tau^{2}\rho_{\min}^{3}p_{\min}^{2}\alpha_{k}}\mathbb{E}[\varepsilon_{k}^{\phi}]+\frac{16|{\cal S}|L^{2}L_{V}^{4}L_{D}^{2}L_{H}^{2}\xi_{k}^{2}}{(1-\gamma)\tau^{2}\rho_{\min}^{3}p_{\min}^{2}\alpha_{k}}\mathbb{E}[\varepsilon_{k}^{\mu}]+\frac{8|{\cal S}|L_{V}^{2}\zeta_{k}^{2}}{(1-\gamma)\tau^{2}\rho_{\min}^{3}p_{\min}^{2}\alpha_{k}}\mathbb{E}[\varepsilon_{l,k}^{V}]$
$\hskip 20.0pt+3L_{V}^{2}\xi_{k}^{2}+2B_{F}(B_{D}+B_{H})\alpha_{k}\xi_{k}+L_{V}(B_{D}^{2}\zeta_{k}^{2}+B_{H}^{2}\xi_{k}^{2})$
$\leq-\frac{\alpha_{k}}{16}\mathbb{E}[\|\nabla_{\theta}J_{f}(\pi_{\theta_{k}},\phi_{\omega_{k}},\hat{\mu}_{k})\|^{2}]+\frac{2|{\cal S}|L_{V}^{2}\eta_{2}^{2}\zeta_{k}^{2}}{(1-\gamma)\tau^{2}\rho_{\min}^{3}p_{\min}^{2}\alpha_{k}}\mathbb{E}[\varepsilon_{k}^{\phi}]+\frac{16|{\cal S}|L^{2}L_{V}^{4}L_{D}^{2}L_{H}^{2}\xi_{k}^{2}}{(1-\gamma)\tau^{2}\rho_{\min}^{3}p_{\min}^{2}\alpha_{k}}\mathbb{E}[\varepsilon_{k}^{\mu}]$
$\hskip 20.0pt+\frac{8|{\cal S}|L_{V}^{2}\zeta_{k}^{2}}{(1-\gamma)\tau^{2}\rho_{\min}^{3}p_{\min}^{2}\alpha_{k}}\mathbb{E}[\varepsilon_{l,k}^{V}]+2L_{V}^{2}\alpha_{k}\mathbb{E}[\varepsilon_{f,k}^{V}]+3B_{F}(B_{D}+B_{H})\alpha_{k}\xi_{k},$
where the second inequality uses the step size condition $\frac{\xi_{k}}{\alpha_{k}}\leq\frac{(1-\gamma)\tau^{2}\rho_{\min}^{3}p_{\min}^{2}}{4|{\cal S}|L_{V}L_{D}L_{H}}$, and the last inequality is due to the step size condition $\frac{\zeta_{k}}{\xi_{k}}\leq\frac{B_{H}}{B_{D}}$ and $\frac{\xi_{k}}{\alpha_{k}}\leq\frac{B_{F}(B_{D}+B_{H})}{L_{V}(2B_{H}^{2}+3L_{V})}$.
$\blacksquare$
C.4 Proof of Proposition 4
We introduce a technical lemma below, which bounds an important cross term in the proof of the proposition.
Lemma 7
Under the assumptions and step sizes of Proposition 4, we have for all $k\geq 0$
$\mathbb{E}[\langle\hat{V}_{f,k}-V_{f}^{\pi_{\theta_{k}},\phi_{\omega_{k}},\,\hat{\mu}_{k}}+\beta_{k}\bar{G}_{f}(\theta_{k},\omega_{k},\hat{\mu}_{k},\hat{V}_{f,k}),V_{f}^{\pi_{\theta_{k}},\phi_{\omega_{k}},\,\hat{\mu}_{k}}-V_{f}^{\pi_{\theta_{k+1}},\phi_{\omega_{k+1}},\,\hat{\mu}_{k+1}}\rangle]$
$\leq\frac{(1-\gamma)\beta_{k}}{2}\mathbb{E}[\|\hat{V}_{f,k}-V_{f}^{\pi_{\theta_{k}},\phi_{\omega_{k}},\,\hat{\mu}_{k}}+\beta_{k}\bar{G}_{f}(\theta_{k},\omega_{k},\hat{\mu}_{k},\hat{V}_{f,k})\|^{2}]$
$\hskip 20.0pt+\frac{6L_{V}^{2}\alpha_{k}^{2}}{(1-\gamma)\beta_{k}}\mathbb{E}[\|\nabla_{\theta}J_{f}(\pi_{\theta_{k}},\phi_{\omega_{k}},\hat{\mu}_{k})\|^{2}]+\frac{6L_{V}^{2}L_{H}^{2}\xi_{k}^{2}}{(1-\gamma)\beta_{k}}\mathbb{E}[\varepsilon_{k}^{\pi}]+\frac{6L_{V}^{2}\eta_{2}^{2}\zeta_{k}^{2}}{(1-\gamma)\beta_{k}}\mathbb{E}[\varepsilon_{k}^{\phi}]$
$\hskip 20.0pt+\frac{16L^{2}L_{V}^{2}L_{H}^{2}\xi_{k}^{2}}{(1-\gamma)\beta_{k}}\mathbb{E}[\varepsilon_{k}^{\mu}]+\frac{6L_{V}^{4}\zeta_{k}^{2}}{(1-\gamma)\beta_{k}}\mathbb{E}[\varepsilon_{f,k}^{V}]+\frac{24L_{V}^{2}L_{D}^{2}\alpha_{k}^{2}}{(1-\gamma)\beta_{k}}\mathbb{E}[\varepsilon_{l,k}^{V}]+L_{VV}(B_{F}+B_{D}+B_{H})^{2}\alpha_{k}^{2}.$
The proof of the lemma can be found in Section D.7.
By the definition of $\varepsilon_{f,k}^{V}$ and the update rule $\hat{V}_{f,k}$,
$\varepsilon_{f,k+1}^{V}$ $=\|\hat{V}_{f,k+1}-V_{f}^{\pi_{\theta_{k+1}},\phi_{\omega_{k+1}},\,\hat{\mu}_{k+1}}\|^{2}$
$=\|\Pi_{B_{V}}\Big(\hat{V}_{f,k}+\beta_{k}G_{f}(\theta_{k},\hat{\mu}_{k},\hat{V}_{f,k},s_{k},a_{k},b_{k},s_{k}^{\prime})\Big)-V_{f}^{\pi_{\theta_{k+1}},\phi_{\omega_{k+1}},\,\hat{\mu}_{k+1}}\|^{2}$
$\leq\Big\|\hat{V}_{f,k}+\beta_{k}G_{f}(\theta_{k},\hat{\mu}_{k},\hat{V}_{f,k},s_{k},a_{k},b_{k},s_{k}^{\prime})-V_{f}^{\pi_{\theta_{k+1}},\phi_{\omega_{k+1}},\,\hat{\mu}_{k+1}}\Big\|^{2}$
$=\Big\|\hat{V}_{f,k}-V_{f}^{\pi_{\theta_{k}},\phi_{\omega_{k}},\,\hat{\mu}_{k}}+\beta_{k}\bar{G}_{f}(\theta_{k},\omega_{k},\hat{\mu}_{k},\hat{V}_{f,k})+\beta_{k}\Big(G_{f}(\theta_{k},\hat{\mu}_{k},\hat{V}_{f,k},s_{k},a_{k},b_{k},s_{k}^{\prime})-\bar{G}_{f}(\theta_{k},\omega_{k},\hat{\mu}_{k},\hat{V}_{f,k})\Big)$
$\hskip 20.0pt+V_{f}^{\pi_{\theta_{k}},\phi_{\omega_{k}},\,\hat{\mu}_{k}}-V_{f}^{\pi_{\theta_{k+1}},\phi_{\omega_{k+1}},\,\hat{\mu}_{k+1}}\Big\|^{2}$
$\leq\Big\|\hat{V}_{f,k}-V_{f}^{\pi_{\theta_{k}},\phi_{\omega_{k}},\,\hat{\mu}_{k}}+\beta_{k}\bar{G}_{f}(\theta_{k},\omega_{k},\hat{\mu}_{k},\hat{V}_{f,k})\Big\|^{2}$
$\hskip 20.0pt+2\beta_{k}^{2}\|G_{f}(\theta_{k},\hat{\mu}_{k},\hat{V}_{f,k},s_{k},a_{k},b_{k},s_{k}^{\prime})-\bar{G}_{f}(\theta_{k},\omega_{k},\hat{\mu}_{k},\hat{V}_{f,k})\|^{2}$
$\hskip 20.0pt+2\Big\|V_{f}^{\pi_{\theta_{k}},\phi_{\omega_{k}},\,\hat{\mu}_{k}}-V_{f}^{\pi_{\theta_{k+1}},\phi_{\omega_{k+1}},\,\hat{\mu}_{k+1}}\Big\|^{2}$
$\hskip 20.0pt+\beta_{k}\langle\hat{V}_{f,k}-V_{f}^{\pi_{\theta_{k}},\phi_{\omega_{k}},\,\hat{\mu}_{k}}+\beta_{k}\bar{G}_{f}(\theta_{k},\omega_{k},\hat{\mu}_{k},\hat{V}_{f,k}),G_{f}(\theta_{k},\hat{\mu}_{k},\hat{V}_{f,k},s_{k},a_{k},b_{k},s_{k}^{\prime})-\bar{G}_{f}(\theta_{k},\omega_{k},\hat{\mu}_{k},\hat{V}_{f,k})\rangle$
$\hskip 20.0pt+\langle\hat{V}_{f,k}-V_{f}^{\pi_{\theta_{k}},\phi_{\omega_{k}},\,\hat{\mu}_{k}}+\beta_{k}\bar{G}_{f}(\theta_{k},\omega_{k},\hat{\mu}_{k},\hat{V}_{f,k}),V_{f}^{\pi_{\theta_{k}},\phi_{\omega_{k}},\,\hat{\mu}_{k}}-V_{f}^{\pi_{\theta_{k+1}},\phi_{\omega_{k+1}},\,\hat{\mu}_{k+1}}\rangle,$ (86)
where the first inequality follows from the fact that the projection to a convex set is non-expansive.
To bound the first term of (86),
$\Big\|\hat{V}_{f,k}-V_{f}^{\pi_{\theta_{k}},\phi_{\omega_{k}},\,\hat{\mu}_{k}}+\beta_{k}\bar{G}_{f}(\theta_{k},\omega_{k},\hat{\mu}_{k},\hat{V}_{f,k})\Big\|^{2}$
$=\|\hat{V}_{f,k}-V_{f}^{\pi_{\theta_{k}},\phi_{\omega_{k}},\,\hat{\mu}_{k}}\|^{2}+\beta_{k}^{2}\|\bar{G}_{f}(\theta_{k},\omega_{k},\hat{\mu}_{k},\hat{V}_{f,k})\|^{2}+2\beta_{k}\langle\hat{V}_{f,k}-V_{f}^{\pi_{\theta_{k}},\phi_{\omega_{k}},\,\hat{\mu}_{k}},\bar{G}_{f}(\theta_{k},\omega_{k},\hat{\mu}_{k},\hat{V}_{f,k})\rangle$
$=\|\hat{V}_{f,k}-V_{f}^{\pi_{\theta_{k}},\phi_{\omega_{k}},\,\hat{\mu}_{k}}\|^{2}+\beta_{k}^{2}\|\bar{G}_{f}(\theta_{k},\omega_{k},\hat{\mu}_{k},\hat{V}_{f,k})-\bar{G}_{f}(\theta_{k},\omega_{k},\hat{\mu}_{k},V_{f}^{\pi_{\theta_{k}},\phi_{\omega_{k}},\,\hat{\mu}_{k}})\|^{2}$
$\hskip 20.0pt+2\beta_{k}\langle\hat{V}_{f,k}-V_{f}^{\pi_{\theta_{k}},\phi_{\omega_{k}},\,\hat{\mu}_{k}},\bar{G}_{f}(\theta_{k},\omega_{k},\hat{\mu}_{k},\hat{V}_{f,k})-\bar{G}_{f}(\theta_{k},\omega_{k},\hat{\mu}_{k},V_{f}^{\pi_{\theta_{k}},\phi_{\omega_{k}},\,\hat{\mu}_{k}})\rangle$
$=\|\hat{V}_{f,k}-V_{f}^{\pi_{\theta_{k}},\phi_{\omega_{k}},\,\hat{\mu}_{k}}\|^{2}+\beta_{k}^{2}\|\bar{G}_{f}(\theta_{k},\omega_{k},\hat{\mu}_{k},\hat{V}_{f,k})-\bar{G}_{f}(\theta_{k},\omega_{k},\hat{\mu}_{k},V_{f}^{\pi_{\theta_{k}},\phi_{\omega_{k}},\,\hat{\mu}_{k}})\|^{2}$
$\hskip 20.0pt+2\beta_{k}\Big(\hat{V}_{f,k}-V_{f}^{\pi_{\theta_{k}},\phi_{\omega_{k}},\,\hat{\mu}_{k}}\Big)^{\top}(\gamma P^{\pi_{\theta_{k}},\phi_{\omega_{k}},\,\hat{\mu}_{k}}-I)\Big(\hat{V}_{f,k}-V_{f}^{\pi_{\theta_{k}},\phi_{\omega_{k}},\,\hat{\mu}_{k}}\Big)$
$\leq\|\hat{V}_{f,k}-V_{f}^{\pi_{\theta_{k}},\phi_{\omega_{k}},\,\hat{\mu}_{k}}\|^{2}+L_{G}^{2}\beta_{k}^{2}\|\hat{V}_{f,k}-V_{f}^{\pi_{\theta_{k}},\phi_{\omega_{k}},\,\hat{\mu}_{k}}\|^{2}+2(\gamma-1)\beta_{k}\|\hat{V}_{f,k}-V_{f}^{\pi_{\theta_{k}},\phi_{\omega_{k}},\,\hat{\mu}_{k}}\|^{2}$
$\leq(1-(1-\gamma)\beta_{k})\varepsilon_{f,k}^{V},$ (87)
where the last inequality follows from the step size condition $\beta_{k}\leq\frac{1-\gamma}{L_{G}^{2}}$.
For the third term of (86),
$2\|V_{f}^{\pi_{\theta_{k}},\phi_{\omega_{k}},\,\hat{\mu}_{k}}-V_{f}^{\pi_{\theta_{k+1}},\phi_{\omega_{k+1}},\,\hat{\mu}_{k+1}}\|^{2}$ $\leq 6L_{V}^{2}\Big(\|\pi_{\theta_{k+1}}-\pi_{\theta_{k}}\|^{2}+\|\phi_{\omega_{k+1}}-\phi_{\omega_{k}}\|^{2}+\|\hat{\mu}_{k+1}-\hat{\mu}_{k}\|^{2}\Big)2$
$\leq 6L_{V}^{2}\Big(B_{F}^{2}\alpha_{k}^{2}+B_{D}^{2}\zeta_{k}^{2}+B_{H}^{2}\xi_{k}^{2}\Big)$
$\leq 6L_{V}^{2}(B_{F}+B_{D}+B_{H})^{2}\alpha_{k}^{2}.$ (88)
Note that the fourth term of (86) vanishes in expectation. Collecting the bounds from (87), (88), and Lemma 7, we have
$\mathbb{E}[\varepsilon_{f,k+1}^{V}]$ $\leq(1-(1-\gamma)\beta_{k})\mathbb{E}[\varepsilon_{f,k}^{V}]+2\beta_{k}^{2}\mathbb{E}[\|G_{f}(\theta_{k},\hat{\mu}_{k},\hat{V}_{f,k},s_{k},a_{k},b_{k},s_{k}^{\prime})-\bar{G}_{f}(\theta_{k},\omega_{k},\hat{\mu}_{k},\hat{V}_{f,k})\|^{2}]$
$\hskip 20.0pt+6L_{V}^{2}(B_{F}+B_{D}+B_{H})^{2}\alpha_{k}^{2}$
$\hskip 20.0pt+\frac{(1-\gamma)\beta_{k}}{2}\mathbb{E}[\|\hat{V}_{f,k}-V_{f}^{\pi_{\theta_{k}},\phi_{\omega_{k}},\,\hat{\mu}_{k}}+\beta_{k}\bar{G}_{f}(\theta_{k},\omega_{k},\hat{\mu}_{k},\hat{V}_{f,k})\|^{2}]$
$\hskip 20.0pt+\frac{6L_{V}^{2}\alpha_{k}^{2}}{(1-\gamma)\beta_{k}}\mathbb{E}[\|\nabla_{\theta}J_{f}(\pi_{\theta_{k}},\phi_{\omega_{k}},\hat{\mu}_{k})\|^{2}]+\frac{6L_{V}^{2}L_{H}^{2}\xi_{k}^{2}}{(1-\gamma)\beta_{k}}\mathbb{E}[\varepsilon_{k}^{\pi}]+\frac{6L_{V}^{2}\eta_{2}^{2}\zeta_{k}^{2}}{(1-\gamma)\beta_{k}}\mathbb{E}[\varepsilon_{k}^{\phi}]$
$\hskip 20.0pt+\frac{16L^{2}L_{V}^{2}L_{H}^{2}\xi_{k}^{2}}{(1-\gamma)\beta_{k}}\mathbb{E}[\varepsilon_{k}^{\mu}]+\frac{6L_{V}^{4}\zeta_{k}^{2}}{(1-\gamma)\beta_{k}}\mathbb{E}[\varepsilon_{f,k}^{V}]+\frac{24L_{V}^{2}L_{D}^{2}\alpha_{k}^{2}}{(1-\gamma)\beta_{k}}\mathbb{E}[\varepsilon_{l,k}^{V}]+L_{VV}(B_{F}+B_{D}+B_{H})^{2}\alpha_{k}^{2}$
$\leq(1-\frac{(1-\gamma)\beta_{k}}{2}+\frac{6L_{V}^{4}\zeta_{k}^{2}}{(1-\gamma)\beta_{k}})\mathbb{E}[\varepsilon_{f,k}^{V}]+8B_{G}^{2}\beta_{k}^{2}+(6L_{V}^{2}+L_{VV})(B_{F}+B_{D}+B_{H})^{2}\alpha_{k}^{2}$
$\hskip 20.0pt+\frac{6L_{V}^{2}\alpha_{k}^{2}}{(1-\gamma)\beta_{k}}\mathbb{E}[\|\nabla_{\theta}J_{f}(\pi_{\theta_{k}},\phi_{\omega_{k}},\hat{\mu}_{k})\|^{2}]+\frac{6L_{V}^{2}L_{H}^{2}\xi_{k}^{2}}{(1-\gamma)\beta_{k}}\mathbb{E}[\varepsilon_{k}^{\pi}]+\frac{6L_{V}^{2}\eta_{2}^{2}\zeta_{k}^{2}}{(1-\gamma)\beta_{k}}\mathbb{E}[\varepsilon_{k}^{\phi}]$
$\hskip 20.0pt+\frac{16L^{2}L_{V}^{2}L_{H}^{2}\xi_{k}^{2}}{(1-\gamma)\beta_{k}}\mathbb{E}[\varepsilon_{k}^{\mu}]+\frac{24L_{V}^{2}L_{D}^{2}\alpha_{k}^{2}}{(1-\gamma)\beta_{k}}\mathbb{E}[\varepsilon_{l,k}^{V}]$
$\leq(1-\frac{(1-\gamma)\beta_{k}}{4})\mathbb{E}[\varepsilon_{f,k}^{V}]+\frac{6L_{V}^{2}\alpha_{k}^{2}}{(1-\gamma)\beta_{k}}\mathbb{E}[\|\nabla_{\theta}J_{f}(\pi_{\theta_{k}},\phi_{\omega_{k}},\hat{\mu}_{k})\|^{2}]+\frac{6L_{V}^{2}L_{H}^{2}\xi_{k}^{2}}{(1-\gamma)\beta_{k}}\mathbb{E}[\varepsilon_{k}^{\pi}]$
$\hskip 20.0pt+\frac{6L_{V}^{2}\eta_{2}^{2}\zeta_{k}^{2}}{(1-\gamma)\beta_{k}}\mathbb{E}[\varepsilon_{k}^{\phi}]+\frac{16L^{2}L_{V}^{2}L_{H}^{2}\xi_{k}^{2}}{(1-\gamma)\beta_{k}}\mathbb{E}[\varepsilon_{k}^{\mu}]+\frac{24L_{V}^{2}L_{D}^{2}\alpha_{k}^{2}}{(1-\gamma)\beta_{k}}\mathbb{E}[\varepsilon_{l,k}^{V}]+12B_{G}^{2}\beta_{k}^{2},$
where the second inequality follows from $\|\hat{V}_{f,k}-V_{f}^{\pi_{\theta_{k}},\phi_{\omega_{k}},\,\hat{\mu}_{k}}+\beta_{k}\bar{G}_{f}(\theta_{k},\omega_{k},\hat{\mu}_{k},\hat{V}_{f,k})\|^{2}\leq\varepsilon_{f,k}^{V}$ established in (87), and the last inequality follows from the step size conditions $\frac{\zeta_{k}}{\beta_{k}}\leq\frac{1-\gamma}{2\sqrt{6}L_{V}^{2}}$ and $\frac{\alpha_{k}}{\beta_{k}}\leq\frac{2B_{G}}{\sqrt{6L_{V}^{2}+L_{VV}}(B_{F}+B_{D}+B_{H})}$.
The bound on $\mathbb{E}[\varepsilon_{l,k}^{V}]$ can be established using an identical argument, under the step size condition $\frac{\alpha_{k}}{\beta_{k}}\leq\frac{1-\gamma}{4\sqrt{6}L_{V}L_{D}}$.
$\blacksquare$
Appendix D Proof of Lemmas
D.1 Proof of Lemma 1
We focus on proving the inequalities under the follower’s reward and note that the same argument can be used for the leader’s cumulative reward and value function.
We show the Lipschitz continuity conditions first. Fixing the mean field and leader’s policy, the follower’s problem reduces to a standard infinite-horizon discounted-reward MDP, in which it is well-known that the value function is Lipschitz and smooth with respect to the follower’s policy parameter. Specifically, adapting Zeng et al. [2021][Lemma B.5], we have
$\|V_{f}^{\pi_{\theta},\phi_{\omega},\mu}-V_{f}^{\pi_{\theta^{\prime}},\phi_{\omega^{\prime}},\mu}\|\leq\frac{2}{(1-\gamma)^{2}}(\|\theta-\theta^{\prime}\|+\|\phi_{\omega}-\phi_{\omega^{\prime}}\|).$ (89)
This means that to show (40) it suffices for us to show the Lipschitz continuity of $V_{f}^{\pi_{\theta},\phi_{\omega},\mu}$ with respect to $\mu$. The value function can be expressed as
$V_{f}^{\pi,\phi,\mu}(s)$ $=\frac{1}{1-\gamma}\sum_{s^{\prime},a,b}d_{\rho}^{\pi,\,\phi,\,\mu}(s^{\prime})\pi(a\mid s^{\prime})\phi(b\mid s^{\prime})r_{f}(s,a,b,\mu)$
$=\frac{1}{1-\gamma}\langle d_{\rho}^{\pi,\,\phi,\,\mu},r_{f}^{\pi,\phi,\mu}\rangle$
$=e_{s}^{\top}(I-\gamma P^{\pi,\phi,\mu})^{-\top}r_{f}^{\pi,\phi,\mu},$ (90)
where $e_{s}\in\mathbb{R}^{|{\cal S}|}$ denotes a vector with 1 at entry $s$ and zero otherwise, and $r_{f}^{\pi,\phi,\mu}\in\mathbb{R}^{|{\cal S}|}$ denotes the marginalized reward such that
$r_{f}^{\pi,\phi,\mu}(s)=\sum_{a,b}r_{f}(s,a,b,\mu)\pi(a\mid s)\phi(b\mid s).$
We have from (90)
$\|V_{f}^{\pi,\phi,\mu}-V_{f}^{\pi,\phi,\mu^{\prime}}\|$
$=\Big((I-\gamma P^{\pi,\phi,\mu})^{-\top}r_{f}^{\pi,\phi,\mu}-(I-\gamma P^{\pi,\phi,\mu^{\prime}})^{-\top}r_{f}^{\pi,\phi,\mu^{\prime}}\Big)$
$=\Big((I-\gamma P^{\pi,\phi,\mu})^{-\top}-(I-\gamma P^{\pi,\phi,\mu^{\prime}})^{-\top}\Big)r_{f}^{\pi,\phi,\mu}+(I-\gamma P^{\pi,\phi,\mu^{\prime}})^{-\top}\Big(r_{f}^{\pi,\phi,\mu}-r_{f}^{\pi,\phi,\mu^{\prime}}\Big)$
$=(I-\gamma P^{\pi,\phi,\mu})^{-\top}\Big((I-\gamma P^{\pi,\phi,\mu^{\prime}})^{\top}-(I-\gamma P^{\pi,\phi,\mu})^{\top}\Big)(I-\gamma P^{\pi,\phi,\mu^{\prime}})^{-\top}r_{f}^{\pi,\phi,\mu}$
$\hskip 20.0pt+(I-\gamma P^{\pi,\phi,\mu^{\prime}})^{-\top}\Big(r_{f}^{\pi,\phi,\mu}-r_{f}^{\pi,\phi,\mu^{\prime}}\Big)$
$\leq\|(I-\gamma P^{\pi,\phi,\mu})^{-\top}\|\Big\|(I-\gamma P^{\pi,\phi,\mu^{\prime}})^{\top}-(I-\gamma P^{\pi,\phi,\mu})^{\top}\Big\|\|(I-\gamma P^{\pi,\phi,\mu^{\prime}})^{-\top}\|\|r_{f}^{\pi,\phi,\mu}\|$
$\hskip 20.0pt+\|(I-\gamma P^{\pi,\phi,\mu^{\prime}})^{-\top}\|\Big\|r_{f}^{\pi,\phi,\mu}-r_{f}^{\pi,\phi,\mu^{\prime}}\Big\|$
$\leq\frac{1}{1-\gamma}\cdot\gamma\|P^{\pi,\phi,\mu}-P^{\pi,\phi,\mu^{\prime}}\|\cdot\frac{1}{1-\gamma}\cdot\sqrt{|{\cal S}|}+\frac{1}{1-\gamma}\Big\|r_{f}^{\pi,\phi,\mu}-r_{f}^{\pi,\phi,\mu^{\prime}}\Big\|$
$\leq\frac{L_{P}\gamma\sqrt{|{\cal S}|}}{(1-\gamma)^{2}}\|\mu-\mu^{\prime}\|+\frac{L_{r}}{1-\gamma}\|\mu-\mu^{\prime}\|$
$\leq\Big(\frac{L_{P}\gamma\sqrt{|{\cal S}|}}{(1-\gamma)^{2}}+\frac{L_{r}}{1-\gamma}\Big)\|\mu-\mu^{\prime}\|,$ (91)
where the second last inequality follows from Assumption 1.
Combining (89) and (91) implies (38), which obviously leads to (37) as $J_{f}(\pi,\phi,\mu)$ is simply $\langle V_{f}^{\pi,\phi,\mu},\rho\rangle$.
To show the smoothness conditions, we note that again adapting Zeng et al. [2021][Lemma B.5], we have
$\|\nabla_{\theta}V_{f}^{\pi_{\theta},\phi_{\omega},\mu}-\nabla_{\theta}V_{f}^{\pi_{\theta^{\prime}},\phi_{\omega},\mu}\|\leq\frac{8}{(1-\gamma)^{3}}\|\theta-\theta^{\prime}\|,\;\|\nabla_{\omega}V_{f}^{\pi_{\theta},\phi_{\omega},\mu}-\nabla_{\omega}V_{f}^{\pi_{\theta},\phi_{\omega^{\prime}},\mu}\|\leq\frac{8}{(1-\gamma)^{3}}\|\omega-\omega^{\prime}\|,$
This implies (39).
To show (40), we differentiate $V_{f}^{\pi,\phi,\mu}$ with respect to $\mu$. It can be seen from (90) that
$\nabla_{\mu}V_{f}^{\pi,\phi,\mu}=\gamma\left(I-\gamma P^{\pi,\phi,\mu}\right)^{-\top}\left(\nabla_{\mu}P^{\pi,\phi,\mu}\right)^{\top}\left(I-\gamma P^{\pi,\phi,\mu}\right)^{-\top}r_{f}^{\pi,\phi,\mu}+\left(I-\gamma P^{\pi,\phi,\mu}\right)^{-\top}\nabla_{\mu}r_{f}^{\pi,\phi,\mu}.$
Therefore, we have
$\|\nabla_{\mu}V_{f}^{\pi,\phi,\mu}-\nabla_{\mu}V_{f}^{\pi,\phi,\mu^{\prime}}\|$
$\leq\gamma\|\left(I-\gamma P^{\pi,\phi,\mu}\right)^{-\top}\hskip-5.0pt\left(\nabla_{\mu}P^{\pi,\phi,\mu}\right)^{\top}\hskip-5.0pt\left(I-\gamma P^{\pi,\phi,\mu}\right)^{-\top}\hskip-5.0ptr_{f}^{\pi,\phi,\mu}\hskip-2.0pt-\hskip-2.0pt\left(I-\gamma P^{\pi,\phi,\mu^{\prime}}\right)^{-\top}\hskip-5.0pt\left(\nabla_{\mu}P^{\pi,\phi,\mu^{\prime}}\right)^{\top}\left(I-\gamma P^{\pi,\phi,\mu^{\prime}}\right)^{-\top}\hskip-5.0ptr_{f}^{\pi,\phi,\mu^{\prime}}\|$
$\hskip 20.0pt+\gamma\|\left(I-\gamma P^{\pi,\phi,\mu}\right)^{-\top}\nabla_{\mu}r_{f}^{\pi,\phi,\mu}-\left(I-\gamma P^{\pi,\phi,\mu^{\prime}}\right)^{-\top}\nabla_{\mu}r_{f}^{\pi,\phi,\mu^{\prime}}\|$
$\leq\gamma\|\left(I-\gamma P^{\pi,\phi,\mu}\right)^{-\top}-\left(I-\gamma P^{\pi,\phi,\mu^{\prime}}\right)^{-\top}\|\cdot L_{P}\cdot\frac{1}{1-\gamma}\cdot\sqrt{|{\cal S}|}$
$\hskip 20.0pt+\gamma\frac{1}{1-\gamma}\cdot\|\left(\nabla_{\mu}P^{\pi,\phi,\mu}\right)^{\top}-\left(\nabla_{\mu}P^{\pi,\phi,\mu^{\prime}}\right)^{\top}\|\cdot\frac{1}{1-\gamma}\cdot\sqrt{|{\cal S}|}$
$\hskip 20.0pt+\gamma\frac{1}{1-\gamma}\cdot L_{P}\cdot\|\left(I-\gamma P^{\pi,\phi,\mu}\right)^{-\top}-\left(I-\gamma P^{\pi,\phi,\mu^{\prime}}\right)^{-\top}\|\cdot\sqrt{|{\cal S}|}$
$\hskip 20.0pt+\gamma\frac{1}{1-\gamma}\cdot L_{P}\cdot\frac{1}{1-\gamma}\cdot\|\nabla_{\mu}r_{f}^{\pi,\phi,\mu}-\nabla_{\mu}r_{f}^{\pi,\phi,\mu^{\prime}}\|$
$\hskip 20.0pt+\gamma\|\left(\nabla_{\mu}P^{\pi,\phi,\mu}\right)^{\top}-\left(\nabla_{\mu}P^{\pi,\phi,\mu^{\prime}}\right)^{\top}\|\cdot L_{r}$
$\hskip 20.0pt+\gamma L_{P}\cdot\|\nabla_{\mu}r_{f}^{\pi,\phi,\mu}-\nabla_{\mu}r_{f}^{\pi,\phi,\mu^{\prime}}\|$
$\leq\frac{2\gamma L_{P}\sqrt{|{\cal S}|}}{1-\gamma}\|\left(I-\gamma P^{\pi,\phi,\mu}\right)^{-\top}-\left(I-\gamma P^{\pi,\phi,\mu^{\prime}}\right)^{-\top}\|+\Big(\frac{\gamma\sqrt{|{\cal S}|}}{(1-\gamma)^{2}}+\gamma L_{r}\Big)\|\nabla_{\mu}P^{\pi,\phi,\mu}-\nabla_{\mu}P^{\pi,\phi,\mu^{\prime}}\|$
$\hskip 20.0pt+\Big(\frac{\gamma L_{P}}{(1-\gamma)^{2}}+\gamma L_{P}\Big)\|\nabla_{\mu}r_{f}^{\pi,\phi,\mu}-\nabla_{\mu}r_{f}^{\pi,\phi,\mu^{\prime}}\|$
$\leq\frac{2\gamma L_{P}\sqrt{|{\cal S}|}}{1-\gamma}\|\left(I-\gamma P^{\pi,\phi,\mu}\right)^{-\top}\|\|\left(I-\gamma P^{\pi,\phi,\mu^{\prime}}\right)^{\top}-\left(I-\gamma P^{\pi,\phi,\mu}\right)^{\top}\|\|\left(I-\gamma P^{\pi,\phi,\mu^{\prime}}\right)^{-\top}\|$
$\hskip 20.0pt+\Big(\frac{\gamma\sqrt{|{\cal S}|}}{(1-\gamma)^{2}}+\gamma L_{r}\Big)L_{P}\|\mu-\mu^{\prime}\|+\Big(\frac{\gamma L_{P}}{(1-\gamma)^{2}}+\gamma L_{P}\Big)L_{r}\|\mu-\mu^{\prime}\|$
$\leq\frac{2\gamma L_{P}\sqrt{|{\cal S}|}}{(1-\gamma)^{3}}\cdot\gamma\|\nabla_{\mu}P^{\pi,\phi,\mu}-\nabla_{\mu}P^{\pi,\phi,\mu^{\prime}}\|+2\Big(\frac{\gamma\sqrt{|{\cal S}|}L_{P}}{(1-\gamma)^{2}}+\gamma L_{P}L_{r}\Big)\|\mu-\mu^{\prime}\|$
$\leq\Big(\frac{2\gamma^{2}\sqrt{|{\cal S}|}L_{P}^{2}}{(1-\gamma)^{3}}+\frac{2\gamma\sqrt{|{\cal S}|}L_{P}}{(1-\gamma)^{2}}+2\gamma L_{P}L_{r}\Big)\|\mu-\mu^{\prime}\|,$
where we have plugged in the smoothness bounds on $P$ and $r_{f}$ from Assumption 1.
$\blacksquare$
D.2 Proof of Lemma 2
Note that the discounted occupancy measure can be expressed as
$d_{\rho}^{\pi,\phi,\mu}=(1-\gamma)(I-\gamma P^{\pi,\phi,\mu})^{-1}\rho.$ (92)
This implies
$\Big\|\sum_{s}\big(d_{\rho}^{\pi_{\theta_{1}},\,\phi_{\omega_{1}}\,\mu_{1}}(s)P^{\pi_{\theta_{1}},\,\phi_{\omega_{1}}\,\mu_{1}}(\cdot\mid s)-d_{\rho}^{\pi_{\theta_{2}},\,\phi_{\omega_{2}}\,\mu_{2}}(s)P^{\pi_{\theta_{2}},\,\phi_{\omega_{2}}\,\mu_{2}}(\cdot\mid s)\big)\Big\|$
$=(1-\gamma)\|P^{\pi_{\theta_{1}},\,\phi_{\omega_{1}}\,\mu_{1}}(I-\gamma P^{\pi_{\theta_{1}},\,\phi_{\omega_{1}}\,\mu_{1}})^{-1}\rho-P^{\pi_{\theta_{2}},\,\phi_{\omega_{2}}\,\mu_{2}}(I-\gamma P^{\pi_{\theta_{2}},\,\phi_{\omega_{2}}\,\mu_{2}})^{-1}\rho\|$
$\leq(1-\gamma)\|\rho\|\|P^{\pi_{\theta_{1}},\,\phi_{\omega_{1}}\,\mu_{1}}-P^{\pi_{\theta_{2}},\,\phi_{\omega_{2}}\,\mu_{2}}\|\|(I-\gamma P^{\pi_{\theta_{1}},\,\phi_{\omega_{1}}\,\mu_{1}})^{-1}\|$
$\hskip 20.0pt+(1-\gamma)\|\rho\|\|P^{\pi_{\theta_{2}},\,\phi_{\omega_{2}}\,\mu_{2}}\|\|(I-\gamma P^{\pi_{\theta_{1}},\,\phi_{\omega_{1}}\,\mu_{1}})^{-1}-(I-\gamma P^{\pi_{\theta_{2}},\,\phi_{\omega_{2}}\,\mu_{2}})^{-1}\|$
$\leq(1-\gamma)\Big(L_{P}\|\mu_{1}-\mu_{2}\|+\|\pi_{\theta_{1}}-\pi_{\theta_{2}}\|+\|\phi_{\omega_{1}}-\phi_{\omega_{2}}\|\Big)\cdot\frac{1}{1-\gamma}$
$\hskip 20.0pt+(1-\gamma)\|(I-\gamma P^{\pi_{\theta_{2}},\,\phi_{\omega_{2}}\,\mu_{2}})^{-1}\|\|(I-\gamma P^{\pi_{\theta_{2}},\,\phi_{\omega_{2}}\,\mu_{2}})-(I-\gamma P^{\pi_{\theta_{1}},\,\phi_{\omega_{1}}\,\mu_{1}})\|\|(I-\gamma P^{\pi_{\theta_{1}},\,\phi_{\omega_{1}}\,\mu_{1}})^{-1}\|$
$\leq\Big(L_{P}\|\mu_{1}-\mu_{2}\|+\|\pi_{\theta_{1}}-\pi_{\theta_{2}}\|+\|\phi_{\omega_{1}}-\phi_{\omega_{2}}\|\Big)+\frac{\gamma}{1-\gamma}\|P^{\pi_{\theta_{1}},\,\phi_{\omega_{1}}\,\mu_{1}}-P^{\pi_{\theta_{2}},\,\phi_{\omega_{2}}\,\mu_{2}}\|$
$\leq\frac{1}{1-\gamma}\Big(L_{P}\|\mu_{1}-\mu_{2}\|+\|\pi_{\theta_{1}}-\pi_{\theta_{2}}\|+\|\phi_{\omega_{1}}-\phi_{\omega_{2}}\|\Big)$
$\leq\frac{1}{1-\gamma}\Big(L_{P}\|\mu_{1}-\mu_{2}\|+\|\theta_{1}-\theta_{2}\|+\|\omega_{1}-\omega_{2}\|\Big),$
where the second inequality follows from Assumption 1 and the fact that the state transition matrix is $1$-Lipschitz in the policies, and the final inequality is due to the $1$-Lipschitz continuity of the softmax function.
$\blacksquare$
D.3 Proof of Lemma 3
Note that $\nabla_{\theta}\log\pi_{\theta}(a\mid s)$ can be expressed entry-wise as
$\frac{\partial\log\pi_{\theta}(a\mid s)}{\partial\theta_{s^{\prime},a^{\prime}}}={\mathbf{1}}\left[s=s^{\prime}\right]\left({\mathbf{1}}\left[a=a^{\prime}\right]-\pi_{\theta}\left(a^{\prime}\mid s\right)\right),$ (93)
which implies
$\|\nabla_{\theta}\log\pi_{\theta}(a\mid s)\|_{2}\leq\|\nabla_{\theta}\log\pi_{\theta}(a\mid s)\|_{1}\leq 1+1=2.$ (94)
By the definition of $F$,
$\|F(\theta,\mu,V_{f},s,a,b,s^{\prime})\|$ $\leq\|\nabla_{\theta}\log\pi_{\theta}(a\mid s)\|\Big|r_{f}(s,a,b,\mu)+\tau E(\pi_{\theta},s)+\gamma V_{f}(s^{\prime})-V_{f}(s)\Big|$
$\leq 2(1+\tau\log|{\cal A}|+\gamma B_{V}+B_{V})$
$=2(1+\gamma)B_{V}+2\tau\log|{\cal A}|+2,$
where the second inequality applies (94).
By a similar argument, we can show
$\|D(\omega,\mu,V_{l},s,b,s^{\prime})\|\leq 2(1+\gamma)B_{V}+2.$
For $H$, we have
$\|H(\mu,s)\|=\left\|e_{s}-\mu\right\|\leq\left\|e_{s}\right\|+\|\mu\|\leq 2.$
For the $G$ operators,
$\|G_{f}(\theta,\mu,V_{f},s,a,b,s^{\prime})\|$ $\leq\|e_{s}\|\Big|r_{f}(s,a,b,\mu)+\tau E(\pi_{\theta},s)+\gamma V_{f}(s^{\prime})-V_{f}(s)\Big|$
$\leq 1\cdot(1+\tau\log|{\cal A}|+\gamma B_{V}+B_{V})$
$=(1+\gamma)B_{V}+\tau\log|{\cal A}|+1.$
Similarly,
$\|G_{l}(\omega,\mu,V_{l},s,b,s^{\prime})\|\leq(1+\gamma)B_{V}+1.$
$\blacksquare$
D.4 Proof of Lemma 4
By the definition of $\bar{F}$,
$\|\bar{F}(\theta_{1},\omega_{1},\mu_{1},V_{1})-\bar{F}(\theta_{2},\omega_{2},\mu_{2},V_{2})\|$
$=\|\mathbb{E}_{s\sim d_{\rho}^{\pi_{\theta_{1}},\,\phi_{\omega_{1}},\,\mu_{1}},a\sim\pi_{\theta_{1}}(\cdot\mid s),b\sim\phi_{\omega_{1}}(\cdot\mid s),s^{\prime}\sim{\cal P}^{\mu_{1}}(\cdot\mid s,a,b)}[F(\theta_{1},\mu_{1},V_{1},s,a,b,s^{\prime})]$
$\hskip 20.0pt-\mathbb{E}_{s\sim d_{\rho}^{\pi_{\theta_{2}},\,\phi_{\omega_{2}},\,\mu_{2}},a\sim\pi_{\theta_{2}}(\cdot\mid s),b\sim\phi_{\omega_{2}}(\cdot\mid s),s^{\prime}\sim{\cal P}^{\mu_{2}}(\cdot\mid s,a,b)}[F(\theta_{2},\mu_{2},V_{2},s,a,b,s^{\prime})]\|$
$=\Big\|\sum_{s,a,b,s^{\prime}}\Big(d_{\rho}^{\pi_{\theta_{1}},\,\phi_{\omega_{1}}\,\mu_{1}}(s)\pi_{\theta_{1}}(a\mid s)\phi_{\omega_{1}}(b\mid s){\cal P}^{\mu_{1}}(s^{\prime}\mid s,a,b)$
$\hskip 50.0pt-d_{\rho}^{\pi_{\theta_{2}},\,\phi_{\omega_{2}}\,\mu_{2}}(s)\pi_{\theta_{2}}(a\mid s)\phi_{\omega_{2}}(b\mid s){\cal P}^{\mu_{2}}(s^{\prime}\mid s,a)\Big)F(\theta_{2},\mu_{2},V_{2},s,a,b,s^{\prime})$
$\hskip 20.0pt+\mathbb{E}_{s\sim d_{\rho}^{\pi_{\theta_{1}},\,\phi_{\omega_{1}},\,\mu_{1}},a\sim\pi_{\theta_{1}}(\cdot\mid s),b\sim\phi_{\omega_{1}}(\cdot\mid s),s^{\prime}\sim{\cal P}^{\mu_{1}}(\cdot\mid s,a,b)}[F(\theta_{1},\mu_{1},V_{1},s,a,b,s^{\prime})-F(\theta_{2},\mu_{2},V_{2},s,a,b,s^{\prime})]\Big\|$
$\leq\Big\|\mathbb{E}_{s\sim d_{\rho}^{\pi_{\theta_{1}},\,\phi_{\omega_{1}},\,\mu_{1}},a\sim\pi_{\theta_{1}}(\cdot\mid s),b\sim\phi_{\omega_{1}}(\cdot\mid s),s^{\prime}\sim{\cal P}^{\mu_{1}}(\cdot\mid s,a,b)}[F(\theta_{1},\mu_{1},V_{1},s,a,b,s^{\prime})-F(\theta_{2},\mu_{2},V_{2},s,a,b,s^{\prime})]\Big\|$
$\hskip 20.0pt+B_{F}\Big|\sum_{s,a,b,s^{\prime}}\big(d_{\rho}^{\pi_{\theta_{1}},\,\phi_{\omega_{1}}\,\mu_{1}}(s)\pi_{\theta_{1}}(a\mid s)\phi_{\omega_{1}}(b\mid s){\cal P}^{\mu_{1}}(s^{\prime}\mid s,a,b)-d_{\rho}^{\pi_{\theta_{2}},\,\phi_{\omega_{2}}\,\mu_{2}}(s)\pi_{\theta_{2}}(a\mid s)\phi_{\omega_{2}}(b\mid s){\cal P}^{\mu_{2}}(s^{\prime}\mid s,a)\big)\Big|$
$\leq\Big\|\mathbb{E}_{s\sim d_{\rho}^{\pi_{\theta_{1}},\,\phi_{\omega_{1}},\,\mu_{1}},a\sim\pi_{\theta_{1}}(\cdot\mid s),b\sim\phi_{\omega_{1}}(\cdot\mid s),s^{\prime}\sim{\cal P}^{\mu_{1}}(\cdot\mid s,a,b)}[F(\theta_{1},\mu_{1},V_{1},s,a,b,s^{\prime})-F(\theta_{2},\mu_{2},V_{2},s,a,b,s^{\prime})]\Big\|$
$\hskip 20.0pt+B_{F}\Big|\sum_{s,s^{\prime}}\big(d_{\rho}^{\pi_{\theta_{1}},\,\phi_{\omega_{1}}\,\mu_{1}}(s)P^{\pi_{\theta_{1}},\,\phi_{\omega_{1}}\,\mu_{1}}(s^{\prime}\mid s)-d_{\rho}^{\pi_{\theta_{2}},\,\phi_{\omega_{2}}\,\mu_{2}}(s)P^{\pi_{\theta_{2}},\,\phi_{\omega_{2}}\,\mu_{2}}(s^{\prime}\mid s)\big)\Big|.$ (95)
To bound the first term of (95),
$\|F(\theta_{1},\mu_{1},V_{1},s,a,b,s^{\prime})-F(\theta_{2},\mu_{2},V_{2},s,a,b,s^{\prime})\|$
$=\|\nabla_{\theta}\log\pi_{\theta_{1}}(a\mid s)(r_{f}(s,a,b,\mu_{1})+\tau E(\pi_{\theta_{1}},s)+\gamma V_{1}(s^{\prime})-V_{1}(s))$
$\hskip 20.0pt-\nabla_{\theta}\log\pi_{\theta_{2}}(a\mid s)(r_{f}(s,a,b,\mu_{2})+\tau E(\pi_{\theta_{2}},s)+\gamma V_{2}(s^{\prime})-V_{2}(s))\|$
$\leq\|\nabla_{\theta}\log\pi_{\theta_{1}}(a\mid s)-\nabla_{\theta}\log\pi_{\theta_{2}}(a\mid s)\|\Big|r_{f}(s,a,b,\mu_{1})+\tau E(\pi_{\theta_{1}},s)+\gamma V_{1}(s^{\prime})-V_{1}(s)\Big|$
$\hskip 20.0pt+\|\nabla_{\theta}\log\pi_{\theta_{2}}(a\mid s)\|\Big|r_{f}(s,a,b,\mu_{1})-r_{f}(s,a,b,\mu_{2})+\tau E(\pi_{\theta_{1}},s)-\tau E(\pi_{\theta_{2}},s)$
$\hskip 150.0pt+\gamma V_{1}(s^{\prime})-\gamma V_{2}(s^{\prime})-V_{1}(s)+V_{2}(s))\Big|$
$\leq\|\theta_{1}-\theta_{2}\|(1+\tau\log|{\cal A}|+(1+\gamma)B_{V})$
$\hskip 20.0pt+2\Big(L_{r}|\mu_{1}-\mu_{2}|+\frac{(4+8\log|{\cal A}|)\tau}{(1-\gamma)^{3}}\|\theta_{1}-\theta_{2}\|+(1+\gamma)\|V_{1}-V_{2}\|\Big)$
$\leq\left(1+\tau\log|{\cal A}|+(1+\gamma)B_{V}+\frac{(4+8\log|{\cal A}|)\tau}{(1-\gamma)^{3}}\right)\|\theta_{1}-\theta_{2}\|+2L_{r}\|\mu_{1}-\mu_{2}\|+2(1+\gamma)\|V_{1}-V_{2}\|,$ (96)
where the second inequality is due to Lipschitz continuity of the entropy function (see Zeng et al. [2022][Lemma 6]) and the boundedness and $1$-Lipschitz continuity of $\nabla_{\theta}\log\pi_{\theta}$, which is obvious from (93) and (94).
By Lemma 2, we have for the second term of (95),
$B_{F}\Big|\sum_{s,s^{\prime}}\big(d_{\rho}^{\pi_{\theta_{1}},\,\phi_{\omega_{1}}\,\mu_{1}}(s)P^{\pi_{\theta_{1}},\,\phi_{\omega_{1}}\,\mu_{1}}(s^{\prime}\mid s)-d_{\rho}^{\pi_{\theta_{2}},\,\phi_{\omega_{2}}\,\mu_{2}}(s)P^{\pi_{\theta_{2}},\,\phi_{\omega_{2}}\,\mu_{2}}(s^{\prime}\mid s)\big)\Big|$
$\leq B_{F}\Big\|\sum_{s}\big(d_{\rho}^{\pi_{\theta_{1}},\,\phi_{\omega_{1}}\,\mu_{1}}(s)P^{\pi_{\theta_{1}},\,\phi_{\omega_{1}}\,\mu_{1}}(\cdot\mid s)-d_{\rho}^{\pi_{\theta_{2}},\,\phi_{\omega_{2}}\,\mu_{2}}(s)P^{\pi_{\theta_{2}},\,\phi_{\omega_{2}}\,\mu_{2}}(\cdot\mid s)\big)\Big\|$
$\leq\frac{B_{F}}{1-\gamma}\Big(L_{P}\|\mu_{1}-\mu_{2}\|+\|\theta_{1}-\theta_{2}\|+\|\omega_{1}-\omega_{2}\|\Big).$ (97)
Substituting (96) and (97) into (95), we have
$\|\bar{F}(\theta_{1},\omega_{1},\mu_{1},V_{1})-\bar{F}(\theta_{2},\omega_{2},\mu_{2},V_{2})\|$
$\leq\left(1+\tau\log|{\cal A}|+(1+\gamma)B_{V}+\frac{(4+8\log|{\cal A}|)\tau}{(1-\gamma)^{3}}\right)\|\theta_{1}-\theta_{2}\|+2L_{r}\|\mu_{1}-\mu_{2}\|+2(1+\gamma)\|V_{1}-V_{2}\|$
$\hskip 20.0pt+\frac{B_{F}}{1-\gamma}\Big(L_{P}\|\mu_{1}-\mu_{2}\|+\|\theta_{1}-\theta_{2}\|+\|\omega_{1}-\omega_{2}\|\Big)$
$=\left(1+\tau\log|{\cal A}|+(1+\gamma)B_{V}+\frac{(4+8\log|{\cal A}|)\tau}{(1-\gamma)^{3}}+\frac{B_{F}}{1-\gamma}\right)\|\theta_{1}-\theta_{2}\|$
$\hskip 20.0pt+\left(2L_{r}+\frac{B_{F}L_{P}}{1-\gamma}\right)\|\mu_{1}-\mu_{2}\|+\frac{B_{F}}{1-\gamma}\|\omega_{1}-\omega_{2}\|+4\|V_{1}-V_{2}\|.$
This obviously leads to the claimed result by setting $L_{F}=\max\left\{1+\tau\log|{\cal A}|+(1+\gamma)B_{V}+\frac{(4+8\log|{\cal A}|)\tau}{(1-\gamma)^{3}}+\frac{B_{F}}{1-\gamma},2L_{r}+\frac{B_{F}L_{P}}{1-\gamma},\frac{B_{F}}{1-\gamma},4\right\}$.
The Lipschitz continuity of $\bar{D}$ can be shown using almost identical steps, which we will skip.
For $\bar{H}$, we can adopt a similar argument
$\|\bar{H}(\pi_{1},\omega_{1},\mu_{1})-\bar{H}(\pi_{2},\omega_{2},\mu_{2})\|$
$\leq\|\mathbb{E}_{\bar{s}\sim\nu^{\pi_{1},\,\phi_{\omega_{1}},\,\mu_{1}}}[H(\mu_{1},\bar{s})-H(\mu_{2},\bar{s})]\|+B_{H}\Big|\sum_{s}\big(d_{\rho}^{\pi_{\theta_{1}},\,\phi_{\omega_{1}}\,\mu_{1}}(s)-d_{\rho}^{\pi_{\theta_{2}},\,\phi_{\omega_{2}}\,\mu_{2}}(s)\big)\Big|$
$\leq\|\mu_{1}-\mu_{2}\|+\Big\|\sum_{s}\big(d_{\rho}^{\pi_{\theta_{1}},\,\phi_{\omega_{1}}\,\mu_{1}}(s)P^{\pi_{\theta_{1}},\,\phi_{\omega_{1}}\,\mu_{1}}(\cdot\mid s)-d_{\rho}^{\pi_{\theta_{2}},\,\phi_{\omega_{2}}\,\mu_{2}}(s)P^{\pi_{\theta_{2}},\,\phi_{\omega_{2}}\,\mu_{2}}(\cdot\mid s)\big)\Big\|$
$\leq\|\mu_{1}-\mu_{2}\|+\frac{1}{1-\gamma}\Big(L_{P}\|\mu_{1}-\mu_{2}\|+\|\theta_{1}-\theta_{2}\|+\|\omega_{1}-\omega_{2}\|\Big)$
$\leq L_{H}(\|\theta_{1}-\theta_{2}\|+\|\omega_{1}-\omega_{2}\|+\|\mu_{1}-\mu_{2}\|),$
with $L_{H}=\max\{\frac{L_{P}}{1-\gamma},1\}$.
For $\bar{G}_{f}$, we again split $\|\bar{G}_{f}(\theta_{1},\omega_{1},\mu_{1},V_{1})-\bar{G}_{f}(\theta_{2},\omega_{2},\mu_{2},V_{2})\|$ into two parts and invoke Lemma 2
$\|\bar{G}_{f}(\theta_{1},\omega_{1},\mu_{1},V_{1})-\bar{G}_{f}(\theta_{2},\omega_{2},\mu_{2},V_{2})\|$
$\leq\|\mathbb{E}_{s\sim d_{\rho}^{\pi_{\theta_{1}},\,\phi_{\omega_{1}}\,\mu_{1}},a\sim\pi_{\theta_{1}}(\cdot\mid s),b\sim\phi_{\omega_{1}}(\cdot\mid s),s^{\prime}\sim{\cal P}^{\mu_{1}}(\cdot\mid s,a,b)}[G_{f}(\theta_{1},\mu_{1},V_{1},s,a,b,s^{\prime})-G_{f}(\theta_{2},\mu_{2},V_{2},s,a,b,s^{\prime})]\|$
$\hskip 20.0pt+B_{G}\Big|\sum_{s,s^{\prime}}\big(d_{\rho}^{\pi_{\theta_{1}},\,\phi_{\omega_{1}}\,\mu_{1}}(s)P^{\pi_{\theta_{1}},\,\phi_{\omega_{1}}\,\mu_{1}}(s^{\prime}\mid s)-d_{\rho}^{\pi_{\theta_{2}},\,\phi_{\omega_{2}}\,\mu_{2}}(s)P^{\pi_{\theta_{2}},\,\phi_{\omega_{2}}\,\mu_{2}}(s^{\prime}\mid s)\big)\Big|$
$\leq\Big|r_{f}(s,a,b,\mu_{1})-r_{f}(s,a,b,\mu_{2})+\tau E(\pi_{\theta_{1}},s)-\tau E(\pi_{\theta_{2}},s)+\gamma V_{1}(s^{\prime})-\gamma V_{2}(s^{\prime})-V_{1}(s)+V_{2}(s))\Big|$
$\hskip 20.0pt+\frac{B_{G}}{1-\gamma}\Big(L_{P}\|\mu_{1}-\mu_{2}\|+\|\theta_{1}-\theta_{2}\|+\|\omega_{1}-\omega_{2}\|\Big)$
$\leq\Big(\frac{(4+8\log|{\cal A}|)\tau}{(1-\gamma)^{3}}+\frac{B_{G}}{1-\gamma}\Big)\|\theta_{1}-\theta_{2}\|+(L_{r}+\frac{B_{G}L_{P}}{1-\gamma})\|\mu_{1}-\mu_{2}\|+\frac{B_{G}}{1-\gamma}\|\omega_{1}-\omega_{2}\|+2\|V_{1}-V_{2}\|.$
Similarly,
$\|\bar{G}_{l}(\omega_{1},\mu_{1},V_{1})-\bar{G}_{l}(\omega_{2},\mu_{2},V_{2})\|$
$\leq\|\mathbb{E}_{s\sim d_{\rho}^{\pi_{\theta_{1}},\,\phi_{\omega_{1}}\,\mu_{1}},a\sim\pi_{\theta_{1}}(\cdot\mid s),b\sim\phi_{\omega_{1}}(\cdot\mid s),s^{\prime}\sim{\cal P}^{\mu_{1}}(\cdot\mid s,a,b)}[G_{l}(\mu_{1},V_{1},s,a,b,s^{\prime})-G_{l}(\mu_{2},V_{2},s,b,s^{\prime})]\|$
$\hskip 20.0pt+B_{G}\Big|\sum_{s,s^{\prime}}\big(d_{\rho}^{\pi_{\theta_{1}},\,\phi_{\omega_{1}}\,\mu_{1}}(s)P^{\pi_{\theta_{1}},\,\phi_{\omega_{1}}\,\mu_{1}}(s^{\prime}\mid s)-d_{\rho}^{\pi_{\theta_{2}},\,\phi_{\omega_{2}}\,\mu_{2}}(s)P^{\pi_{\theta_{2}},\,\phi_{\omega_{2}}\,\mu_{2}}(s^{\prime}\mid s)\big)\Big|$
$\leq\Big|r_{l}(s,b,\mu_{1})-r_{l}(s,b,\mu_{2})+\gamma V_{1}(s^{\prime})-\gamma V_{2}(s^{\prime})-V_{1}(s)+V_{2}(s))\Big|$
$\hskip 20.0pt+\frac{B_{G}}{1-\gamma}\Big(L_{P}\|\mu_{1}-\mu_{2}\|+\|\theta_{1}-\theta_{2}\|+\|\omega_{1}-\omega_{2}\|\Big)$
$\leq\frac{B_{G}}{1-\gamma}\|\theta_{1}-\theta_{2}\|+(L_{r}+\frac{B_{G}L_{P}}{1-\gamma})\|\mu_{1}-\mu_{2}\|+\frac{B_{G}}{1-\gamma}\|\omega_{1}-\omega_{2}\|+2\|V_{1}-V_{2}\|.$
$\blacksquare$
D.5 Proof of Lemma 5
Within the proof of Lemma 5, we use the shorthand notation
$y_{k}=\hat{\mu}_{k}-\mu^{\star}(\phi_{\omega_{k}})+\xi_{k}\bar{H}(\pi^{\star}(\phi_{\omega_{k}},\hat{\mu}_{k}),\omega_{k},\hat{\mu}_{k}).$
As the operator $\mu^{\star}$ is Lipschitz, we know from the mean-value theorem that there exists $\omega_{k+1}^{z}=z\omega_{k}+(1-z)\omega_{k+1}$ for some scalar $z\in[0,1]$ such that
$\mu^{\star}(\phi_{\omega_{k+1}})-\mu^{\star}(\phi_{\omega_{k}})$ $=\nabla_{\omega}\mu^{\star}(\phi_{\omega_{k+1}^{z}})^{\top}(\omega_{k+1}-\omega_{k})$
$=\zeta_{k}\nabla_{\omega}\mu^{\star}(\phi_{\omega_{k+1}^{z}})^{\top}D(\omega_{k},\hat{\mu}_{k},\hat{V}_{l,k},s_{k},b_{k},s_{k}^{\prime})$
$=\zeta_{k}\nabla_{\omega}\mu^{\star}(\phi_{\omega_{k+1}^{z}})^{\top}\bar{D}(\omega_{k},\hat{\mu}_{k},\hat{V}_{l,k})$
$\hskip 20.0pt+\zeta_{k}\nabla_{\omega}\mu^{\star}(\phi_{\omega_{k+1}^{z}})^{\top}\Big(D(\omega_{k},\hat{\mu}_{k},\hat{V}_{l,k},s_{k},b_{k},s_{k}^{\prime})-\bar{D}(\omega_{k},\hat{\mu}_{k},\hat{V}_{l,k})\Big).$
This implies
$\mathbb{E}[\langle\hat{\mu}_{k}-\mu^{\star}(\phi_{\omega_{k}})+\xi_{k}\bar{H}(\pi^{\star}(\phi_{\omega_{k}},\hat{\mu}_{k}),\omega_{k},\hat{\mu}_{k}),\mu^{\star}(\phi_{\omega_{k+1}})-\mu^{\star}(\phi_{\omega_{k}})\rangle]$
$=\zeta_{k}\mathbb{E}[\langle y_{k},\nabla_{\omega}\mu^{\star}(\phi_{\omega_{k+1}^{z}})^{\top}\bar{D}(\omega_{k},\hat{\mu}_{k},\hat{V}_{l,k})\rangle]$
$\hskip 20.0pt+\zeta_{k}\mathbb{E}[\langle y_{k},\nabla_{\omega}\mu^{\star}(\phi_{\omega_{k+1}^{z}})^{\top}\Big(D(\omega_{k},\hat{\mu}_{k},\hat{V}_{l,k},s_{k},b_{k},s_{k}^{\prime})-\bar{D}(\omega_{k},\hat{\mu}_{k},\hat{V}_{l,k})\Big)\rangle]$
$=\zeta_{k}\mathbb{E}[\langle y_{k},\nabla_{\omega}\mu^{\star}(\phi_{\omega_{k+1}^{z}})^{\top}\bar{D}(\omega_{k},\hat{\mu}_{k},\hat{V}_{l,k})\rangle]$
$\hskip 20.0pt+\zeta_{k}\mathbb{E}[\langle y_{k},\Big(\nabla_{\omega}\mu^{\star}(\phi_{\omega_{k+1}^{z}})-\nabla_{\omega}\mu^{\star}(\phi_{\omega_{k}})\Big)^{\top}\Big(D(\omega_{k},\hat{\mu}_{k},\hat{V}_{l,k},s_{k},b_{k},s_{k}^{\prime})-\bar{D}(\omega_{k},\hat{\mu}_{k},\hat{V}_{l,k})\Big)\rangle]$
$\hskip 20.0pt+\zeta_{k}\mathbb{E}[\langle y_{k},\nabla_{\omega}\mu^{\star}(\phi_{\omega_{k}})^{\top}\mathbb{E}[D(\omega_{k},\hat{\mu}_{k},\hat{V}_{l,k},s_{k},b_{k},s_{k}^{\prime})-\bar{D}(\omega_{k},\hat{\mu}_{k},\hat{V}_{l,k})\mid{\cal F}_{k-1}]\rangle]$
$=\zeta_{k}\mathbb{E}[\langle y_{k},\nabla_{\omega}\mu^{\star}(\phi_{\omega_{k+1}^{z}})^{\top}\bar{D}(\omega_{k},\hat{\mu}_{k},\hat{V}_{l,k})\rangle]$
$\hskip 20.0pt+\zeta_{k}\mathbb{E}[\langle y_{k},\Big(\nabla_{\omega}\mu^{\star}(\phi_{\omega_{k+1}^{z}})-\nabla_{\omega}\mu^{\star}(\phi_{\omega_{k}})\Big)^{\top}\Big(D(\omega_{k},\hat{\mu}_{k},\hat{V}_{l,k},s_{k},b_{k},s_{k}^{\prime})-\bar{D}(\omega_{k},\hat{\mu}_{k},\hat{V}_{l,k})\Big)\rangle].$ (98)
To bound the first term of (98), note that by Assumption 2 and the fact that the softmax operator is $1$-Lipschitz
$\zeta_{k}\langle y_{k},\nabla_{\omega}\mu^{\star}(\phi_{\omega_{k+1}^{z}})^{\top}\bar{D}(\omega_{k},\hat{\mu}_{k},\hat{V}_{l,k})\rangle$
$\leq L\zeta_{k}\|y_{k}\|\|\bar{D}(\omega_{k},\hat{\mu}_{k},\hat{V}_{l,k})\|$
$\leq\frac{(1-\delta)\xi_{k}}{16}\|y_{k}\|^{2}+\frac{4L^{2}\zeta_{k}^{2}}{(1-\delta)\xi_{k}}\|\bar{D}(\omega_{k},\hat{\mu}_{k},\hat{V}_{l,k})\|^{2}$
$\leq\frac{(1-\delta)\xi_{k}}{16}\|y_{k}\|^{2}+\frac{8L^{2}\zeta_{k}^{2}}{(1-\delta)\xi_{k}}\|\bar{D}(\omega_{k},\mu^{\star}(\phi_{\omega_{k}}),V_{l}^{\phi_{\omega_{k}},\mu^{\star}(\phi_{\omega_{k}})})\|^{2}$
$\hskip 20.0pt+\frac{8L^{2}\zeta_{k}^{2}}{(1-\delta)\xi_{k}}\|\bar{D}(\omega_{k},\hat{\mu}_{k},\hat{V}_{l,k})-\bar{D}(\omega_{k},\mu^{\star}(\phi_{\omega_{k}}),V_{l}^{\phi_{\omega_{k}},\mu^{\star}(\phi_{\omega_{k}})})\|^{2}$
$\leq\frac{(1-\delta)\xi_{k}}{16}\|y_{k}\|^{2}+\frac{8L^{2}\zeta_{k}^{2}}{(1-\delta)\xi_{k}}\|\nabla_{\omega}J_{l}(\phi_{\omega_{k}},\mu)\mid_{\mu=\mu^{\star}(\phi_{\omega_{k}})}\|^{2}$
$\hskip 20.0pt+\frac{16L^{2}L_{D}^{2}\zeta_{k}^{2}}{(1-\delta)\xi_{k}}\|\hat{\mu}_{k}-\mu^{\star}(\phi_{\omega_{k}})\|^{2}+\frac{16L^{2}L_{D}^{2}\zeta_{k}^{2}}{(1-\delta)\xi_{k}}\|\hat{V}_{l,k}-V_{l}^{\phi_{\omega_{k}},\mu^{\star}(\phi_{\omega_{k}})}\|^{2}$
$\leq\frac{(1-\delta)\xi_{k}}{16}\|y_{k}\|^{2}+\frac{8L^{2}\eta_{2}^{2}\zeta_{k}^{2}}{(1-\delta)\xi_{k}}\|\nabla_{\omega}J_{l}(\phi_{\omega_{k}},\mu^{\star}(\phi_{\omega_{k}}))\|^{2}+\frac{16L^{2}L_{D}^{2}\zeta_{k}^{2}}{(1-\delta)\xi_{k}}\|\hat{\mu}_{k}-\mu^{\star}(\phi_{\omega_{k}})\|^{2}$
$\hskip 20.0pt+\frac{32L^{2}L_{D}^{2}\zeta_{k}^{2}}{(1-\delta)\xi_{k}}\|V_{l}^{\phi_{\omega_{k}},\hat{\mu}_{k}}-V_{l}^{\phi_{\omega_{k}},\mu^{\star}(\phi_{\omega_{k}})}\|^{2}+\frac{32L^{2}L_{D}^{2}\zeta_{k}^{2}}{(1-\delta)\xi_{k}}\|\hat{V}_{l,k}-V_{l}^{\phi_{\omega_{k}},\hat{\mu}_{k}}\|^{2}$
$\leq\frac{(1-\delta)\xi_{k}}{16}\|y_{k}\|^{2}+\frac{8L^{2}\eta_{2}^{2}\zeta_{k}^{2}}{(1-\delta)\xi_{k}}\varepsilon_{k}^{\phi}+\frac{48L^{2}L_{V}^{2}L_{D}^{2}\zeta_{k}^{2}}{(1-\delta)\xi_{k}}\varepsilon_{k}^{\mu}+\frac{32L^{2}L_{D}^{2}\zeta_{k}^{2}}{(1-\delta)\xi_{k}}\varepsilon_{l,k}^{V},$ (99)
where the fifth inequality follows from Assumption 6.
To bound the second term of (98),
$\zeta_{k}\mathbb{E}[\langle y_{k},\Big(\nabla_{\omega}\mu^{\star}(\phi_{\omega_{k+1}^{z}})-\nabla_{\omega}\mu^{\star}(\phi_{\omega_{k}})\Big)^{\top}\Big(D(\omega_{k},\hat{\mu}_{k},\hat{V}_{l,k},s_{k},b_{k},s_{k}^{\prime})-\bar{D}(\omega_{k},\hat{\mu}_{k},\hat{V}_{l,k})\Big)\rangle]$
$\leq 2B_{D}\zeta_{k}\mathbb{E}[\|y_{k}\|\|\nabla_{\omega}\mu^{\star}(\phi_{\omega_{k+1}^{z}})-\nabla_{\omega}\mu^{\star}(\phi_{\omega_{k}})\|]$
$\leq 2B_{D}L\zeta_{k}\mathbb{E}[\|y_{k}\|\|\phi_{\omega^{z}_{k+1}}-\phi_{\omega_{k}}\|]$
$\leq 2B_{D}L\zeta_{k}\mathbb{E}[\|y_{k}\|\|\phi_{\omega_{k+1}}-\phi_{\omega_{k}}\|]$
$\leq 2B_{D}^{2}L\zeta_{k}^{2}\mathbb{E}[\|y_{k}\|]$
$\leq B_{D}^{2}L\zeta_{k}^{2}\mathbb{E}[\|y_{k}\|^{2}]+B_{D}^{2}L\zeta_{k}^{2}.$ (100)
Substituting (99) and (100) into (98),
$\mathbb{E}[\langle\hat{\mu}_{k}-\mu^{\star}(\phi_{\omega_{k}})+\xi_{k}\bar{H}(\pi^{\star}(\phi_{\omega_{k}},\hat{\mu}_{k}),\omega_{k},\hat{\mu}_{k}),\mu^{\star}(\phi_{\omega_{k+1}})-\mu^{\star}(\phi_{\omega_{k}})\rangle]$
$\leq\frac{(1-\delta)\xi_{k}}{16}\mathbb{E}[\|y_{k}\|^{2}]+\frac{8L^{2}\eta_{2}^{2}\zeta_{k}^{2}}{(1-\delta)\xi_{k}}\mathbb{E}[\varepsilon_{k}^{\phi}]+\frac{48L^{2}L_{V}^{2}L_{D}^{2}\zeta_{k}^{2}}{(1-\delta)\xi_{k}}\mathbb{E}[\varepsilon_{k}^{\mu}]+\frac{32L^{2}L_{D}^{2}\zeta_{k}^{2}}{(1-\delta)\xi_{k}}\mathbb{E}[\varepsilon_{l,k}^{V}]$
$\hskip 20.0pt+B_{D}^{2}L\zeta_{k}^{2}\mathbb{E}[\|y_{k}\|^{2}]+B_{D}^{2}L\zeta_{k}^{2}$
$\leq\frac{(1-\delta)\xi_{k}}{8}\mathbb{E}[\|y_{k}\|^{2}]+\frac{8L^{2}\eta_{2}^{2}\zeta_{k}^{2}}{(1-\delta)\xi_{k}}\mathbb{E}[\varepsilon_{k}^{\phi}]+\frac{48L^{2}L_{V}^{2}L_{D}^{2}\zeta_{k}^{2}}{(1-\delta)\xi_{k}}\mathbb{E}[\varepsilon_{k}^{\mu}]+\frac{32L^{2}L_{D}^{2}\zeta_{k}^{2}}{(1-\delta)\xi_{k}}\mathbb{E}[\varepsilon_{l,k}^{V}]+B_{D}^{2}L\zeta_{k}^{2},$
where the last inequality follows from the step size conditions $\zeta_{k}\leq 1$ and $\frac{\zeta_{k}}{\xi_{k}}\leq\frac{1-\delta}{16B_{D}^{2}L}$.
$\blacksquare$
D.6 Proof of Lemma 6
Within the proof of Lemma 6, we denote $x_{k}=(\phi_{\omega_{k}},\hat{\mu}_{k})$. By the law of total expectation,
$-\mathbb{E}[\langle\nabla_{x}J_{f}(\pi_{\theta_{k}},\phi_{\omega_{k}},\hat{\mu}_{k})-\nabla_{x}J_{f}(\pi,\phi_{\omega_{k}},\hat{\mu}_{k})\mid_{\pi=\pi^{\star}(\phi_{\omega_{k}},\hat{\mu}_{k})},x_{k+1}-x_{k}\rangle]$
$=-\mathbb{E}[\langle\nabla_{x}J_{f}(\pi_{\theta_{k}},\phi_{\omega_{k}},\hat{\mu}_{k})-\nabla_{x}J_{f}(\pi,\phi_{\omega_{k}},\hat{\mu}_{k})\mid_{\pi=\pi^{\star}(\phi_{\omega_{k}},\hat{\mu}_{k})},\mathbb{E}\left[\left.\begin{array}[]{c}\zeta_{k}D(\omega_{k},\hat{\mu}_{k},\hat{V}_{l,k},s_{k},b_{k},s_{k}^{\prime})\\ \xi_{k}H(\hat{\mu}_{k},\bar{s}_{k})\end{array}\right|{\cal F}_{k-1}\right]\rangle]$ (103)
$=\mathbb{E}[\langle\nabla_{x}J_{f}(\pi_{\theta_{k}},\phi_{\omega_{k}},\hat{\mu}_{k})-\nabla_{x}J_{f}(\pi,\phi_{\omega_{k}},\hat{\mu}_{k})\mid_{\pi=\pi^{\star}(\phi_{\omega_{k}},\hat{\mu}_{k})},\left[\begin{array}[]{c}\zeta_{k}\bar{D}(\omega_{k},\hat{\mu}_{k},\hat{V}_{l,k})\\ \xi_{k}H(\pi_{\theta_{k}},\omega_{k},\hat{\mu}_{k})\end{array}\right]\rangle]$ (106)
$\leq\frac{(1-\gamma)\tau^{2}\rho_{\min}^{3}p_{\min}^{2}\alpha_{k}}{4|{\cal S}|L_{V}^{2}}\mathbb{E}[\|\nabla_{x}J_{f}(\pi_{\theta_{k}},\phi_{\omega_{k}},\hat{\mu}_{k})-\nabla_{x}J_{f}(\pi,\phi_{\omega_{k}},\hat{\mu}_{k})\mid_{\pi=\pi^{\star}(\phi_{\omega_{k}},\hat{\mu}_{k})}\|^{2}]$
$\hskip 20.0pt+\frac{|{\cal S}|L_{V}^{2}\zeta_{k}^{2}}{(1-\gamma)\tau^{2}\rho_{\min}^{3}p_{\min}^{2}\alpha_{k}}\mathbb{E}[\|\bar{D}(\omega_{k},\hat{\mu}_{k},\hat{V}_{l,k})\|^{2}]+\frac{|{\cal S}|L_{V}^{2}\xi_{k}^{2}}{(1-\gamma)\tau^{2}\rho_{\min}^{3}p_{\min}^{2}\alpha_{k}}\mathbb{E}[\|\bar{H}(\pi_{\theta_{k}},\omega_{k},\hat{\mu}_{k})\|^{2}]$
$\leq\frac{(1-\gamma)\tau^{2}\rho_{\min}^{3}p_{\min}^{2}\alpha_{k}}{4|{\cal S}|}\mathbb{E}[\|\pi^{\star}(\phi_{\omega_{k}},\hat{\mu}_{k})-\pi_{\theta_{k}}\|^{2}]$
$\hskip 20.0pt+\frac{2|{\cal S}|L_{V}^{2}\zeta_{k}^{2}}{(1-\gamma)\tau^{2}\rho_{\min}^{3}p_{\min}^{2}\alpha_{k}}\mathbb{E}[\|\bar{D}(\omega_{k},\mu^{\star}(\phi_{\omega_{k}}),V_{l}^{\phi_{\omega_{k}},\mu^{\star}(\phi_{\omega_{k}})})\|^{2}]$
$\hskip 20.0pt+\frac{2|{\cal S}|L_{V}^{2}\zeta_{k}^{2}}{(1-\gamma)\tau^{2}\rho_{\min}^{3}p_{\min}^{2}\alpha_{k}}\mathbb{E}[\|\bar{D}(\omega_{k},\hat{\mu}_{k},\hat{V}_{l,k})-\bar{D}(\omega_{k},\mu^{\star}(\phi_{\omega_{k}}),V_{l}^{\phi_{\omega_{k}},\mu^{\star}(\phi_{\omega_{k}})})\|^{2}]$
$\hskip 20.0pt+\frac{2|{\cal S}|L_{V}^{2}\zeta_{k}^{2}}{(1-\gamma)\tau^{2}\rho_{\min}^{3}p_{\min}^{2}\alpha_{k}}\mathbb{E}[\|\bar{H}(\pi^{\star}(\phi_{\omega_{k}},\hat{\mu}_{k}),\omega_{k},\hat{\mu}_{k})-\bar{H}(\pi^{\star}(\phi_{\omega_{k}}),\omega_{k},\mu^{\star}(\phi_{\omega_{k}}))\|^{2}]$
$\hskip 20.0pt+\frac{2|{\cal S}|L_{V}^{2}\zeta_{k}^{2}}{(1-\gamma)\tau^{2}\rho_{\min}^{3}p_{\min}^{2}\alpha_{k}}\mathbb{E}[\|\bar{H}(\pi_{\theta_{k}},\omega_{k},\hat{\mu}_{k})-\bar{H}(\pi^{\star}(\phi_{\omega_{k}},\hat{\mu}_{k}),\omega_{k},\hat{\mu}_{k})\|^{2}]$
$\leq\frac{(1-\gamma)\tau^{2}\rho_{\min}^{3}p_{\min}^{2}\alpha_{k}}{4|{\cal S}|}\mathbb{E}[\|\pi^{\star}(\phi_{\omega_{k}},\hat{\mu}_{k})-\pi_{\theta_{k}}\|^{2}]$
$\hskip 20.0pt+\frac{2|{\cal S}|L_{V}^{2}\zeta_{k}^{2}}{(1-\gamma)\tau^{2}\rho_{\min}^{3}p_{\min}^{2}\alpha_{k}}\mathbb{E}[\|\nabla_{\omega}J_{l}(\phi_{\omega_{k}},\mu)\mid_{\mu=\mu^{\star}(\phi_{\omega_{k}})}\|^{2}]$
$\hskip 20.0pt+\frac{4|{\cal S}|L_{V}^{2}L_{D}^{2}\zeta_{k}^{2}}{(1-\gamma)\tau^{2}\rho_{\min}^{3}p_{\min}^{2}\alpha_{k}}\mathbb{E}[\|\hat{\mu}_{k}-\mu^{\star}(\phi_{\omega_{k}})\|^{2}]+\frac{4|{\cal S}|L_{V}^{2}L_{D}^{2}\zeta_{k}^{2}}{(1-\gamma)\tau^{2}\rho_{\min}^{3}p_{\min}^{2}\alpha_{k}}\mathbb{E}[\|\hat{V}_{l,k}-V_{l}^{\phi_{\omega_{k}},\mu^{\star}(\phi_{\omega_{k}})}\|^{2}]$
$\hskip 20.0pt+\frac{2|{\cal S}|L_{V}^{2}L_{D}^{2}\xi_{k}^{2}}{(1-\gamma)\tau^{2}\rho_{\min}^{3}p_{\min}^{2}\alpha_{k}}\cdot L_{H}^{2}\mathbb{E}[\|\hat{\mu}_{k}-\mu^{\star}(\phi_{\omega_{k}})\|^{2}+L^{2}\|\hat{\mu}_{k}-\mu^{\star}\left(\phi_{\omega_{k}}\right)\|^{2}]$
$\hskip 20.0pt+\frac{2|{\cal S}|L_{V}^{2}L_{D}^{2}\xi_{k}^{2}}{(1-\gamma)\tau^{2}\rho_{\min}^{3}p_{\min}^{2}\alpha_{k}}\cdot L_{H}^{2}\mathbb{E}[\|\pi^{\star}(\phi_{\omega_{k},\hat{\mu}_{k}})-\pi_{\theta_{k}}\|^{2}]$
$\leq\frac{(1-\gamma)\tau^{2}\rho_{\min}^{3}p_{\min}^{2}\alpha_{k}}{4|{\cal S}|}\mathbb{E}[\|\pi^{\star}(\phi_{\omega_{k}},\hat{\mu}_{k})-\pi_{\theta_{k}}\|^{2}]$
$\hskip 20.0pt+\frac{2|{\cal S}|L_{V}^{2}\eta_{2}^{2}\zeta_{k}^{2}}{(1-\gamma)\tau^{2}\rho_{\min}^{3}p_{\min}^{2}\alpha_{k}}\mathbb{E}[\|\nabla_{\omega}J_{l}(\phi_{\omega_{k}},\mu^{\star}(\phi_{\omega_{k}}))\|^{2}]+\frac{4|{\cal S}|L_{V}^{2}L_{D}^{2}\zeta_{k}^{2}}{(1-\gamma)\tau^{2}\rho_{\min}^{3}p_{\min}^{2}\alpha_{k}}\mathbb{E}[\|\hat{\mu}_{k}-\mu^{\star}(\phi_{\omega_{k}})\|^{2}]$
$\hskip 20.0pt+\frac{8|{\cal S}|L_{V}^{2}L_{D}^{2}\zeta_{k}^{2}}{(1-\gamma)\tau^{2}\rho_{\min}^{3}p_{\min}^{2}\alpha_{k}}\mathbb{E}[\|V_{l}^{\phi_{\omega_{k}},\hat{\mu}_{k}}-V_{l}^{\phi_{\omega_{k}},\mu^{\star}(\phi_{\omega_{k}})}\|^{2}]+\frac{8|{\cal S}|L_{V}^{2}L_{D}^{2}\zeta_{k}^{2}}{(1-\gamma)\tau^{2}\rho_{\min}^{3}p_{\min}^{2}\alpha_{k}}\mathbb{E}[\|\hat{V}_{l,k}-V_{l}^{\phi_{\omega_{k}},\hat{\mu}_{k}}\|^{2}]$
$\hskip 20.0pt+\frac{4|{\cal S}|L^{2}L_{V}^{2}L_{D}^{2}L_{H}^{2}\xi_{k}^{2}}{(1-\gamma)\tau^{2}\rho_{\min}^{3}p_{\min}^{2}\alpha_{k}}\mathbb{E}[\|\hat{\mu}_{k}-\mu^{\star}(\phi_{\omega_{k}})\|^{2}]+\frac{2|{\cal S}|L_{V}^{2}L_{D}^{2}L_{H}^{2}\xi_{k}^{2}}{(1-\gamma)\tau^{2}\rho_{\min}^{3}p_{\min}^{2}\alpha_{k}}\mathbb{E}[\|\pi^{\star}(\phi_{\omega_{k},\hat{\mu}_{k}})-\pi_{\theta_{k}}\|^{2}]$
$\leq\frac{(1-\gamma)\tau^{2}\rho_{\min}^{3}p_{\min}^{2}\alpha_{k}}{4|{\cal S}|}\mathbb{E}[\|\pi^{\star}(\phi_{\omega_{k}},\hat{\mu}_{k})-\pi_{\theta_{k}}\|^{2}]+\frac{2|{\cal S}|L_{V}^{2}L_{D}^{2}L_{H}^{2}\xi_{k}^{2}}{(1-\gamma)\tau^{2}\rho_{\min}^{3}p_{\min}^{2}\alpha_{k}}\mathbb{E}[\|\pi^{\star}(\phi_{\omega_{k},\hat{\mu}_{k}})-\pi_{\theta_{k}}\|^{2}]$
$\hskip 20.0pt+\frac{2|{\cal S}|L_{V}^{2}\eta_{2}^{2}\zeta_{k}^{2}}{(1-\gamma)\tau^{2}\rho_{\min}^{3}p_{\min}^{2}\alpha_{k}}\mathbb{E}[\varepsilon_{k}^{\phi}]+\frac{16|{\cal S}|L^{2}L_{V}^{4}L_{D}^{2}L_{H}^{2}\xi_{k}^{2}}{(1-\gamma)\tau^{2}\rho_{\min}^{3}p_{\min}^{2}\alpha_{k}}\mathbb{E}[\varepsilon_{k}^{\mu}]+\frac{8|{\cal S}|L_{V}^{2}\zeta_{k}^{2}}{(1-\gamma)\tau^{2}\rho_{\min}^{3}p_{\min}^{2}\alpha_{k}}\mathbb{E}[\varepsilon_{l,k}^{V}],$ (107)
where the second inequality follows from the fact that $\bar{H}(\pi^{\star}(\phi),\omega_{k},\mu^{\star}(\phi))=0$, and the last inequality simplifies and combines terms under the step size condition $\frac{\zeta_{k}}{\xi_{k}}\leq LL_{H}$.
$\blacksquare$
D.7 Proof of Lemma 7
Within the proof of this lemma, we employ the shorthand notation $x_{k}=[\theta_{k},\omega_{k},\hat{\mu}_{k}]$, $\ell(x_{k})=V_{f}^{\pi_{\theta_{k}},\phi_{\omega_{k}},\,\hat{\mu}_{k}}$, and
$y_{k}=\hat{V}_{f,k}-V_{f}^{\pi_{\theta_{k}},\phi_{\omega_{k}},\,\hat{\mu}_{k}}+\beta_{k}\bar{G}_{f}(\theta_{k},\omega_{k},\hat{\mu}_{k},\hat{V}_{f,k}).$
The value function is smooth, i.e. has Lipschitz gradients. As a result, we have from the mean-value theorem that there exists $x^{z}_{k+1}=zx_{k}+(1-z)x_{k+1}$ for some scalar $z\in[0,1]$ such that
$\ell(x_{k})-\ell(x_{k+1})$
$=\nabla_{x}\ell(x^{z}_{k+1})^{\top}\Big(x_{k}-x_{k+1}\Big)$
$=\Big(\nabla_{\theta}V_{f}^{\pi_{\theta^{z}_{k+1}},\phi_{\omega^{z}_{k+1}},\,\hat{\mu}^{z}_{k+1}}\Big)^{\top}\Big(\theta_{k}-\theta_{k+1}\Big)+\Big(V_{f}^{\pi_{\theta^{z}_{k+1}},\phi_{\omega^{z}_{k+1}},\,\hat{\mu}^{z}_{k+1}}\Big)^{\top}\Big(\omega_{k}-\omega_{k+1}\Big)$
$\hskip 20.0pt+\Big(V_{f}^{\pi_{\theta^{z}_{k+1}},\phi_{\omega^{z}_{k+1}},\,\hat{\mu}^{z}_{k+1}}\Big)^{\top}\Big(\hat{\mu}_{k}-\hat{\mu}_{k+1}\Big)$
$=\alpha_{k}\Big(\nabla_{\theta}V_{f}^{\pi_{\theta^{z}_{k+1}},\phi_{\omega^{z}_{k+1}},\,\hat{\mu}^{z}_{k+1}}\Big)^{\top}\bar{F}(\theta_{k},\omega_{k},\hat{\mu}_{k},\hat{V}_{f,k})$
$\hskip 20.0pt+\alpha_{k}\Big(\nabla_{\theta}V_{f}^{\pi_{\theta^{z}_{k+1}},\phi_{\omega^{z}_{k+1}},\,\hat{\mu}^{z}_{k+1}}\Big)^{\top}\Big(F(\theta_{k},\omega_{k},\hat{\mu}_{k},\hat{V}_{f,k},s_{k},a_{k},b_{k},s_{k}^{\prime})-\bar{F}(\theta_{k},\omega_{k},\hat{\mu}_{k},\hat{V}_{f,k})\Big)$
$\hskip 20.0pt+\zeta_{k}\Big(\nabla_{\omega}V_{f}^{\pi_{\theta^{z}_{k+1}},\phi_{\omega^{z}_{k+1}},\,\hat{\mu}^{z}_{k+1}}\Big)^{\top}\bar{D}(\omega_{k},\hat{\mu}_{k},\hat{V}_{l,k})$
$\hskip 20.0pt+\zeta_{k}\Big(\nabla_{\theta}V_{f}^{\pi_{\theta^{z}_{k+1}},\phi_{\omega^{z}_{k+1}},\,\hat{\mu}^{z}_{k+1}}\Big)^{\top}\Big(D(\omega_{k},\hat{\mu}_{k},\hat{V}_{l,k},s_{k},b_{k},s_{k}^{\prime})-\bar{D}(\omega_{k},\hat{\mu}_{k},\hat{V}_{l,k})\Big)$
$\hskip 20.0pt+\xi_{k}\Big(\nabla_{\mu}V_{f}^{\pi_{\theta^{z}_{k+1}},\phi_{\omega^{z}_{k+1}},\,\hat{\mu}^{z}_{k+1}}\Big)^{\top}\bar{H}(\pi_{\theta_{k}},\omega_{k},\hat{\mu}_{k})$
$\hskip 20.0pt+\xi_{k}\Big(\nabla_{\theta}V_{f}^{\pi_{\theta^{z}_{k+1}},\phi_{\omega^{z}_{k+1}},\,\hat{\mu}^{z}_{k+1}}\Big)^{\top}\Big(H(\hat{\mu}_{k},\bar{s}_{k})-\bar{H}(\pi_{\theta_{k}},\omega_{k},\hat{\mu}_{k})\Big),$ (108)
where we denote $\theta^{z}_{k+1}=z\theta_{k}+(1-z)\theta_{k+1},\omega^{z}_{k+1}=z\omega_{k}+(1-z)\omega_{k+1},\hat{\mu}^{z}_{k+1}=z\hat{\mu}_{k}+(1-z)\hat{\mu}_{k+1}$.
Plugging (108) into the cross term of interest, we have
$\langle\hat{V}_{f,k}-V_{f}^{\pi_{\theta_{k}},\phi_{\omega_{k}},\,\hat{\mu}_{k}}+\beta_{k}\bar{G}_{f}(\theta_{k},\omega_{k},\hat{\mu}_{k},\hat{V}_{f,k}),V_{f}^{\pi_{\theta_{k}},\phi_{\omega_{k}},\,\hat{\mu}_{k}}-V_{f}^{\pi_{\theta_{k+1}},\phi_{\omega_{k+1}},\,\hat{\mu}_{k+1}}\rangle$
$=\langle y_{k},\ell(x_{k})-\ell(x_{k+1})\rangle$
$=\alpha_{k}\langle y_{k},\Big(\nabla_{\theta}V_{f}^{\pi_{\theta^{z}_{k+1}},\phi_{\omega^{z}_{k+1}},\,\hat{\mu}^{z}_{k+1}}\Big)^{\top}\bar{F}(\theta_{k},\omega_{k},\hat{\mu}_{k},\hat{V}_{f,k})\rangle$
$\hskip 20.0pt+\alpha_{k}\langle y_{k},\Big(\nabla_{\theta}V_{f}^{\pi_{\theta^{z}_{k+1}},\phi_{\omega^{z}_{k+1}},\,\hat{\mu}^{z}_{k+1}}\Big)^{\top}\Big(F(\theta_{k},\omega_{k},\hat{\mu}_{k},\hat{V}_{f,k}),s_{k},a_{k},b_{k},s_{k}^{\prime})-\bar{F}(\theta_{k},\omega_{k},\hat{\mu}_{k},\hat{V}_{f,k})\Big)\rangle$
$\hskip 20.0pt+\zeta_{k}\langle y_{k},\Big(\nabla_{\omega}V_{f}^{\pi_{\theta^{z}_{k+1}},\phi_{\omega^{z}_{k+1}},\,\hat{\mu}^{z}_{k+1}}\Big)^{\top}\bar{D}(\omega_{k},\hat{\mu}_{k},\hat{V}_{l,k})\rangle$
$\hskip 20.0pt+\zeta_{k}\langle y_{k},\Big(\nabla_{\omega}V_{f}^{\pi_{\theta^{z}_{k+1}},\phi_{\omega^{z}_{k+1}},\,\hat{\mu}^{z}_{k+1}}\Big)^{\top}\Big(D(\omega_{k},\hat{\mu}_{k},\hat{V}_{l,k},s_{k},b_{k},s_{k}^{\prime})-\bar{D}(\omega_{k},\hat{\mu}_{k},\hat{V}_{l,k})\Big)\rangle$
$\hskip 20.0pt+\xi_{k}\langle y_{k},\Big(\nabla_{\mu}V_{f}^{\pi_{\theta^{z}_{k+1}},\phi_{\omega^{z}_{k+1}},\,\hat{\mu}^{z}_{k+1}}\Big)^{\top}\bar{H}(\pi_{\theta_{k}},\omega_{k},\hat{\mu}_{k})\rangle$
$\hskip 20.0pt+\xi_{k}\langle y_{k},\Big(\nabla_{\mu}V_{f}^{\pi_{\theta^{z}_{k+1}},\phi_{\omega^{z}_{k+1}},\,\hat{\mu}^{z}_{k+1}}\Big)^{\top}\Big(H(\hat{\mu}_{k},\bar{s}_{k})-\bar{H}(\pi_{\theta_{k}},\omega_{k},\hat{\mu}_{k})\Big)\rangle.$ (109)
We bound each term of (109) individually. First, by Young’s inequality
$\alpha_{k}\langle y_{k},\Big(\nabla_{\theta}V_{f}^{\pi_{\theta^{z}_{k+1}},\phi_{\omega^{z}_{k+1}},\,\hat{\mu}^{z}_{k+1}}\Big)^{\top}\bar{F}(\theta_{k},\omega_{k},\hat{\mu}_{k},\hat{V}_{f,k})\rangle$
$\leq L_{V}\alpha_{k}\|y_{k}\|\|\bar{F}(\theta_{k},\omega_{k},\hat{\mu}_{k},\hat{V}_{f,k})\|$
$\leq\frac{(1-\gamma)\beta_{k}}{12}\|y_{k}\|^{2}+\frac{3L_{V}^{2}\alpha_{k}^{2}}{(1-\gamma)\beta_{k}}\|\bar{F}(\theta_{k},\omega_{k},\hat{\mu}_{k},\hat{V}_{f,k})\|^{2}$
$\leq\frac{(1-\gamma)\beta_{k}}{12}\|y_{k}\|^{2}+\frac{6L_{V}^{2}\alpha_{k}^{2}}{(1-\gamma)\beta_{k}}\|\bar{F}(\theta_{k},\omega_{k},\hat{\mu}_{k},V_{f}^{\pi_{\theta_{k}},\phi_{\omega_{k}},\hat{\mu}_{k}})\|^{2}$
$\hskip 20.0pt+\frac{6L_{V}^{2}\alpha_{k}^{2}}{(1-\gamma)\beta_{k}}\|\bar{F}(\theta_{k},\omega_{k},\hat{\mu}_{k},\hat{V}_{f,k})-\bar{F}(\theta_{k},\omega_{k},\hat{\mu}_{k},V_{f}^{\pi_{\theta_{k}},\phi_{\omega_{k}},\hat{\mu}_{k}})\|^{2}$
$\leq\frac{(1-\gamma)\beta_{k}}{12}\|y_{k}\|^{2}+\frac{6L_{V}^{2}\alpha_{k}^{2}}{(1-\gamma)\beta_{k}}\|\nabla_{\theta}J_{f}(\pi_{\theta_{k}},\phi_{\omega_{k}},\hat{\mu}_{k})\|^{2}+\frac{6L_{V}^{4}\alpha_{k}^{2}}{(1-\gamma)\beta_{k}}\varepsilon_{f,k}^{V}.$ (110)
For the second term of (109), we take the expectation
$\alpha_{k}\mathbb{E}\left[\langle y_{k},\Big(\nabla_{\theta}V_{f}^{\pi_{\theta^{z}_{k+1}},\phi_{\omega^{z}_{k+1}},\,\hat{\mu}^{z}_{k+1}}\Big)^{\top}\Big(F(\theta_{k},\omega_{k},\hat{\mu}_{k},\hat{V}_{f,k},s_{k},a_{k},b_{k},s_{k}^{\prime})-\bar{F}(\theta_{k},\omega_{k},\hat{\mu}_{k},\hat{V}_{f,k})\Big)\rangle\right]$
$=\alpha_{k}\mathbb{E}[\langle y_{k},\Big(\nabla_{\theta}V_{f}^{\pi_{\theta^{z}_{k+1}},\phi_{\omega^{z}_{k+1}},\,\hat{\mu}^{z}_{k+1}}-\nabla_{\theta}V_{f}^{\pi_{\theta_{k}},\phi_{\omega_{k}},\,\hat{\mu}_{k}}\Big)^{\top}\Big(F(\theta_{k},\omega_{k},\hat{\mu}_{k},\hat{V}_{f,k},s_{k},a_{k},b_{k},s_{k}^{\prime})-\bar{F}(\theta_{k},\omega_{k},\hat{\mu}_{k},\hat{V}_{f,k})\Big)]$
$\hskip 20.0pt+\alpha_{k}\mathbb{E}[\langle y_{k},\Big(\nabla_{\theta}V_{f}^{\pi_{\theta_{k}},\phi_{\omega_{k}},\,\hat{\mu}_{k}}\Big)^{\top}\Big(F(\theta_{k},\omega_{k},\hat{\mu}_{k},\hat{V}_{f,k},s_{k},a_{k},b_{k},s_{k}^{\prime})-\bar{F}(\theta_{k},\omega_{k},\hat{\mu}_{k},\hat{V}_{f,k})\Big)]$
$=\alpha_{k}\mathbb{E}[\langle y_{k},\Big(\nabla_{\theta}V_{f}^{\pi_{\theta^{z}_{k+1}},\phi_{\omega^{z}_{k+1}},\,\hat{\mu}^{z}_{k+1}}-\nabla_{\theta}V_{f}^{\pi_{\theta_{k}},\phi_{\omega_{k}},\,\hat{\mu}_{k}}\Big)^{\top}\Big(F(\theta_{k},\omega_{k},\hat{\mu}_{k},\hat{V}_{f,k},s_{k},a_{k},b_{k},s_{k}^{\prime})-\bar{F}(\theta_{k},\omega_{k},\hat{\mu}_{k},\hat{V}_{f,k})\Big)]$
$\leq 2B_{F}\alpha_{k}\mathbb{E}[\|y_{k}\|\|\nabla_{\theta}V_{f}^{\pi_{\theta^{z}_{k+1}},\phi_{\omega^{z}_{k+1}},\,\hat{\mu}^{z}_{k+1}}-\nabla_{\theta}V_{f}^{\pi_{\theta_{k}},\phi_{\omega_{k}},\,\hat{\mu}_{k}}\|]$
$\leq 2B_{F}L_{VV}\alpha_{k}\mathbb{E}[\|y_{k}\|\Big(\|\pi_{\theta^{z}_{k+1}}-\pi_{\theta_{k}}\|+\|\phi_{\omega^{z}_{k+1}}-\phi_{\omega_{k}}\|+\|\hat{\mu}_{k+1}^{z}-\hat{\mu}_{k}\|\Big)]$
$\leq 2B_{F}L_{VV}\alpha_{k}\mathbb{E}[\|y_{k}\|\Big(\|\pi_{\theta_{k+1}}-\pi_{\theta_{k}}\|+\|\phi_{\omega_{k+1}}-\phi_{\omega_{k}}\|+\|\hat{\mu}_{k+1}-\hat{\mu}_{k}\|\Big)]$
$\leq 2B_{F}L_{VV}\alpha_{k}\mathbb{E}[\|y_{k}\|]\Big(\alpha_{k}B_{F}+\zeta_{k}B_{D}+\xi_{k}B_{H}\Big)$
$\leq 2B_{F}L_{VV}(B_{F}+B_{D}+B_{H})\alpha_{k}^{2}\mathbb{E}[\|y_{k}\|]$
$\leq B_{F}L_{VV}(B_{F}+B_{D}+B_{H})\alpha_{k}^{2}\mathbb{E}[\|y_{k}\|^{2}]+B_{F}L_{VV}(B_{F}+B_{D}+B_{H})\alpha_{k}^{2},$ (111)
where the fifth inequality follows from the step size condition $\zeta_{k}\leq\xi_{l}\leq\alpha_{k}$, and the second equation follows from
$\mathbb{E}[\langle y_{k},\Big(\nabla_{\theta}V_{f}^{\pi_{\theta_{k}},\phi_{\omega_{k}},\,\hat{\mu}_{k}}\Big)^{\top}\Big(F(\theta_{k},\omega_{k},\hat{\mu}_{k},\hat{V}_{f,k},s_{k},a_{k},b_{k},s_{k}^{\prime})-\bar{F}(\theta_{k},\omega_{k},\hat{\mu}_{k},\hat{V}_{f,k})\Big)]$
$=\mathbb{E}[\langle y_{k},\Big(\nabla_{\theta}V_{f}^{\pi_{\theta_{k}},\phi_{\omega_{k}},\,\hat{\mu}_{k}}\Big)^{\top}\mathbb{E}\Big[F(\theta_{k},\omega_{k},\hat{\mu}_{k},\hat{V}_{f,k},s_{k},a_{k},b_{k},s_{k}^{\prime})-\bar{F}(\theta_{k},\omega_{k},\hat{\mu}_{k},\hat{V}_{f,k})\mid{\cal F}_{k-1}\Big]]$
$=0.$
The third term of (109) can be bounded similar to the first term,
$\zeta_{k}\langle y_{k},\Big(\nabla_{\omega}V_{f}^{\pi_{\theta^{z}_{k+1}},\phi_{\omega^{z}_{k+1}},\,\hat{\mu}^{z}_{k+1}}\Big)^{\top}\bar{D}(\omega_{k},\hat{\mu}_{k},\hat{V}_{l,k})\rangle$
$\leq L_{V}\zeta_{k}\|y_{k}\|\|\bar{D}(\omega_{k},\hat{\mu}_{k},\hat{V}_{l,k})\|$
$\leq\frac{(1-\gamma)\beta_{k}}{12}\|y_{k}\|^{2}+\frac{3L_{V}^{2}\zeta_{k}^{2}}{(1-\gamma)\beta_{k}}\|\bar{D}(\omega_{k},\hat{\mu}_{k},\hat{V}_{l,k})\|^{2}$
$\leq\frac{(1-\gamma)\beta_{k}}{12}\|y_{k}\|^{2}+\frac{6L_{V}^{2}\zeta_{k}^{2}}{(1-\gamma)\beta_{k}}\|\bar{D}(\omega_{k},\mu^{\star}(\phi_{\omega_{k}}),V_{l}^{\phi_{\omega_{k}},\mu^{\star}(\phi_{\omega_{k}})})\|^{2}$
$\hskip 20.0pt+\frac{6L_{V}^{2}\zeta_{k}^{2}}{(1-\gamma)\beta_{k}}\|\bar{D}(\omega_{k},\hat{\mu}_{k},\hat{V}_{l,k})-\bar{D}(\omega_{k},\mu^{\star}(\phi_{\omega_{k}}),V_{l}^{\phi_{\omega_{k}},\mu^{\star}(\phi_{\omega_{k}})})\|^{2}$
$\leq\frac{(1-\gamma)\beta_{k}}{12}\|y_{k}\|^{2}+\frac{6L_{V}^{2}\zeta_{k}^{2}}{(1-\gamma)\beta_{k}}\|\nabla_{\omega}J_{l}(\phi_{\omega_{k}},\mu)\mid_{\mu=\mu^{\star}(\phi_{\omega_{k}})}\|^{2}$
$\hskip 20.0pt+\frac{12L_{V}^{2}L_{D}^{2}\zeta_{k}^{2}}{(1-\gamma)\beta_{k}}\|\hat{\mu}_{k}-\mu^{\star}(\phi_{\omega_{k}})\|^{2}+\frac{12L_{V}^{2}L_{D}^{2}\zeta_{k}^{2}}{(1-\gamma)\beta_{k}}\|\hat{V}_{l,k}-V_{l}^{\phi_{\omega_{k}},\mu^{\star}(\phi_{\omega_{k}})}\|^{2}$
$\leq\frac{(1-\gamma)\beta_{k}}{12}\|y_{k}\|^{2}+\frac{6L_{V}^{2}\eta_{2}^{2}\zeta_{k}^{2}}{(1-\gamma)\beta_{k}}\|\nabla_{\omega}J_{l}(\phi_{\omega_{k}},\mu^{\star}(\phi_{\omega_{k}}))\|^{2}+\frac{12L_{V}^{2}L_{D}^{2}\zeta_{k}^{2}}{(1-\gamma)\beta_{k}}\|\hat{\mu}_{k}-\mu^{\star}(\phi_{\omega_{k}})\|^{2}$
$\hskip 20.0pt+\frac{24L_{V}^{2}L_{D}^{2}\zeta_{k}^{2}}{(1-\gamma)\beta_{k}}\|V_{l}^{\phi_{\omega_{k}},\hat{\mu}_{k}}-V_{l}^{\phi_{\omega_{k}},\mu^{\star}(\phi_{\omega_{k}})}\|^{2}+\frac{24L_{V}^{2}L_{D}^{2}\zeta_{k}^{2}}{(1-\gamma)\beta_{k}}\|\hat{V}_{l,k}-V_{l}^{\phi_{\omega_{k}},\hat{\mu}_{k}}\|^{2}$
$\leq\frac{(1-\gamma)\beta_{k}}{12}\|y_{k}\|^{2}+\frac{6L_{V}^{2}\eta_{2}^{2}\zeta_{k}^{2}}{(1-\gamma)\beta_{k}}\varepsilon_{k}^{\phi}+\frac{36L_{V}^{4}L_{D}^{2}\zeta_{k}^{2}}{(1-\gamma)\beta_{k}}\varepsilon_{k}^{\mu}+\frac{24L_{V}^{2}L_{D}^{2}\zeta_{k}^{2}}{(1-\gamma)\beta_{k}}\varepsilon_{l,k}^{V},$ (112)
where the fifth inequality follows from Assumption 6.
For the fourth term of (109), we again take the expectation and use the technique in (111)
$\zeta_{k}\mathbb{E}\langle y_{k},\Big(\nabla_{\omega}V_{f}^{\pi_{\theta^{z}_{k+1}},\phi_{\omega^{z}_{k+1}},\,\hat{\mu}^{z}_{k+1}}\Big)^{\top}\Big(D(\omega_{k},\hat{\mu}_{k},\hat{V}_{f,k},s_{k},b_{k},s_{k}^{\prime})-\bar{D}(\omega_{k},\hat{\mu}_{k},\hat{V}_{f,k})\Big)\rangle$
$=\zeta_{k}\mathbb{E}[\langle y_{k},\Big(\nabla_{\theta}V_{f}^{\pi_{\theta^{z}_{k+1}},\phi_{\omega^{z}_{k+1}},\,\hat{\mu}^{z}_{k+1}}-\nabla_{\theta}V_{f}^{\pi_{\theta_{k}},\phi_{\omega_{k}},\,\hat{\mu}_{k}}\Big)^{\top}\Big(D(\omega_{k},\hat{\mu}_{k},\hat{V}_{l,k},s_{k},b_{k},s_{k}^{\prime})-\bar{D}(\omega_{k},\hat{\mu}_{k},\hat{V}_{l,k})\Big)]$
$\hskip 20.0pt+\zeta_{k}\mathbb{E}[\langle y_{k},\Big(\nabla_{\theta}V_{f}^{\pi_{\theta_{k}},\phi_{\omega_{k}},\,\hat{\mu}_{k}}\Big)^{\top}\Big(D(\omega_{k},\hat{\mu}_{k},\hat{V}_{l,k},s_{k},b_{k},s_{k}^{\prime})-\bar{D}(\omega_{k},\hat{\mu}_{k},\hat{V}_{l,k})\Big)]$
$=\zeta_{k}\mathbb{E}[\langle y_{k},\Big(\nabla_{\theta}V_{f}^{\pi_{\theta^{z}_{k+1}},\phi_{\omega^{z}_{k+1}},\,\hat{\mu}^{z}_{k+1}}-\nabla_{\theta}V_{f}^{\pi_{\theta_{k}},\phi_{\omega_{k}},\,\hat{\mu}_{k}}\Big)^{\top}\Big(D(\omega_{k},\hat{\mu}_{k},\hat{V}_{l,k},s_{k},b_{k},s_{k}^{\prime})-\bar{D}(\omega_{k},\hat{\mu}_{k},\hat{V}_{l,k})\Big)]$
$\leq 2B_{D}\zeta_{k}\mathbb{E}[\|y_{k}\|\|\nabla_{\theta}V_{f}^{\pi_{\theta^{z}_{k+1}},\phi_{\omega^{z}_{k+1}},\,\hat{\mu}^{z}_{k+1}}-\nabla_{\theta}V_{f}^{\pi_{\theta_{k}},\phi_{\omega_{k}},\,\hat{\mu}_{k}}\|]$
$\leq 2B_{D}L_{VV}\zeta_{k}\mathbb{E}[\|y_{k}\|\Big(\|\pi_{\theta^{z}_{k+1}}-\pi_{\theta_{k}}\|+\|\phi_{\omega^{z}_{k+1}}-\phi_{\omega_{k}}\|+\|\hat{\mu}_{k+1}^{z}-\hat{\mu}_{k}\|\Big)]$
$\leq 2B_{D}L_{VV}\zeta_{k}\mathbb{E}[\|y_{k}\|\Big(\|\pi_{\theta_{k+1}}-\pi_{\theta_{k}}\|+\|\phi_{\omega_{k+1}}-\phi_{\omega_{k}}\|+\|\hat{\mu}_{k+1}-\hat{\mu}_{k}\|\Big)]$
$\leq 2B_{D}L_{VV}\zeta_{k}\mathbb{E}[\|y_{k}\|]\Big(\alpha_{k}B_{F}+\zeta_{k}B_{D}+\xi_{k}B_{H}\Big)$
$\leq B_{D}L_{VV}(B_{F}+B_{D}+B_{H})\alpha_{k}^{2}\mathbb{E}[\|y_{k}\|^{2}]+B_{F}L_{VV}(B_{F}+B_{D}+B_{H})\alpha_{k}^{2}.$ (113)
For the fifth term of (109), recall the definition of $\pi^{\star}$ in (2) and (4)
$\xi_{k}\langle y_{k},\Big(\nabla_{\mu}V_{f}^{\pi_{\theta^{z}_{k+1}},\phi_{\omega^{z}_{k+1}},\,\hat{\mu}^{z}_{k+1}}\Big)^{\top}\bar{H}(\pi_{\theta_{k}},\omega_{k},\hat{\mu}_{k})\rangle$
$\leq L_{V}\xi_{k}\|y_{k}\|\|\bar{H}(\pi_{\theta_{k}},\omega_{k},\hat{\mu}_{k})\|$
$\leq\frac{(1-\gamma)\beta_{k}}{12}\|y_{k}\|^{2}+\frac{3L_{V}^{2}\xi_{k}^{2}}{(1-\gamma)\beta_{k}}\|\bar{H}(\pi_{\theta_{k}},\omega_{k},\hat{\mu}_{k})\|^{2}$
$\leq\frac{(1-\gamma)\beta_{k}}{12}\|y_{k}\|^{2}+\frac{6L_{V}^{2}\xi_{k}^{2}}{(1-\gamma)\beta_{k}}\|\bar{H}(\pi^{\star}(\phi_{\omega_{k}},\hat{\mu}_{k}),\omega_{k},\hat{\mu}_{k})-\bar{H}(\pi^{\star}(\phi_{\omega_{k}}),\omega_{k},\mu^{\star}(\phi_{\omega_{k}}))\|^{2}$
$\hskip 20.0pt+\frac{6L_{V}^{2}\xi_{k}^{2}}{(1-\gamma)\beta_{k}}\|\bar{H}(\pi_{\theta_{k}},\omega_{k},\hat{\mu}_{k})-\bar{H}(\pi^{\star}(\phi_{\omega_{k}},\hat{\mu}_{k}),\omega_{k},\hat{\mu}_{k})\|^{2}$
$\leq\frac{(1-\gamma)\beta_{k}}{12}\|y_{k}\|^{2}+\frac{6L_{V}^{2}\xi_{k}^{2}}{(1-\gamma)\beta_{k}}\cdot L_{H}^{2}\Big(\|\hat{\mu}_{k}-\mu^{\star}(\phi_{\omega_{k}})\|^{2}+L^{2}\|\hat{\mu}_{k}-\mu^{\star}(\phi_{\omega_{k}})\|^{2}\Big)$
$\hskip 20.0pt+\frac{6L_{V}^{2}L_{H}^{2}\xi_{k}^{2}}{(1-\gamma)\beta_{k}}\|\pi_{\theta_{k}}-\pi^{\star}(\phi_{\omega_{k}},\hat{\mu}_{k})\|^{2}$
$\leq\frac{(1-\gamma)\beta_{k}}{12}\|y_{k}\|^{2}+\frac{6L_{V}^{2}L_{H}^{2}\xi_{k}^{2}}{(1-\gamma)\beta_{k}}\varepsilon_{k}^{\pi}+\frac{12L^{2}L_{V}^{2}L_{H}^{2}\xi_{k}^{2}}{(1-\gamma)\beta_{k}}\varepsilon_{k}^{\mu},$ (114)
where the third inequality follows from the fact that $\bar{H}(\pi^{\star}(\phi),\omega_{k},\mu^{\star}(\phi))=0$.
Finally, for the last term of (109), we have in expectation
$\xi_{k}\mathbb{E}[\langle y_{k},\Big(\nabla_{\mu}V_{f}^{\pi_{\theta^{z}_{k+1}},\phi_{\omega^{z}_{k+1}},\,\hat{\mu}^{z}_{k+1}}\Big)^{\top}\Big(H(\hat{\mu}_{k},\bar{s}_{k})-\bar{H}(\pi_{\theta_{k}},\omega_{k},\hat{\mu}_{k})\Big)\rangle]$
$=\xi_{k}\mathbb{E}[\langle y_{k},\Big(\nabla_{\theta}V_{f}^{\pi_{\theta^{z}_{k+1}},\phi_{\omega^{z}_{k+1}},\,\hat{\mu}^{z}_{k+1}}-\nabla_{\theta}V_{f}^{\pi_{\theta_{k}},\phi_{\omega_{k}},\,\hat{\mu}_{k}}\Big)^{\top}\Big(H(\hat{\mu}_{k},\bar{s}_{k})-\bar{H}(\pi_{\theta_{k}},\omega_{k},\hat{\mu}_{k})\Big)]$
$\hskip 20.0pt+\xi_{k}\mathbb{E}[\langle y_{k},\Big(\nabla_{\theta}V_{f}^{\pi_{\theta_{k}},\phi_{\omega_{k}},\,\hat{\mu}_{k}}\Big)^{\top}\Big(H(\hat{\mu}_{k},\bar{s}_{k})-\bar{H}(\pi_{\theta_{k}},\omega_{k},\hat{\mu}_{k})\Big)]$
$=\xi_{k}\mathbb{E}[\langle y_{k},\Big(\nabla_{\theta}V_{f}^{\pi_{\theta^{z}_{k+1}},\phi_{\omega^{z}_{k+1}},\,\hat{\mu}^{z}_{k+1}}-\nabla_{\theta}V_{f}^{\pi_{\theta_{k}},\phi_{\omega_{k}},\,\hat{\mu}_{k}}\Big)^{\top}\Big(H(\hat{\mu}_{k},\bar{s}_{k})-\bar{H}(\pi_{\theta_{k}},\omega_{k},\hat{\mu}_{k})\Big)]$
$\leq 2B_{H}\xi_{k}\mathbb{E}[\|y_{k}\|\|\nabla_{\theta}V_{f}^{\pi_{\theta^{z}_{k+1}},\phi_{\omega^{z}_{k+1}},\,\hat{\mu}^{z}_{k+1}}-\nabla_{\theta}V_{f}^{\pi_{\theta_{k}},\phi_{\omega_{k}},\,\hat{\mu}_{k}}\|]$
$\leq 2B_{H}L_{VV}\xi_{k}\mathbb{E}[\|y_{k}\|\Big(\|\pi_{\theta^{z}_{k+1}}-\pi_{\theta_{k}}\|+\|\phi_{\omega^{z}_{k+1}}-\phi_{\omega_{k}}\|+\|\hat{\mu}_{k+1}^{z}-\hat{\mu}_{k}\|\Big)]$
$\leq 2B_{H}L_{VV}\xi_{k}\mathbb{E}[\|y_{k}\|\Big(\|\pi_{\theta_{k+1}}-\pi_{\theta_{k}}\|+\|\phi_{\omega_{k+1}}-\phi_{\omega_{k}}\|+\|\hat{\mu}_{k+1}-\hat{\mu}_{k}\|\Big)]$
$\leq 2B_{H}L_{VV}\xi_{k}\mathbb{E}[\|y_{k}\|]\Big(\alpha_{k}B_{F}+\zeta_{k}B_{D}+\xi_{k}B_{H}\Big)$
$\leq B_{H}L_{VV}(B_{F}+B_{D}+B_{H})\alpha_{k}^{2}\mathbb{E}[\|y_{k}\|^{2}]+B_{F}L_{VV}(B_{F}+B_{D}+B_{H})\alpha_{k}^{2}.$ (115)
Substituting (110)-(115) into (109), we have
$\mathbb{E}[\langle\hat{V}_{f,k}-V_{f}^{\pi_{\theta_{k}},\phi_{\omega_{k}},\,\hat{\mu}_{k}}+\beta_{k}\bar{G}_{f}(\theta_{k},\omega_{k},\hat{\mu}_{k},\hat{V}_{f,k}),V_{f}^{\pi_{\theta_{k}},\phi_{\omega_{k}},\,\hat{\mu}_{k}}-V_{f}^{\pi_{\theta_{k+1}},\phi_{\omega_{k+1}},\,\hat{\mu}_{k+1}}\rangle]$
$\leq\frac{(1-\gamma)\beta_{k}}{12}\mathbb{E}[\|y_{k}\|^{2}]+\frac{6L_{V}^{2}\alpha_{k}^{2}}{(1-\gamma)\beta_{k}}\mathbb{E}[\|\nabla_{\theta}J_{f}(\pi_{\theta_{k}},\phi_{\omega_{k}},\hat{\mu}_{k})\|^{2}]+\frac{6L_{V}^{4}\alpha_{k}^{2}}{(1-\gamma)\beta_{k}}\mathbb{E}[\varepsilon_{f,k}^{V}]$
$\hskip 20.0pt+B_{F}L_{VV}(B_{F}+B_{D}+B_{H})\alpha_{k}^{2}\mathbb{E}[\|y_{k}\|^{2}]+B_{F}L_{VV}(B_{F}+B_{D}+B_{H})\alpha_{k}^{2}$
$\hskip 20.0pt+\frac{(1-\gamma)\beta_{k}}{12}\mathbb{E}[\|y_{k}\|^{2}]+\frac{6L_{V}^{2}\eta_{2}^{2}\zeta_{k}^{2}}{(1-\gamma)\beta_{k}}\mathbb{E}[\varepsilon_{k}^{\phi}]+\frac{36L_{V}^{4}L_{D}^{2}\zeta_{k}^{2}}{(1-\gamma)\beta_{k}}\mathbb{E}[\varepsilon_{k}^{\mu}]+\frac{24L_{V}^{2}L_{D}^{2}\zeta_{k}^{2}}{(1-\gamma)\beta_{k}}\mathbb{E}[\varepsilon_{l,k}^{V}]$
$\hskip 20.0pt+B_{D}L_{VV}(B_{F}+B_{D}+B_{H})\alpha_{k}^{2}\mathbb{E}[\|y_{k}\|^{2}]+B_{F}L_{VV}(B_{F}+B_{D}+B_{H})\alpha_{k}^{2}$
$\hskip 20.0pt+\frac{(1-\gamma)\beta_{k}}{12}\mathbb{E}[\|y_{k}\|^{2}]+\frac{6L_{V}^{2}L_{H}^{2}\xi_{k}^{2}}{(1-\gamma)\beta_{k}}\mathbb{E}[\varepsilon_{k}^{\pi}]+\frac{12L^{2}L_{V}^{2}L_{H}^{2}\xi_{k}^{2}}{(1-\gamma)\beta_{k}}\mathbb{E}[\varepsilon_{k}^{\mu}]$
$\hskip 20.0pt+B_{H}L_{VV}(B_{F}+B_{D}+B_{H})\alpha_{k}^{2}\mathbb{E}[\|y_{k}\|^{2}]+B_{F}L_{VV}(B_{F}+B_{D}+B_{H})\alpha_{k}^{2}$
$\leq\frac{(1-\gamma)\beta_{k}}{4}\mathbb{E}[\|y_{k}\|^{2}]+\frac{6L_{V}^{2}\alpha_{k}^{2}}{(1-\gamma)\beta_{k}}\mathbb{E}[\|\nabla_{\theta}J_{f}(\pi_{\theta_{k}},\phi_{\omega_{k}},\hat{\mu}_{k})\|^{2}]+\frac{6L_{V}^{2}L_{H}^{2}\xi_{k}^{2}}{(1-\gamma)\beta_{k}}\mathbb{E}[\varepsilon_{k}^{\pi}]+\frac{6L_{V}^{2}\eta_{2}^{2}\zeta_{k}^{2}}{(1-\gamma)\beta_{k}}\mathbb{E}[\varepsilon_{k}^{\phi}]$
$\hskip 20.0pt+\left(\frac{36L_{V}^{4}L_{D}^{2}\zeta_{k}^{2}}{(1-\gamma)\beta_{k}}+\frac{12L^{2}L_{V}^{2}L_{H}^{2}\xi_{k}^{2}}{(1-\gamma)\beta_{k}}\right)\mathbb{E}[\varepsilon_{k}^{\mu}]+\frac{6L_{V}^{4}\zeta_{k}^{2}}{(1-\gamma)\beta_{k}}\mathbb{E}[\varepsilon_{f,k}^{V}]+\frac{24L_{V}^{2}L_{D}^{2}\alpha_{k}^{2}}{(1-\gamma)\beta_{k}}\mathbb{E}[\varepsilon_{l,k}^{V}]$
$\hskip 20.0pt+L_{VV}(B_{F}+B_{D}+B_{H})^{2}\alpha_{k}^{2}\mathbb{E}[\|y_{k}\|^{2}]+L_{VV}(B_{F}+B_{D}+B_{H})^{2}\alpha_{k}^{2}$
$\leq\frac{(1-\gamma)\beta_{k}}{2}\mathbb{E}[\|y_{k}\|^{2}]+\frac{6L_{V}^{2}\alpha_{k}^{2}}{(1-\gamma)\beta_{k}}\mathbb{E}[\|\nabla_{\theta}J_{f}(\pi_{\theta_{k}},\phi_{\omega_{k}},\hat{\mu}_{k})\|^{2}]+\frac{6L_{V}^{2}L_{H}^{2}\xi_{k}^{2}}{(1-\gamma)\beta_{k}}\mathbb{E}[\varepsilon_{k}^{\pi}]+\frac{6L_{V}^{2}\eta_{2}^{2}\zeta_{k}^{2}}{(1-\gamma)\beta_{k}}\mathbb{E}[\varepsilon_{k}^{\phi}]$
$\hskip 20.0pt+\frac{16L^{2}L_{V}^{2}L_{H}^{2}\xi_{k}^{2}}{(1-\gamma)\beta_{k}}\mathbb{E}[\varepsilon_{k}^{\mu}]+\frac{6L_{V}^{4}\zeta_{k}^{2}}{(1-\gamma)\beta_{k}}\mathbb{E}[\varepsilon_{f,k}^{V}]+\frac{24L_{V}^{2}L_{D}^{2}\alpha_{k}^{2}}{(1-\gamma)\beta_{k}}\mathbb{E}[\varepsilon_{l,k}^{V}]+L_{VV}(B_{F}+B_{D}+B_{H})^{2}\alpha_{k}^{2},$
where the last inequality is due to the step size condition $\frac{\zeta_{k}}{\xi_{k}}\leq\frac{LL_{H}}{3L_{V}L_{D}}$ and $\alpha_{k}\leq\min\{1,\frac{1-\gamma}{4L_{VV}(B_{F}+B_{D}+B_{H})^{2}}\}$.
$\blacksquare$
Appendix E Details on Example Satisfying Assumption 6
Example 1
[…]
${\cal P}_{f}^{\mu}(s_{f}^{\prime}\mid s_{f},a,b)={\mathbf{1}}(s_{f}^{\prime}=1),\qquad r_{l}(b,\mu)={\mathbf{1}}(b=1)+\mu(b),$
where ${\mathbf{1}}(\cdot)$ denotes the indicator function (1 if the condition holds, 0 otherwise). We drop $s_{l}$ as an argument from $r_{l}$, since the leader’s state is vacuous. The follower’s reward $r_{f}$ can be any function of $s_{f},a,b$. This MFG satisfies Assumption 6 with $\eta_{1}=2,\eta_{2}=1/2$. Clearly, the leader-follower independence assumption in Cui et al. [2024] is violated as the leader’s reward depends on the mean field and the follower’s reward may depend on the leader’s action.
Below we justify that Example 1 indeed satisfies Assumption 6 but not the leader-follower independence assumption in Cui et al. [2024].
The mean field induced by any leader’s policy is the same and satisfies the following
$\mu^{\star}(\phi)(s)={\mathbf{1}}(s=1).$
As the leader is stateless, we have
$\Phi(\phi)$ $=\sum_{b\in{\cal B}}\phi(b)r_{l}(b,\mu^{\star}(\phi))=\sum_{b\in{\cal B}}\phi(b)\Big({\mathbf{1}}(b=1)+\mu^{\star}(\phi)(b)\Big)$
$=\sum_{b\in{\cal B}}\phi(b)\Big({\mathbf{1}}(b=1)+{\mathbf{1}}(b=1)\Big)=2\phi(1).$
As a result,
$\nabla_{\omega}\Phi(\phi_{\omega})=2\nabla_{\omega}\phi_{\omega}(1).$
Similarly, we have
$J_{l}(\phi,\mu)=\sum_{b\in{\cal B}}\phi(b)\Big({\mathbf{1}}(b=1)+\mu(b)\Big)=\phi(1)+\sum_{b}\phi(b)\mu(b),$
which leads to
$\nabla_{\omega}J_{l}(\phi_{\omega},\mu)\mid_{\mu=\mu^{\star}(\phi_{\omega})}=\nabla_{\omega}\phi_{\omega}(1).$
Since $\nabla_{\omega}\Phi(\phi_{\omega})=2\nabla_{\omega}J_{l}(\phi_{\omega},\mu)\mid_{\mu=\mu^{\star}(\phi_{\omega})}$, the inequalities in Assumption 6 obviously hold with $\eta_{1}=2$ and $\eta_{2}=1/2$.
Appendix F Appendix: Environment Definitions
F.1 Market Entrance
Followers Followers decide whether or not to buy (left) or sell (right) $A=\{\text{buy, sell}\}$ a good at each timestep, leading to the corresponding state $s_{k+1}=a_{k},S=A$. The payoff for followers is based on the proportion of other agents choosing the same action:
$r_{f}(a_{k},b_{k},s_{k},\mu_{k})=-\mu_{k}(a_{k})+\text{bonus}*\mathbbm{1}_{a_{k}=b_{k}},$
where a bonus is awarded for choosing the same action as the leader, as discussed below. The transition kernel is independent of the mean field and a deterministic function of the action of the representative agent:
$\mathcal{P}(s_{k+1}=s^{\prime}|s_{k}=s,a_{k}=a)=\mathbbm{1}_{s^{\prime}=a}.$
Leader We introduce a leader into this market, where the leader is trying to steer the mean field towards some desirable market state goal, for example, buying ($\text{goal}=\text{buy}$), selling ($\text{goal}=\text{sell}$), or balanced market ($\text{goal}\in_{R}\{\text{buy},\text{sell}\}$), by taking actions to bonus these states $b_{k}\in B=\{\text{buy},\text{sell}\}$. This can be seen as, for example, a market maker skewing prices to make certain trades more desirable, e.g., for managing their inventory. The reward for the leader is:
$r_{l}(s_{k},b_{k},\mu_{k})=\mu_{k}(\text{goal})+\upsilon(b_{k}-\text{goal}).$
where $0<\upsilon\ll 1$ is a small reward shaping term, providing more immediate feedback to the leader than the delayed feedback from the mean field.
F.2 Shop position
Figure 5: Beach Bar Environment. The blue stars indicates the fixed bar position, and the golden star shows the leader can add a new bar anywhere within the existing state.
The state space $S$ covers all the potential locations in a 1-dimensional ring, where there is a (fixed) desirable location $\star$ in one of these states, as visualized in Fig.˜5.
Followers The followers are rewarded based on their proximity to the nearest desirable location (either $b_{k}$ or $\star$), as well as the crowdedness of the state they are currently in:
$r_{f}(s_{k},a_{k},b_{k},\mu_{k})=\min(\text{dist}(s_{k},\star),\text{dist}(s_{k},b_{k}))-c\times\log(\mu_{k}(s_{k}))$
where $c\geq 0$ is a coefficient specifying the follower’s preference to avoid the crowd.
Leader The leader is rewarded based on the business of their chosen location: $r_{l}(s_{k},b_{k},\mu_{k})=\mu_{k}(b_{k})$ e.g., more customers visiting a store can lead to higher profits, making the positioning more desirable.
F.3 Equilibrium price
In this environment, the followers are homogeneous firms producing an identical good, where the price is determined by the (endogenous) supply-demand equilibrium around some baseline demand $d$. Each timestep, the representative firm chooses a production quantity $q_{k}$ and replenishment quantity $h_{k}$ based on their inventory level captured by the state $s_{k}$. The state/inventory transitions according to: $s_{k+1}=s_{k}-\text{min}(q_{k},s_{k})+h_{k}$.
The followers are rewarded based on the resulting (endogenous) price and their chosen quantities as:
$r_{f}(s,a,b,\mu)=(p_{k}-c_{0})q_{k}-c_{1}q_{k}^{2}-c_{2}h_{k}-(c_{2}+c_{3})\text{max}(q_{k}-s_{k},0)-c_{4}s_{k}$ (116)
where $c_{x}$ are cost terms penalizing different inventory management and production capacities, and the price is determined based on the supply-demand equilibrium as $p_{k}=\frac{d}{E[q]}^{1/\sigma}$ where
$E[q]=\sum_{\text{inv}}\sum_{q}\sum_{h}\mu(\text{inv})\cdot\pi[(q,h),\text{inv}]\cdot\text{inv}$
is the expected inventory of the mean field.
We introduce a leader into this environment, where the leader attempts to encourage firms to maintain some target inventory level $i$ by taking actions to set the inventory holding cost $c_{2}$. For example, a large target inventory level would encourage firms to build buffers to withstand future supply shocks, or an inventory of zero would encourage optimal production under known and predictable demands. The leader is rewarded by
$r_{F}=-(E[q]-i)^{2}+\upsilon[-c_{2}\cdot(E[q]-i)]$ (117)
where the first term is based on the inventory gap, and the second term is again a reward shaping term for more immediate feedback.
Appendix G Additional Experimental Results
We provide the evolution of the mean fields and resulting metrics across the environments in Figs.˜6, 7 and 8, for market entrance, shop positioning, and equilibrium price respectively.
(a) Goal=buy
(b) Goal=balance
(c) Goal=sell
Figure 6: Market Entrance: Evolution of follower mean fields $\mu$ (top row) and resulting skew (bottom row) under different leader goals.
(a) $c=0$
(b) $c=2.5$
(c) $c\to\infty$
Figure 7: Beach bar: Evolution of follower mean field $\mu$ (top row) and distance between the chosen bar position $b$ (leaders action) and the fixed bar position $\star$ (bottom row) under varying crowd aversions $c$.
(a) Lean (Minimise inventory)
(b) Medium target inventory
(c) Robust (Maximize inventory)
Figure 8: Equilibrium pricing: Evolution of follower mean fields $\mu$ (top row) and resulting inventories (bottom row) under different leader goals.
Appendix H Experimental Details
We run each approach for $20,000$ iterations (plots show every 10th iteration). Each run is repeated 30 times, with seeds set based on the run index for reproducibility.
Proposed. The proposed is run with $\zeta_{0}=0.5,\alpha_{0}=0.25,\beta_{0}=0.02,\xi_{0}=0.25$.
PPO. For the PPO implementation, we base the implementation off CleanRL (in Torch), using a batch size of 256, hidden layer shape of (64, 64), learning rate of 3e43e-4, TanH activation functions for the hidden layers, and a clipping epsilon of 0.2. ADAM is used as the optimiser (as implemented in torch).
Nested. To ensure fair comparisons, the nested comparisons are run with the same hyperparamaters as the proposed. However, follower (resp. leader) specific parameters are decayed at a rate based only on the number of follower (resp. leader) updates.
All approaches are run on a CPU, with Python3, and on an Amazon EC2 with R6i.large.